### v2

# Discrete Approximate Circle Bundles

Abstract
In this paper, we introduce discrete approximate circle bundles, a class of objects designed to serve as the data science analog of circle bundles from algebraic topology. We show that, under appropriate conditions, one can meaningfully and stably identify a discrete approximate circle bundle with an isomorphism class of true circle bundles. We also describe two cohomology invariants which uniquely determine the isomorphism class of a circle bundle, and provide algorithms to compute them given a discrete approximate representative. Finally, we propose a novel methodology for coordinatization and dimensionality reduction of circle bundle data. To illustrate the practical utility and viability of our algorithms, we present applications to both real and synthetic datasets from computer vision (e.g., modeling optical flow).
1 Background And Problem Statement
1.1 Motivation And Context
In recent years, it has been observed that high-dimensional datasets arising in fields such as computer vision, computational chemistry, and motion tracking often lie near highly non-linear, low-dimensional manifolds [TSL00, RS00, BN03]. These manifolds can exhibit complex geometric and topological structure, motivating the development of a range of techniques for identifying and leveraging such structures in tasks including dimensionality reduction, statistical analysis, pattern recognition, and data visualization [CL06, Sap01]. By modeling the intrinsic manifold geometry, these methods have been shown to improve compression, denoising, and feature extraction in domains where traditional linear techniques fail to capture the underlying structure of the data.
In this paper, we focus on a particular class of datasets which can be modeled as circle bundles. Roughly speaking, a circle bundle is a continuously-varying family of (topological) circles parameterized by another space, called the base space of the bundle. The individual circles are called fibers of the bundle, and the union of the fibers is the total space. The parametrization is given by a projection map $\pi:E\to B$ from the total space to the base space – $E$ is locally homeomorphic to the product $B\times\mathbb{S}^{1}$, but globally the structure may be twisted. For instance, the torus and the Klein bottle can each be characterized as the total space of a circle bundle with base space $\mathbb{S}^{1}$, but the two are not homeomorphic (see Figure 1). Similarly, the unit tangent bundle on the 2-sphere is a circle bundle with base $\mathbb{S}^{2}$, but one can show that there is no bundle isomorphism between this bundle and the product space $\mathbb{S}^{2}\times\mathbb{S}^{1}$.
[…]
Figure 1: The Klein bottle $K$ as a circle bundle over $\mathbb{S}^{1}$. The highlighted region $\pi^{-1}(U)$ in $K$ shows all fibers above the neighborhood $U$ in the base space and is homeomorphic to $U\times\mathbb{S}^{1}$. Globally, however, $K$ is not homeomorphic to $\mathbb{S}^{1}\times\mathbb{S}^{1}$.
On the machine learning side, various datasets related to computer vision and graphics have been shown to have circle bundle structures. For instance, in [Ada+20], Adams et al. use local measurements and a feature map to support a torus model for high-contrast optical flow data. Similarly, in [Car+08], the authors propose that a certain space of high-contrast natural image patches is concentrated around a Klein bottle embedded in high dimension. Beyond these examples, datasets arising from 3D objects with rotational symmetries — such as projection images of macromolecules in cryo-electron microscopy — naturally give rise to circle bundle structures, where each fiber corresponds to rotations of the object about a fixed axis [SS14].
The possible circle bundle structures over a metric space $B$ are completely described by a pair of discrete invariants called characteristic classes. In particular, a non-vanishing class indicates the bundle is non-trivial, so there is no hope of constructing a coordinatization for the whole space as a product. Additionally, characteristic classes are invariant under continuous perturbations, so one might hope to be able to stably recover them from a noisy finite sampling. Another key feature of characteristic classes is that they can be computed from local measurements alone, which makes them an ideal tool for working with complex, high-dimensional data not amenable to more direct techniques. Lastly, the characteristic classes of a bundle can be used to construct a so-called bundle map from the total space to a more familiar space which respects topological structure. This observation has recently been leveraged in works such as [Lee+25] and [SP23] to develop non-linear dimensionality reduction techniques based on vector bundles. In this paper, we extend this approach to circle bundles.
1.2 Our Contributions
Here we summarize our contributions:
  1. 1.
    We define notions of discrete approximate circle bundles and local trivializations and show how these relate to other well-studied objects such as discrete approximate Čech cocycles. Building on results of [Tin22] and [SP23], we give conditions under which a discrete approximate circle bundle can be meaningfully identified with a true class of circle bundles.
  2. 2.
    We show that every true circle bundle is uniquely characterized by an orientation class and a twisted Euler class, and provide an algorithm for computing these classes from a family of discrete approximate local trivializations. We further introduce a metric on each space of discrete approximate local trivializations and show that two trivializations which are sufficiently close produce the same characteristic classes. In particular, we prove that if a discrete approximate local trivialization is sufficiently close to a true one, the characteristic classes obtained from our algorithm are the same.
  3. 3.
    Given a discrete approximate circle bundle over a base space $B$ and an open cover $\mathcal{U}$ for $B$, we introduce a weights filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and a corresponding notion of persistence for characteristic class representatives. We then propose a scheme for coordinatization and dimensionality reduction which produces a map from the dataset to the total space of the associated bundle $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, where $V(2,d)$ denotes the Stiefel manifold of 2-planes in $\mathbb{R}^{d}$. Our pipeline can be tuned to be compatible with any stage of the weights filtration on $\mathcal{N}(\mathcal{U})$, yielding a point cloud with a different global topology at each stage.
  4. 4.
    We provide implementations of our algorithms on several real and synthetic datasets related to imaging and optical flow. In Section 7.1, we use these tools to confirm the torus model proposed in [Ada+20]. A more detailed analysis of our findings will appear in an upcoming companion paper.
1.3 Prior Work
A number of approaches to local-to-global inference for analyzing geometrically and topologically complex datasets have recently emerged. The basic principle in each case is to model small pieces of the dataset and analyze how these parts glue together to determine the global shape. Given a dataset $X$, a common technique is to choose a feature map $\pi:X\to S$ to some metric space $S$ (e.g., $\mathbb{R}$) and study how the level sets vary. The Mapper pipeline takes this approach to represent a point cloud as a graph with nodes representing local clusters. Similarly, in [SP22] the authors describe a pipeline called FibeRed which uses local PCA together with a feature map $\pi:X\to S$ to obtain a low-dimensional representation of $X$ which is faithful to both local geometry and global topology.
Another aspect of local-to-global analysis is the synchronization of local coordinate systems. Cohomological obstructions to synchronization are explored in works such as [GBM19] using the language of principal $G$-bundles. We note, however, that the theory provided there makes the highly restrictive assumption that the underlying bundle is flat. In [SP23], the authors introduce a theory of discrete approximate Euclidean vector bundles which handles non-flat cases in the context of $O(d)$-synchronization. Building on the language and technology developed in [Tin22], they show that under good conditions, one can uniquely and stably identify discrete approximate Euclidean vector bundles with true ones. They also provide stable algorithms for computing characteristic classes of vector bundles from discrete approximate cocycle representatives.
1.4 Structure Of Paper
The rest of the paper is structured as follows:
Section 3: Mathematical Background. We review the essential mathematical foundations required for our results, including the general theory of fiber bundles, the classification of circle bundles and cohomology with local coefficients. This section establishes notation and key concepts that will be used throughout the paper.
Section 4: Discrete Approximate Circle Bundles. We introduce the notion of discrete approximate circle bundles and explore their relationship with discrete approximate Čech cocycles and classifying maps as defined in [SP23]. We also discuss how these discrete constructions relate to classical circle bundles.
Section 5: Algorithms for Characteristic Classes. We present algorithms to compute characteristic classes from discrete approximate local trivializations, and prove that our algorithms are stable.
Section 6: Weights Filtration and Persistence. Given a discrete approximate circle bundle over a metric space $B$ and an open cover $\mathcal{U}$ of $B$, we introduce a simplexwise filtration on $\mathcal{N}(\mathcal{U})$ and describe a type of persistent cohomology for characteristic class representatives with respect to this filtration.
Section 7: Coordinatization Pipeline. We provide a pipeline for coordinatization and dimensionality reduction which integrates the Principal Stiefel Coordinates algorithm [Lee+25]. The result is a map from the dataset into a Stiefel manifold which is, in some sense which we explain, compatible with a chosen stage of the associated weights filtration.
Section 8: Implementations and Experiments. We demonstrate our methods on both synthetic and real datasets, applying the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations and showcasing the resulting computations of characteristic classes and coordinatization.
Section 9: Future Work. We conclude with a discussion of potential extensions, open questions, and directions for further exploration in both theory and applications.
2 Preliminaries
2.1 Theory Of Fiber Bundles
In this section, we briefly review the key aspects of the theory of fiber bundles which are relevant for us. Our exposition emphasizes a computational perspective. For additional details, see [Hus94].
Definition 2.1.
Let $F$ be a topological space. A fiber bundle with fiber $F$ is a triple $\xi=(E,\pi,B)$, where $E$ and $B$ are topological spaces and $\pi:E\to B$ is a continuous surjection with the following property: for each $b\in B$, there exists an open neighborhood $U_{b}\subseteq B$ of $b$ and a homeomorphism $\varphi:\pi^{-1}(U_{b})\to U_{b}\times F$ such that if $p_{1}:U_{b}\times F\to U_{b}$ is the projection onto the first factor, then the following diagram commutes:
$\pi^{-1}(U_{b})$$U_{b}\times F$$U_{b}$$\scriptstyle{\varphi}$$\scriptstyle{\pi}$$\scriptstyle{p_{1}}$
Any such map $\varphi$ is called a local trivialization of $\xi$. An open cover $\mathcal{U}$ of $B$ is called a trivializing cover for $\xi$ if there exists a local trivialization of $\xi$ over each $U_{j}\in\mathcal{U}$. Two bundles $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are isomorphic if there exists a homeomorphism $f:E\to E^{\prime}$ such that $\pi^{\prime}\circ f=\pi$. Let $\text{Bun}_{F}(B)$ denote the set of isomorphism classes.
Transition Maps: Given a fiber bundle $\xi=(E,\pi,B)$ with fiber $F$, an open cover $\mathcal{U}$ of $B$ and a family $\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times F\}_{j\in J}$ of local trivializations subordinate to $\mathcal{U}$, there exist unique continuous maps $\Omega_{jk}:U_{j}\cap U_{k}\to\text{Homeo}(F)$ such that $\varphi_{j}\circ\varphi_{k}^{-1}(b,v)=(b,\Omega_{jk}(b)(v))$ for any $(b,v)\in(U_{j}\cap U_{k})\times F$. The maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ are called the transition maps associated with the family of local trivializations. Note that the transition functions associated with any family of local trivializations must satisfy the cocycle condition:
$\Omega_{jk}(b)\circ\Omega_{kl}(b)=\Omega_{jl}(b)\hskip 28.45274pt(b\in U_{jkl}\;,\;(jkl)\in\mathcal{N}(\mathcal{U}))$ (1)
If we require the transition maps to be $G$-valued for some subgroup $G\leq\text{Homeo}(F)$, we say that $G$ is the structure group of the bundle.
Fiber Bundle Construction Lemma: The following well-known lemma essentially states that a fiber bundle is completely determined up to isomorphism by the transition maps associated with a family of local trivializations. It also gives a computationally-friendly way to determine whether two bundles over the same space are isomorphic:
Part A: Let $B$ and $F$ be topological spaces, and let $G$ be a topological group with a continuous left action on $F$. Given an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a collection $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ satisfying the cocycle condition, there exists a fiber bundle $\pi:E\to B$ with fiber $F$ and structure group $G$ that is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$.
Part B: Suppose $\xi^{\prime}=(E^{\prime},\pi^{\prime},B^{\prime})$ is another fiber bundle with fiber $F$ and structure group $G$ which is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}^{\prime}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$. If the action of $G$ on $F$ is faithful, then $\xi\cong\xi^{\prime}$ if and only if there exist maps $\{\mu_{j}:U_{j}\to G\}_{j\in J}$ such that $\Omega_{jk}^{\prime}=\mu_{j}^{-1}\Omega_{jk}\mu_{k}$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
Note that if $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are fiber bundles with fiber $F$ over a common paracompact space $B$ with trivializing covers $\mathcal{U}$ and $\mathcal{U}^{\prime}$, then any common refinement $\mathcal{V}$ of $\mathcal{U}$ and $\mathcal{U}^{\prime}$ is a trivializing cover for $\xi$ and $\xi^{\prime}$, so we can compare the two bundles using this refinement (and the result does not depend on the choice of refinement).
Associated Bundles: Let $\xi=(E,\pi,B)$ be a fiber bundle with fiber $F$ and structure group $G$, and let $F^{\prime}$ be a left $G$-space. The associated bundle to $\xi$ with fiber $F^{\prime}$ is the bundle $\xi^{\prime}=(E\times_{G}F^{\prime},\pi^{\prime},B)$ whose existence is guaranteed by the fiber bundle construction lemma (using the transition maps from any trivialization of $\xi$). The special case of $F^{\prime}=G$ with left action by group multiplication is called the principal $G-\textbf{bundle}$ associated with $\xi$. Note that for any space $B$, the associated bundle construction gives a bijective correspondence between fiber bundles with fiber $F$ and principal $\text{Homeo}(F)$-bundles over $B$.
If $\xi=(E,\pi,B)$ is a principal $G$-bundle and $\varphi:G\to H$ is a continuous group homomorphism, then $H$ has a continuous left $G$-action given by $g\cdot h=\varphi(g)h$. The associated bundle, denoted by $\varphi_{*}\xi$, is then a principal $H$-bundle. In particular, if $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ are transition maps for $\xi$ subordinate to some open cover $\mathcal{U}$ of $B$, then $\{\varphi\circ\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ are transition maps for $\varphi_{*}\xi$.
Čech Cohomology: The cocycle condition (1) and Part B of the Fiber Bundle Construction Lemma suggest a cohomological perspective to fiber bundle classification. This perspective can be articulated using the language of Čech cohomology, which we briefly describe below.
Let $\mathcal{U}$ be an open cover of a topological space $B$ and let $\mathcal{F}$ be a presheaf of abelian groups on $B$. For $k\geq 0$, define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}:\check{C}^{k}(\mathcal{U};\mathcal{F})\to\check{C}^{k+1}(\mathcal{U};\mathcal{F})$ by
$\check{C}^{k}(\mathcal{U};\mathcal{F})=\prod_{\alpha_{0}<\cdots<\alpha_{k}}\mathcal{F}(U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k}})$
$(\delta^{k}\Omega)_{\alpha_{0}\cdots\alpha_{k+1}}=\sum_{j=0}^{k+1}(-1)^{j}\Omega_{\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}}|_{U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k+1}}}$
Then $(\check{C}^{*}(\mathcal{U};\mathcal{F}),\delta^{*})$ is a cochain complex, and the associated cohomology is called Čech cohomology subordinate to $\mathcal{U}$ with coefficients in $\mathcal{F}$.
Note that if $\mathcal{F}$ is a presheaf of non-abelian groups, we can still define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}$ as above, replacing negation by group inversion in the coboundary formula. In general, however, we do not have $\ker\delta^{k+1}\subseteq\text{im }\delta^{k}$ so $\check{H}^{k}(\mathcal{U};\mathcal{F})$ is not well-defined for $k>0$. For $k=1$, we still have a well-defined cohomology set $\check{H}^{1}(\mathcal{U};\mathcal{F})$, though the classes do not have a natural group structure.
Let $\mathcal{F}$ be a presheaf of groups over $B$, and let $\mathcal{U}$ be an open cover of $B$. If $\mathcal{V}$ is a refinement of $\mathcal{U}$, one has an induced map $\check{H}^{*}(\mathcal{U};\mathcal{F})\to\check{H}^{*}(\mathcal{V};\mathcal{F})$ in cohomology. The open covers of $B$ form a directed set with respect to refinements, so the corresponding Čech cohomology groups form a direct system of abelian groups. The direct limit is denoted $\check{H}^{*}(B;\mathcal{F})$ and called the Čech cohomology of $B$. If $B$ is paracompact, then for any topological group $G$, the fiber bundle construction lemma implies a one-to-one correspondence between $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(F)})$ and $\text{Bun}_{F}(B)$, where $\mathcal{C}_{\text{Homeo}(F)}$ denotes the sheaf of continuous $\text{Homeo}(F)$-valued functions on $B$.
We state the following additional properties of Čech cohomology without proof. For further details, see [Bre97].
Proposition 2.2.
Čech cohomology has the following properties:
  1. 1.
    If $\mathcal{U}$ is a good open cover of $B$, then $\check{H}^{*}(\mathcal{U};\mathcal{C}_{G})\cong\check{H}^{*}(B;\mathcal{C}_{G})$ for any topological group $G$.
  2. 2.
    If $B$ has the homotopy type of a CW complex and $G$ is abelian, then $\check{H}^{*}(B;\underline{G})\cong H^{*}(B;G)$, where $\underline{G}$ denotes the sheaf of locally constant $G$-valued functions on $B$.
  3. 3.
    If $\mathcal{F}$ is a flabby sheaf over a paracompact space $B$, then $\check{H}^{k}(B;\mathcal{F})=0$ for all $k>0$.
  4. 4.
    If $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced cochain map $\varphi_{*}:\check{Z}^{*}(B;\mathcal{C}_{G})\to\check{Z}^{*}(B;\mathcal{C}_{H})$ given by $(\varphi_{*}\Omega)_{\alpha_{0}\cdots\alpha_{k}}=\varphi(\Omega_{\alpha_{0}\cdots\alpha_{k}})$. Moreover, if $\varphi$ is a homotopy equivalence, then $\varphi_{*}$ induces an isomorphism in cohomology.
  5. 5.
    If $F\to G\to H$ is a central extension of topological groups, there is an induced short exact sequence $\mathcal{C}_{F}\to\mathcal{C}_{G}\to\mathcal{C}_{H}$ of sheaves, which in turn gives rise to a short exact sequence in Čech cohomology. Moreover, if $F$ is abelian, there is an induced connecting homomorphism $\Delta:\check{H}^{1}(B;\mathcal{C}_{H})\to\check{H}^{2}(B;\mathcal{C}_{F})$ which yields a long exact sequence.
Pullback Bundles: In addition to associated bundles, another key way to obtain new fiber bundles from existing ones is via the pullback construction. Given a fiber bundle $\xi=(E,\pi,B)$ and a continuous map $f:B^{\prime}\to B$, the pullback bundle $f^{*}\xi$ is the triple $(f^{*}E,\pi^{\prime},B^{\prime})$, where
$f^{*}E=\{(b^{\prime},e)\in B^{\prime}\times E:\ f(b^{\prime})=\pi(e)\}\\$
and $\pi^{\prime}:f^{*}E\to B^{\prime}$ is given by $\pi^{\prime}(b^{\prime},e)=b^{\prime}$. Note that if $\xi$ has fiber $F$ then so does $f^{*}\xi$. Moreover, one can show that homotopic maps yield isomorphic pullback bundles.
Universal Bundles: A universal $G$-bundle is a principal $G$-bundle $(EG,p,BG)$ such that $EG$ is (weakly) contractible. The space $BG$ is called a classifying space for $G$. It can be shown that a universal $G$-bundle exists for any topological group $G$. The construction is essentially unique in the sense that any two universal $G$-bundles are isomorphic and any two classifying spaces for $G$ are (weakly) homotopy equivalent.
For any paracompact space $B$, there is a bijective correspondence between $[B,BG]$ and $\text{Prin}_{G}(B)$ given by $f]\mapsto[f^{*}\xi_{G}]$, where $\xi_{G}$ is a universal $G$-bundle. Given any $\xi]\in\text{Prin}_{G}(B)$, a map $f:B\to BG$ such that $\xi\cong f^{*}\xi_{G}$ is called a classifying map for $\xi$.
Lastly, we note that the classifying space construction is functorial in the following sense: if $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced map $B\varphi:BG\to BH$ with the property that if $f:B\to BG$ classifies a principal $G$-bundle $\xi$, then $B\varphi\circ f:B\to BH$ classifies $\varphi^{*}\xi$.
2.2 Cohomology With Local Coefficients
We provide a brief overview of cohomology with local coefficients which we use in our classification of circle bundles. For additional details, see [Eil47], [Hat02].
Definition 2.3.
Given a topological space $B$ and an abelian group $\Gamma$, a local system for $B$ with coefficients in $\Gamma$ is a locally constant sheaf $\mathcal{L}$ with stalk $\Gamma$.
We will primarily be interested in the case where $\Gamma$ is a topological group and $B$ a path-connected paracompact space (in particular, $B$ has a universal cover $p:\widetilde{B}\to B$). In this setting, a local system on $B$ admits a variety of important equivalent characterizations:
Proposition 2.4.
Suppose $G$ is a topological group and $B$ is a path-connected space with universal cover $p:\widetilde{B}\to B$. A local system on $B$ with coefficients in $G$ can be equivalently defined as any one of the following:
  1. 1.
    A sheaf $\mathcal{L}$ with stalk $G$ such that $p^{*}\mathcal{L}$ is a constant sheaf on $\widetilde{B}$,
  2. 2.
    A monodromy representation $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$,
  3. 3.
    A flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$,
  4. 4.
    A Čech cocycle $\omega\in\check{Z}^{1}(B;\underline{\text{Aut}(G)})$.
Proof.
If $\mathcal{L}$ is a locally-constant sheaf, then $p^{*}\mathcal{L}$ is also locally constant. Since $\widetilde{B}$ is simply-connected, $\mathcal{L}$ must be constant. Conversely, if $\mathcal{L}$ is a sheaf on $B$ and $p^{*}\mathcal{L}$ is constant, then clearly $\mathcal{L}$ is locally-constant.
1 $\to$ 2: Since the unit interval $I=[0,1]$ is contractible, the sheaf $\gamma^{*}\mathcal{L}$ on $I$ is constant for any loop $\gamma:I\to B$ based at $b_{0}$. We therefore have an induced automorphism of $G$ by identifying $(\gamma^{*}\mathcal{L})_{0}$ with $(\gamma^{*}\mathcal{L})_{1}$, and the automorphisms induced by homotopic loops coincide. We therefore have a map $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$, and one can easily check that this map is a representation.
2 $\to$ 3: Note that $p:\widetilde{B}\to B$ is a principal $\pi_{1}(B,b_{0})$-bundle, where $\pi_{1}(B,b_{0})$ acts on $\widetilde{B}$ by deck transformations. In particular, this bundle is flat since $\pi_{1}(B,b_{0})$ is a discrete group. A monodromy representation gives a left $\pi_{1}(B,b_{0})$-action on $G$ by automorphisms, so we can form the associated bundle $\widetilde{B}\times_{\rho}G\to B$.
3 $\leftrightarrow$ 4: This equivalence is immediately implied by the fiber bundle construction lemma.
3 $\to$ 1: If $\xi$ is a flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$, the sheaf of locally-constant sections of $\xi$ is a locally-constant sheaf on $B$ with stalk $L$.
Remark 2.5.
The equivalences above all descend to the level of isomorphism classes. For example, two locally-constant sheaves on $B$ are isomorphic if and only if they induce equivalent monodromy representations.
We will primarily be interested in local systems with stalk $\mathbb{Z}$:
Definition 2.6.
A local system with stalk $\mathbb{Z}$ is called an orientation sheaf. Given a cocycle $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, we will denote the orientation sheaf associated with $\omega$ by $\mathbb{Z}_{\omega}$. In particular, if $B$ is a smooth manifold, the orientation sheaf associated with $w_{1}(TB)$ is called the orientation sheaf of the manifold.
By Proposition 2.4, we can think of an orientation sheaf over $B$ as the sheaf of locally-constant sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}\mathbb{Z}$, where $B_{\omega}$ is a double cover of $B$. More generally, given a topological group $G$ with a $\mathbb{Z}_{2}$-action, we will be interested in sheaves of sections of associated bundles of the form $B_{\omega}\times_{\mathbb{Z}_{2}}G$:
Definition 2.7.
Let $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. Given a topological group $G$ with a $\mathbb{Z}_{2}$-action, let $\mathcal{P}_{G,\omega}$ be the presheaf over $B$ defined by
$\mathcal{P}_{G,\omega}(V)=\{\text{families }\{f_{j}\in\mathcal{C}_{G}(U_{j}\cap V)\}:\ f_{j}|_{U_{j}\cap U_{k}\cap V}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}\cap V}\}$
The sheaf $\mathcal{C}_{G,\omega}$ of continuous $G$-valued functions on $B$ twisted by $\omega$ is the sheafification of $\mathcal{P}_{G,\omega}$. Note that $\mathcal{C}_{G,\omega}$ is equivalently defined as the sheaf of sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}G$, where $p_{\omega}:B_{\omega}\to B$ denotes the orientation double cover represented by $\omega$.
If $G$ is a discrete group, we will write $G_{\omega}$ instead of $\mathcal{C}_{G,\omega}$.
Proposition 2.8.
Suppose $\omega,\omega^{\prime}\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$. Then, for any topological group $G$ with left $\mathbb{Z}_{2}$-action, the sheaves $\mathcal{C}_{G,\omega}$ and $\mathcal{C}_{G,\omega^{\prime}}$ are isomorphic if any only if $\omega]=[\omega^{\prime}]$. Explicitly, if $\omega=\omega^{\prime}\cdot\delta^{0}\phi$, then the map which sends $f_{j}\in\mathcal{C}_{G,\omega}(U_{j})$ to $\phi_{j}\cdot f_{j}\in\mathcal{C}_{G,\omega^{\prime}}(U_{j})$ is an isomorphism of sheaves.
Proof.
Let $\{f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$ such that $f_{j}|_{U_{j}\cap U_{k}}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}}$ whenever $U_{j}\cap U_{k}$ is nonempty, and let $\{f^{\prime}_{j}=\phi_{j}\cdot f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$. Since $\omega=\omega^{\prime}\cdot\delta^{0}\phi$, we have
$\tilde{f}_{j}|_{U_{j}\cap U_{k}}=\phi_{j}\cdot f_{j}|_{U_{j}\cap U_{k}}=(\phi_{j}\omega_{jk})\cdot f_{k}|_{U_{j}\cap U_{k}}=\phi_{k}\cdot f_{k}|_{U_{j}\cap U_{k}}=\tilde{f}_{k}|_{U_{j}\cap U_{k}}$
whenever $U_{j}\cap U_{k}$ is nonempty. It follows that the assignment $\{f_{j}\}_{j\in J}\longmapsto\{\tilde{f}_{j}\}_{j\in J}$ induces a sheaf morphism $\mathcal{C}_{G,\omega}\to\mathcal{C}_{G,\omega^{\prime}}$. Since the map on each open set is clearly a bijection, the resulting sheaf morphism is an isomorphism.
For $G$ abelian, a base space $B$ and a local system $\mathcal{L}$ with stalk $G$, one can define cohomology groups $H^{k}(B;\mathcal{L})$, called cohomology with local coefficients in several ways, which are equivalent under mild assumptions. Most directly (though most abstractly), $H^{k}(B;\mathcal{L})$ can be interpreted as sheaf cohomology groups defined in terms of derived functors. If $B$ is paracompact, then by Proposition 2.2 we have $\check{H}^{k}(\mathcal{U};\mathcal{L})\cong H^{k}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of $B$, so we can compute these groups with Čech cohomology.
One can also define a notion of singular (co)homology with local coefficients (in fact, any (co)homology theory can be promoted to a theory with local coefficients). The details are not important for us, but we note that when $B$ is paracompact, the singular and Čech cohomology groups with coefficients in $\mathcal{L}$ are isomorphic. We also have the following generalization of Poincare duality:
Proposition 2.9.
Let $B$ be a closed connected $n$-manifold, and let $\mathcal{L}$ be a local system of abelian groups. Then $H^{k}(B;\mathcal{L})\cong H_{n-k}(B;\mathcal{L}\otimes\mathcal{O})$ for all $k$, where $\mathcal{O}$ denotes the orientation sheaf of $B$.
2.3 Classification Of Circle Bundles
In this paper, we use the term circle bundle to mean a fiber bundle with fiber $\mathbb{S}^{1}$ (note that some authors reserve this term for principal $\mathbb{S}^{1}$-bundles). A priori, the structure group for a circle bundle is $\text{Homeo}(\mathbb{S}^{1})$. We have the following two central extensions of topological groups:
$1\to\text{Homeo}_{+}(\mathbb{S}^{1})\to\text{Homeo}(\mathbb{S}^{1})\xrightarrow{\deg}\mathbb{Z}_{2}\to 1$ (2)
$1\to\mathbb{Z}\to\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})\xrightarrow{\widetilde{\exp}}\text{Homeo}_{+}(\mathbb{S}^{1})\to 1\\$ (3)
Here $\text{Homeo}_{+}(\mathbb{S}^{1})$ denotes the group of orientation-preserving homeomorphisms of $\mathbb{S}^{1}$, and $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ denotes the universal cover of $\text{Homeo}_{+}(\mathbb{S}^{1})$, which can be modeled explicitly as
$\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})=\{f\in\text{Homeo}_{+}(\mathbb{R}):\ f(x+1)=f(x)+1\ \text{ for all }x\}$ (4)
The projection map $\widetilde{\exp}:\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is defined by $\widetilde{\exp}(\tilde{f})=\text{exp}\circ\tilde{f}\circ\text{exp}^{-1}$, where $\exp:\mathbb{R}\to\mathbb{S}^{1}$ is given by $\exp(\theta)=e^{2\pi i\theta}$.
Notice that sequence (2) splits, and the section $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ induced by complex conjugation gives rise to a semidirect decomposition $\text{Homeo}(\mathbb{S}^{1})\cong\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}$. In more detail, we obtain a right $\mathbb{Z}_{2}$-action on $\text{Homeo}_{+}(\mathbb{S}^{1})$ via the assignment $\pm 1\longmapsto\text{Adj}_{r(\pm 1)}$ which induces a group multiplication law $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\text{Adj}_{r(\omega_{1})}(\Lambda_{2}),\omega_{1}\omega_{2})$. The map $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$ is then a group isomorphism.
Thus, for any open cover $\mathcal{U}$ of a paracompact space $B$, $\varphi$ induces an isomorphism on cohomology, where the cocycle condition is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
$\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$
Consequently, for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, there is a natural bijection between cocycles $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ satisfying $\deg_{*}(\Omega)=\omega$ and cochains $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$ satisfying $\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U})$. Alternatively, we can interpret $\omega$ as a local system and $\Lambda$ as a Cech cocycle in $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})$:
$\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})=\coprod_{\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\\$ (5)
The advantage of this perspective is suggested by the following proposition:
Proposition 2.10.
Suppose $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. The following sequence is short exact:
$1\to\mathbb{Z}_{\omega}\to\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}\xrightarrow{\widetilde{\exp}_{*}}\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega}\to 1$ (6)
Moreover, the sheaf $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so the induced connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ is an isomorphism.
Proof.
[…]
At the level of stalks, sequence (6) reduces to exact sequence (3), hence it is also exact as a sequence of sheaves of abelian groups. By Proposition 2.2, we therefore obtain a long exact sequence in Cech cohomology, and in particular, a connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$. Moreover, $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ is contractible, so $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so Proposition 2.2 implies $\Delta_{\omega}$ is an isomorphism by exactness.
The proposition above immediately leads to the following classification result:
Proposition 2.11.
For any open cover $\mathcal{U}$ of a paracompact space $B$, there is a bijective correspondence
$\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})\ \cong\ \coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}}_{2})}\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
via the assignment $\Omega]\longmapsto\left(w_{1}(\Omega),\tilde{e}(\Omega)\right)$ given by $w_{1}(\Omega)=[\deg_{*}(\Omega)]$ and $\tilde{e}(\Omega)=\Delta_{\omega}([\Omega])$.
We therefore have a complete classification of circle bundles over paracompact spaces in terms of a pair of discrete invariants. As we will later show, these two invariants are closely connected to the first Stiefel-Whitney class and Euler class for real plane bundles, so we will refer to $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ respectively as the Stiefel-Whitney class (or orientation class) and the twisted Euler class of the circle bundle represented by $\Omega$.
Remark 2.12.
The notation in Proposition 2.11 is slightly abusive, since each cocycle $\omega$ within a cohomology class determines a different local system $\mathbb{Z}_{\omega}$. On the other hand, by Proposition 2.8, the local systems determined by any two cohomologous cocycles are isomorphic. In particular, if $\omega$ is a coboundary, then $\mathbb{Z}_{\omega}\cong\underline{\mathbb{Z}}$.
Observe that when $B$ is a locally finite 1-dimensional CW complex (e.g., a graph), the Proposition 2.11 implies that circle bundles over $B$ are completely determined by their Stiefel-Whitney classes. In the case when $B$ is a compact connected surface, the classification is also especially simple and explicit:
Proposition 2.13.
Let $B$ be a closed connected 2-manifold, and let $\omega\in H^{1}(B;\mathbb{Z}_{2})$ determine a double cover $p_{\omega}:B_{\omega}\to B$. Then
[…]
We therefore have the following classification of circle bundles over $B$:
[…]
Proof.
Twisted Poincare duality asserts that for any compact connected $n$-manifold $M$ and locally constant sheaf $\mathcal{L}$ of Abelian groups on $M$ (with fiber $G$), one has $H^{*}(M;\mathcal{L})\cong H_{n-*}(M;\mathcal{L}\otimes\mathcal{O})$, where $\mathcal{O}$ denotes the orientation sheaf of $M$. Moreover, $H_{0}(M;\mathcal{L})$ is isomorphic to the coinvariants of the monodromy action $\omega_{\mathcal{L}}:\pi_{1}(M)\to\text{Aut}(G)$ induced by $\mathcal{L}$, i.e.,
$H_{0}(M;\mathcal{L})\cong G/\langle\omega_{\mathcal{L}}(\gamma)\cdot a-a\ |\ a\in G,\gamma\in\pi_{1}(M)\rangle$
Thus, for any compact connected surface $B$ and orientation character $\omega\in H^{1}(B;\mathbb{Z}_{2})$, one has
[…]
Since $\mathbb{Z}_{\omega}\otimes\mathcal{O}\cong\underline{\mathbb{Z}}$ if and only if $\omega+w_{1}(B)=0$ (or, equivalently, $\omega=w_{1}(B)^{-1}=w_{1}(B)$), the proposition follows.
In light of the proposition above, for any compact connected surface $B$ and $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, one can define a twisted fundamental class $B]_{\omega}$ which generates $\check{H}_{0}(B;\mathbb{Z}_{\omega})$ – this choice is unique when $\omega\neq w_{1}(B)$ and unique up to sign when $\omega=w_{1}(B)$. The pairing $\langle\tilde{e},[B]_{\omega}\rangle$ then gives an explicit bijection between $\check{H}^{2}(B;\mathbb{Z}_{\omega})$ and its image in $\mathbb{Z}$. We call the integer $\langle\tilde{e}(\Omega),[B]_{\omega}\rangle$ the twisted Euler number of the bundle classified by $\Omega$.
2.4 Reduction Of Structure Group
The structure group $\text{Homeo}(\mathbb{S}^{1})$ is rather unwieldly – in particular, it is disconnected, non-abelian and infinite-dimensional, making it challenging to work with from a computational perspective. In this section, we identify reductions of this structure group to simpler ones.
Definition 2.14.
We say that a circle bundle $\xi$ is orientable if any of the following equivalent conditions are met:
  1. 1.
    $\xi$ admits a family of local trivializations with $\text{Homeo}_{+}(\mathbb{S}^{1})$-valued transition maps,
  2. 2.
    The cohomology class $\Omega_{\xi}\in\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ lifts to a class in $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$,
  3. 3.
    The Stiefel-Whitney class $w_{1}(\Omega_{\xi})\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ vanishes.
The next proposition implies another useful characterization of orientable circle bundles:
Proposition 2.15.
The inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is a homotopy equivalence.
Proof.
Consider the map $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined by $\tau(f)=f(1)$. Observe that $\tau$ is a continuous surjection, and the fiber of $\tau$ over each $z\in\mathbb{S}^{1}$ contains the orientation-preserving homeomorphisms which map $1$ to $z$. One can easily check that each $\tau^{-1}(z)$ is contractible, so $\tau$ is a homotopy inverse for $\iota$.
An immediate consequence of the proposition above is that one can always reduce the structure group of an orientable circle bundle $\xi$ to $\mathbb{S}^{1}$ itself, which means every orientable circle bundle admits a principal $\mathbb{S}^{1}$-structure. In fact, one always has two non-isomorphic choices for the principal structure on $\xi$ corresponding to the two choices of orientation, though the underlying fiber bundles are isomorphic. This is consistent with our earlier observation that the (twisted) Euler number of a circle bundle is only well-defined up to sign until an orientation class representative for the bundle is specified.
More generally, Proposition 2.15 implies that the inclusion of $\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}=\text{Iso}(\mathbb{S}^{1})$ into $\text{Homeo}(\mathbb{S}^{1})$ is a homotopy equivalence. It follows that over paracompact spaces, any circle bundle, no matter how wild, admits a family of trivializations with $\text{Iso}(\mathbb{S}^{1})$-valued transition maps.
While the inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ induces a one-to-one correspondence between oriented circle bundles and principal $\mathbb{S}^{1}$-bundles, we note that the retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined in the proof of Proposition 2.15 is not a group homomorphism, hence does not induce a cochain map $\check{C}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})\to\check{C}^{1}(B;\mathcal{C}_{\mathbb{S}^{1}})$ which descends to an isomorphism on cohomology. In particular, the following proposition shows that no retract can induce such a map:
Proposition 2.16.
There does not exist a retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ which is also a group homomorphism.
Proof.
One can show $\text{Homeo}_{+}(\mathbb{S}^{1})$ is a perfect group in the sense that every element is expressible as a product of commutators (see [Ghy01] Proposition 5.11). On the other hand, if $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ is a group homomorphism, we must have $\tau([f,g])=[\tau(f),\tau(g)]$ for all $g,h\in\text{Homeo}_{+}(\mathbb{S}^{1})$. Since $\mathbb{S}^{1}$ is Abelian and $\text{Homeo}_{+}(\mathbb{S}^{1})$ is perfect, it follows immediately that $\tau(f)=1$ for all $f$. We conclude that $\tau$ cannot be a retract onto $\mathbb{S}^{1}$.
Thus, given an arbitrary family of trivializations for a circle bundle, it may be difficult to find explicit homotopies so that the new transitions maps take values in $\text{Iso}(\mathbb{S}^{1})$. On the other hand, if we define the supremum distance on $\text{Homeo}(\mathbb{S}^{1})$ by
$d_{\infty}(f,g)=\sup_{z\in\mathbb{S}^{1}}d_{\mathbb{S}^{1}}(f(z),g(z))\\$ (7)
where $d_{\mathbb{S}^{1}}$ denotes the geodesic distance metric on $\mathbb{S}^{1}$, the proposition below states that well-behaved homeomorphisms of $\mathbb{S}^{1}$ are close to isometries. Since the proof requires some technical digressions, we defer it to the appendix.
Proposition 2.17.
If $f\in\text{Homeo}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\leq 2\pi(\frac{L-1}{L})$.
Note in particular that $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\to 0$ as $L\to 1$.
From a computational perspective, the structure group $\text{Iso}(\mathbb{S}^{1})\cong\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}$ is vastly preferable to $\text{Homeo}(\mathbb{S}^{1})$. In particular, $\text{Iso}(\mathbb{S}^{1})$ is a 1-dimensional Lie group (with two connected components), and it has a natural identification with the matrix group $O(2)$ which gives rise to the following commutative diagram of split extensions:
[…]
Note that $\mathbb{S}^{1}$ is identified with $SO(2)$, giving rise to the following commutative diagram of exponential sequences:
$\mathbb{Z}$$\mathbb{R}$${SO(2)}{\mathbb{Z}}$$\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$$\text{Homeo}_{+}(\mathbb{S}^{1})$$\scriptstyle{\exp}$$\scriptstyle{\iota}$$\scriptstyle{\iota}$$\scriptstyle{\widetilde{\exp}}$ (9)
where $\exp:\mathbb{R}\to SO(2)$ is given by $\exp(\theta)=\left(\begin{array}[]{cc}\cos(2\pi\theta)&-\sin(2\pi\theta)\\ \sin(2\pi\theta)&\cos(2\pi\theta)\\ \end{array}\right)$. Passing to Cech cohomology, it is clear that the Stiefel-Whitney and twisted Euler classes we defined to classify circle bundles correspond precisely to their classical counterparts for vector bundles. However, it is useful to know how these classes arise and can be explicitly computed from arbitrary $\text{Homeo}(\mathbb{S}^{1})$-valued transition data.
2.5 The Universal Circle Bundle And Dimensionality Reduction
A classifying map can be used to model the total space of a fiber bundle as an embedded submanifold of a more familiar space. Specifically, if $G$ is a topological group and $\mu_{G}=(EG,p,BG)$ is a universal $G$-bundle, then for any principal $G$-bundle $\xi$ over a paracompact space $B$, we can construct a classifying map $f:B\to BG$ such that such that $f^{*}\mu_{G}=\xi$. This map is covered by a bundle map $F:E(\xi)\to EG$ which embeds the total space of $\xi$ into $EG$. More generally, if $\xi_{F}$ is the associated bundle of $\xi$ with fiber $F$, then $\xi_{F}$ is isomorphic to the pullback of the associated bundle $EG\times_{G}F\to BG$ along the same classifying map.
In the context of circle bundles, the associated bundle construction gives bijective correspondences
$\text{Bun}_{\mathbb{S}^{1}}(B)\quad\leftrightarrow\quad\text{Prin}_{O(2)}(B)\quad\leftrightarrow\quad\text{Vect}_{2,\mathbb{R}}(B)$
for any paracompact space $B$. It follows that up to homotopy, every circle bundle $\xi$ over $B$ can be realized (up to isomorphism) as the sphere bundle of a rank-2 real vector bundle $\xi_{\text{vect}}$.
A model for the classifying space of real rank-2 vector bundles is the infinite Grassmann manifold $\text{Gr}(2)=\text{Gr}_{2}(\mathbb{R}^{\infty})$ consisting of all 2-planes in $\mathbb{R}^{\infty}$, topologized as the direct limit $\text{Gr}_{2}(\mathbb{R}^{2})\subset\text{Gr}_{2}(\mathbb{R}^{3})\subset\cdots$ (see [MS74] for more details). Identifying each plane with its projection matrix gives an embedding of $\text{Gr}(2)$ into $\mathbb{R}^{\infty\times\infty}$. The total space of the universal plane bundle $\mu$ over $\text{Gr}(2)$ can be modeled by the Stiefel manifold $V(2)=V(2,\infty)$:
$V(2,\infty)=\{A\in\mathbb{R}^{\infty\times 2}:A^{T}A=I\}$
The columns of each $A\in V(2)$ form an orthonormal basis for a plane in $\text{Gr}(2)$ whose projection matrix is given by $AA^{T}$. Thus, the universal plane bundle is then defined by the surjection $\pi_{Gr}:V(2)\to\text{Gr}(2)$ given by $\pi_{Gr}(A)=AA^{T}$.
Given a circle bundle $\xi$ over a paracompact space $B$, the associated plane bundle $\xi_{\text{vect}}$ can be expressed as the pullback of $\mu$ along a classifying map $f:B\to\text{Gr}(2)$. It follows that $\xi$ is isomorphic to the pullback of the associated circle bundle $\mu_{\circ}=V(2)\times_{O(2)}\mathbb{S}^{1}\to\text{Gr}(2)$ along $f$. In this sense $\mu_{\circ}$ is a universal circle bundle, and $\xi$ is classified by the same map as $\xi_{\text{vect}}$.
From the perspective of data science and dimensionality reduction, a major advantage of choosing $\mu_{\circ}$ as our model for the universal circle bundle is that we can apply known techniques to obtain low-dimensional representations of point clouds in $V(2)\times_{O(2)}\mathbb{S}^{1}$. Our coordinatization pipeline for discrete approximate circle bundles integrates Principal Stiefel Coordinates [Lee+25] for this purpose.
3 Discrete Approximate Circle Bundles
In this section, we introduce discrete approximate notions of local trivializations, Cech cohomology and classifying maps for circle bundles. The purpose of these objects is to represent true circle bundles with finite data which can be algorithmically, efficiently and stably computed from a point cloud and a feature map.
3.1 Three Characterizations
Definition 3.1.
A discrete $(\varepsilon,\delta)$-approximate circle bundle is a map $\pi:X\to B$ between metric spaces together with an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a family of maps $\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ with the following properties:
  1. 1.
    $d_{H}(f_{l}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})<\delta$ (with respect to Euclidean distance) for every $U_{jkl}\in\mathcal{N}(\mathcal{U})$.
  2. 2.
    There exists a simplicial cochain $\Omega\in C^{1}(\mathcal{N}(\mathcal{U});O(2))$ such that
    $|f_{j}(x)-\Omega_{jk}f_{k}(x)|<\varepsilon$ (10)
    whenever $x\in\pi^{-1}(U_{jk})\in\mathcal{N}(\mathcal{U})$.
The family $\{f_{j}\}_{j\in J}$ is called a discrete $(\varepsilon,\delta)$-approximate local trivialization, and the set of discrete $(\varepsilon,\delta)$-approximate local trivializations for $\pi$ subordinate to an open cover $\mathcal{U}$ is denoted by $T_{\varepsilon,\delta}(\pi,\mathcal{U})$. Any $\Omega$ satisfying (10) is said to be a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
Given a discrete approximate local trivialization, we will refer to the quantity $\alpha\equiv\frac{\varepsilon}{1-\delta}$ as the trivialization quality of the trivialization and write $T_{\alpha}(\pi,\mathcal{U})=\bigcup_{\frac{\varepsilon}{1-\delta}<\alpha}T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for any $\alpha\geq 0$. We define the following metric on $T_{\infty}(\pi,\mathcal{U})$:
$d_{T}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})=\sup_{j\in J}\sup_{x\in\pi^{-1}(U_{j})}|f_{j}(x)-g_{j}(x)|\\$ (11)
Notice that we require witnesses to take values in $O(2)$ rather than $\text{Homeo}(\mathbb{S}^{1})$.
Next, we introduce discrete and approximate relaxations of classical Cech cohomology. As we have seen, Cech cohomology provides a useful framework for computing obstructions and classifications, so these objects will play a central role in our classification and coordinatization algorithms.
Definition 3.2.
Given a collection of open sets $\mathcal{U}=\{U_{j}\}_{j\in J}$ and a metric group $G$, the set of $G$-valued $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})=\{\Omega\in\check{C}^{k}(\mathcal{U};\mathcal{C}_{G}):\ d_{Z}(\delta^{k}\Omega,\mathbbm{1})<\varepsilon\}$ (12)
where $d_{Z}$ is the supremum metric on $\check{C}^{k}(\mathcal{U};\mathcal{C}_{G})$:
$d_{Z}(\Omega,\Lambda)=\sup_{U_{\sigma}\in\mathcal{N}(\mathcal{U})^{(k)}}\sup_{b\in U_{\sigma}}d_{G}(\Omega_{\sigma}(b),\Lambda_{\sigma}(b))$ (13)
The set of $G$-valued discrete $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\underline{G})=\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})\cap\check{C}^{k}(\mathcal{U};\underline{G})$ (14)
For $k=1$, we define the $\varepsilon$-approximate cohomology set $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})$ as the orbit space of $\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})$ with respect to the left action $\check{C}^{0}(\mathcal{U};\mathcal{C}_{G})\curvearrowright\check{C}^{1}(\mathcal{U};\mathcal{C}_{G})$ given by $(\phi\cdot\Omega)_{jk}=\phi_{j}\Omega_{jk}\phi_{k}^{-1}$.
We will primarily be interested in the case where $G=O(2)$, since every circle bundle over a paracompact space $B$ is represented by a class in $\check{H}^{1}(B;\mathcal{C}_{O(2)})$. Here we endow $O(2)$ with the Frobenius metric, since this is the most natural choice for a number of key results.
Finally, we introduce an approximate version of classifying maps. These will play an important role in our coordinatization pipeline.
Definition 3.3.
Given a paracompact space $B$ and $\varepsilon\geq 0$, an $\varepsilon$-approximate classifying map is a map $f:B\to\text{Gr}(2)^{\varepsilon}$.
3.2 Relationships Between Notions
In this section, we describe two key constructions which produce functions
$\text{wit}:T_{\alpha}(\pi,\mathcal{U})\to\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})\hskip 28.45274pt\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(2)^{\sqrt{2}\varepsilon})\\$
We first show that discrete approximate local trivializations are witnessed by discrete approximate cocycles:
Proposition 3.4.
Suppose $\Omega$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. Then $\Omega\in\check{Z}^{1}_{3\alpha\sqrt{2}}(\mathcal{U};\underline{O(2)})$.
Proof.
Assume $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$, where $\alpha=\frac{\varepsilon}{1-\delta}$. Suppose $x\in\pi^{-1}(U_{j}\cap U_{k}\cap U_{l})$. Let $v_{\tau\tau^{\prime}}=\Omega_{\tau\tau^{\prime}}f_{\tau^{\prime}}(x)-f_{\tau}(x)$ for $\tau,\tau^{\prime}\in\{j,k,l\}$. Then,
$\left|\Omega_{jk}\Omega_{kl}f_{l}(x)-\Omega_{jl}f_{l}(x)\right|=\left|\Omega_{jk}(v_{kl}+f_{k}(x))-\Omega_{jl}f_{l}(x)\right|$
$\left|\Omega_{jk}v_{kl}+(f_{j}(x)+v_{jk})-(f_{j}(x)-v_{jl})\right|=\left|\Omega_{jk}v_{kl}+v_{jk}-v_{jl}\right|$
$\leq|v_{kl}|+|v_{jk}|+|v_{jl}|<3\varepsilon\\$
The result then follows immediately from Proposition A.10.
Next, we show how to algorithmically compute a witness for a discrete approximate circle bundle.
Construction 3.5.
Let $X=\{x_{s}\}_{s=1}^{n}$ be a finite set, and let $f,g:S\to\mathbb{S}^{1}$. Compute the values $\{\alpha_{s}=\cos^{-1}(f(x_{s}))\}^{n}_{s=1}$ and $\{\beta_{s}=\cos^{-1}(g(x_{s}))\}^{n}_{s=1}$, then define $\gamma_{s}=\alpha_{s}-\beta_{s}$ and $\tilde{\gamma}_{s}=\alpha_{s}+\beta_{s}$ for all $s$. Find the respective midpoints $\theta$ and $\widetilde{\theta}$ of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ and the shortest arc $\widetilde{C}$ containing $\{\widetilde{\gamma}_{s}\}_{s=1}^{n}$. Compute the values
$d_{\text{rot}}=\text{max}_{s}|f(x_{s})-R_{\theta}g(x_{s})|,\hskip 28.45274ptd_{\text{ref}}=\text{max}_{s}|f(x_{s})-R_{\widetilde{\theta}}g(x_{s})|$
If $d_{\text{rot}}<d_{\text{ref}}$, set $\Omega=R_{\theta}$; otherwise, set $\Omega=R_{\widetilde{\theta}}r$, where $R_{\theta}=\left(\begin{array}[]{cc}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\\ \end{array}\right)$ and $r=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$.
Proposition 3.6.
If $n>1$ and $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ in Construction 3.5 above, then $\Omega\in O(2)$ is the unique minimizer of $\text{max}_{s}|f(x_{s})-\Omega g(x_{s})|$.
Proof.
Let $L:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L(\Omega)=\max_{s}|f(x_{s})-\Omega g(x_{s})|^{2}$. We have
$|f(x_{s})-\Omega g(x_{s})|^{2}$ $=|f(x_{s})|^{2}+|g(x_{2})|^{2}-2f(x_{s})\cdot\Omega g(x_{s})$
$=2-2f(x_{s})\cdot\Omega g(x_{s})$
for all $s$. If $\Omega=R_{\theta}$ is a counter-clockwise rotation by $\theta$, then $f(x_{s})\cdot\Omega g(x_{s})=\cos(\gamma_{s}-\theta)$, so to minimize $L$ over $SO(2)$ we need to identify $\theta\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\gamma_{s}$ is minimized. If the measure of $C$ is less than $\pi$, then this is precisely achieved at the midpoint of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ (furthermore, the requirement that $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ guarantees that $C$ is unique). To see this, first note that the optimal $\theta$ certainly lies in $C$, and for fixed $\theta\in C$, $\cos(\gamma_{s}-\theta)$ is smallest when $\gamma_{s}$ is an endpoint of $C$. By symmetry, the $\theta$ which maximizes the minimum of the values $\{\cos(\gamma_{s}-\theta)\}_{s=1}^{n}$ must be half way between the two end points (otherwise, a small perturbation of $\theta$ towards the midpoint would move $\theta$ closer to the more distant endpoint while still keeping $\theta$ nearer to the closer endpoint. This would yield a more optimal choice for $\theta$).
Analogously, if $\Omega=R_{\widetilde{\theta}}r$ for some $\widetilde{\theta}\in\mathbb{R}$, one can check that $f(x_{s})\cdot\Omega g(x_{s})=\cos(\widetilde{\gamma}_{s}-\theta)$. Thus, to minimize $L$ over the component of $O(2)$ with matrices of negative determinant, we need to identify $\widetilde{\theta}\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\widetilde{\gamma}_{s}$ is minimized. Finally, comparing $d_{\text{rot}}$ and $d_{\text{ref}}$ determines which component of $O(2)$ contains the optimal $\Omega$. To see that these two values cannot be equal, observe that if $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})>\pi$. It follows that then $d_{\text{rot}}<\frac{\pi}{2}$ and $d_{\text{ref}}>\frac{\pi}{2}$. Similarly, if $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})>\pi$, so $d_{\text{ref}}<d_{\text{rot}}$.
Corollary 3.7.
If $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for $\varepsilon<\sqrt{2}$ and $|\pi^{-1}(U_{jk})|>1$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, then Construction 3.5 yields a witness.
Proof.
The hypotheses imply
$|f_{j}(x)-\Omega_{jk}f_{k}(x)|^{2}<\varepsilon^{2}$
or, equivalently,
$\gamma^{jk}_{s}=(f_{j}(x_{s}))\cdot(\Omega_{jk}f_{k}(x_{s}))>1-\frac{\varepsilon^{2}}{2}$
for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$. Then $|\gamma^{jk}_{s}|>\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$ for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $\text{diam}(\{\gamma^{jk}_{s}\})<2\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$. We conclude that if $\cos^{-1}(1-\frac{\varepsilon^{2}}{2})<\frac{\pi}{2}$ (or, equivalently, if $\varepsilon<\sqrt{2}$), then each $\Omega_{jk}$ obtained using Construction 3.5 is the unique minimizer of $\text{max}_{x_{s}\in\pi^{-1}(U_{j}\cap U_{k})}|f_{j}(x_{s})-f_{k}(x_{s})|$. Thus by Proposition 3.4, $\Omega$ is a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
We state the following technical result for later reference:
Proposition 3.8.
Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\}\in T_{\varepsilon_{f},\infty}(\pi,\mathcal{U})$ and $\{g_{j}\}\in T_{\varepsilon_{g},\infty}(\pi,\mathcal{U})$, and $d_{h}(f_{k}(\pi^{-1}(U_{jk})),\mathbb{S}^{1})<\delta_{1}$ (with respect to the Euclidean metric) for every ordered 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$. If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta_{2}$, then $d_{Z}(\Omega,\Lambda)<\frac{\sqrt{2}}{1-\delta_{1}}(\varepsilon_{f}+\varepsilon_{g}+2\delta_{2})$.
Proof.
Given any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, the triangle inequality implies
$|\Omega_{jk}f_{k}(x)-\Lambda_{jk}f_{k}(x)|\leq$
$|\Omega_{jk}f_{k}(x)-f_{j}(x)|\ +\ |f_{j}(x)-g_{j}(x)|\ +\ |g_{j}(x)-\Lambda_{jk}g_{k}(x)|\ +\ |\Lambda_{jk}g_{k}(x)-\Lambda_{jk}f_{k}(x)|$
$\varepsilon_{f}+\delta_{2}+\varepsilon_{g}+\delta_{2}\\$
The result then follows immediately from Proposition A.10.
Corollary 3.9.
If $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ both witness $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ then $d_{Z}(\Omega,\Lambda)<(2\sqrt{2})\alpha$.
Corollary 3.10.
Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}(\alpha+\delta)$.
The following construction and subsequent result are due to [SP23]:
Construction 3.11.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some locally-finite open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of a paracompact space $B$. Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$. For each $j$, define $\Phi_{j}:U_{j}\to V(2)$ by
$\Phi_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho_{0}(b)}\Omega_{0j}(b)\\ \sqrt{\rho_{1}(b)}\Omega_{1j}(b)\\ \vdots\\ \end{array}\right)$ (15)
Note that when $b\in U_{j}\cap U_{k}^{c}$, $\rho_{k}(b)=0$, so this map is well-defined. Let $f:B\to\mathbb{R}^{\infty\times\infty}$ be defined by
$\widetilde{f}(b)=\sum_{j\in J}\rho_{j}(b)\Phi_{j}(b)\Phi_{j}(b)^{T}$ (16)
Since each $b$ is contained in a finite number of sets in $\mathcal{U}$, this map is well-defined.
For a proof of the proposition below, see [SP23]:
Theorem 3.12.
For $\varepsilon<\frac{1}{2}$ and any partition of unity $\{\rho_{j}\}_{j\in J}$, Construction 3.11 yields a map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(d)^{\sqrt{2}\varepsilon})$ with the following properties:
  1. 1.
    The induced map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to[B,\text{Gr}(d)^{\sqrt{2}\varepsilon}]$ factors through $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ and does not depend on choice of $\{\rho_{j}\}_{j\in J}$.
  2. 2.
    If $\Omega,\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Lambda)<\delta$, then $\text{cl}(\Omega)$ and $\text{cl}(\Lambda)$ become equal in $B,\text{Gr}(d)^{\sqrt{2}(\delta+\varepsilon)}]$.
3.3 Relationship To True Circle Bundles
In this section, we show that for sufficiently small $\alpha$, any discrete approximate local trivialization $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ corresponds to a true circle bundle. We build on previous results of [Tin22] and [SP23] which provide conditions under which approximate $O(2)$-cocycles and classifying maps can be uniquely identified with true ones.
The following construction and subsequent result, from [Tin22], describe a procedure for obtaining a classifying map from an approximate one:
Construction 3.13.
Given $A\in\mathbb{R}^{n\times n}$, compute the symmetrization $A_{s}=\frac{A+A^{T}}{2}$, then compute an orthogonal diagonalization $\Lambda D\Lambda^{T}$ of $A_{s}$ such that the eigenvalues appear in decreasing order along the diagonal of $D$. Finally, set $\Pi(A)=\Lambda J_{2}\Lambda^{T}$, where $J_{2}$ is the $n\times n$ matrix whose only non-zero entries are the first two diagonal entries, both equal to 1. Note that this construction naturally extends to the case $n=\infty$.
Proposition 3.14.
Let $\Pi:\mathbb{R}^{\infty\times\infty}\to\text{Gr}(2)$ be the function described in Construction 3.13. For $\varepsilon\leq\frac{\sqrt{2}}{2}$, the restriction of $\Pi$ to $\text{Gr}(2)^{\varepsilon}$ is a retraction map. Furthermore, the induced map $\Pi_{*}:\text{Maps}(B,\text{Gr}(2)^{\varepsilon})\to\text{Maps}(B,\text{Gr}(2))$ descends to a bijection $\Pi_{*}:[B,\text{Gr}(2)^{\varepsilon}]\to[B,\text{Gr}(2)]$.
For later use, we also record the following additional construction and result from the same paper:
Construction 3.15.
Let $S_{n}=\{(P,A)\in\text{Gr}(2,n)\times V(2,n):\text{rank}(PA)=2\}$. Given any $(P,A)\in S_{n}$, let $\widehat{\Pi}_{n}(P,A)\in V(2,n)$ denote the $U$ matrix from the polar decomposition $UH=PA$. Explicitly, $H=((PA)^{T}(PA))^{\frac{1}{2}}$ and $U=(PA)H^{-1}$.
Proposition 3.16.
Construction 3.15 defines a map $\widehat{\Pi}_{n}:S_{n}\to V(2,n)$ for each $n$, thereby inducing a map $\widehat{\Pi}:S\to V(2)$ (where $S$ is the direct limit of the sequence $S_{1}\subset S_{2}\subset\cdots$). Furthermore, $\widehat{\Pi}(P,A)$ is the closest matrix to $A$ in $\pi_{Gr}^{-1}(P)$ with respect to the Frobenius metric.
Finally, we have the following construction and subsequent result from [SP23]:
Construction 3.17.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$. Let $\{\Phi_{j}:U_{j}\to V(2)\}_{j\in J}$ be the maps defined by (15), and let $\Pi$ be as described in Construction 3.13. Define $f_{\Omega}=\Pi_{*}\circ\text{cl}(\Omega)$, then for each $j$, define $\widetilde{\Phi}_{j}:U_{j}\to V(2)$ by $\widetilde{\Phi}_{j}(b)=\hat{\Pi}(f_{\Omega}(b),\Phi_{j}(b))$. Finally, define $\widetilde{\Omega}\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ by $\widetilde{\Omega}_{jk}(b)=\widetilde{\Phi}_{j}^{T}(b)\widetilde{\Phi}_{k}(b)$.
Proposition 3.18.
For $\varepsilon<\frac{1}{2}$, Construction 3.17 yields a map $\Pi_{Z}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that the following diagram commutes:
[…]
We also have an induced commutative diagram at the level of cohomology and homotopy classes of maps. Furthermore, if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for $\varepsilon<\frac{\sqrt{2}}{4}$, then $d_{Z}(\Omega,\Pi_{Z}(\Omega))<9\varepsilon$.
The following corollary is immediate:
Corollary 3.19.
If $\varepsilon<\frac{\sqrt{2}}{2}$, any $\varepsilon$-approximate classifying map uniquely determines a true classifying map, hence can be identified with an isomorphism class of true circle bundles. Similarly, if $\varepsilon<\frac{1}{2}$, any $\varepsilon$-approximate $O(2)$ Cech 1-cocycle $\Omega$ uniquely determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
Proposition 3.20.
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$. Then any witness determines a true circle bundle, and the circle bundles determined by any two witnesses coincide. In this sense, $\{f_{j}\}$ uniquely determines a true circle bundle.
Proof.
Any witness $\Omega$ lies in $\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ by Proposition 3.4. Thus, if $\alpha<\frac{1}{6\sqrt{2}}$, then $(3\sqrt{2})\alpha<\frac{1}{2}$, so by Proposition 3.19 $\Omega$ determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
By Corollary 3.9, any two witnesses $\Omega,\Omega^{\prime}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ satisfy $d_{Z}(\Omega,\Omega^{\prime})<2\varepsilon\sqrt{2}$. Thus, by Proposition 3.12, $\Pi_{*}\circ\text{cl}(\Omega)=\Pi_{*}\circ\text{cl}(\Omega^{\prime})$ as long as $\sqrt{2}(3\varepsilon\sqrt{2}+2\varepsilon\sqrt{2})=10\varepsilon\leq\frac{\sqrt{2}}{2}$, or equivalently, $\varepsilon\leq\frac{\sqrt{2}}{20}$.
We now discuss conditions under which a discrete approximate circle bundle can be uniquely and stably identified with a class of true circle bundles.
Proposition 3.21.
Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$. If $5\alpha+2\delta<\frac{\sqrt{2}}{4}$, then $\{f_{j}\}$ and $\{g_{j}\}$ both determine the same true circle bundle.
Proof.
By Proposition 3.10, any two witnesses $\Omega_{1},\Omega_{2}\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ that $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ satisfy $d_{Z}(\Omega_{1},\Omega_{2})<2\sqrt{2}(\alpha+\delta)$. By Proposition 3.4, we have $\Omega_{1},\Omega_{2}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, so Proposition 3.18 implies $\text{cl}(\Omega_{1})]$ and $\text{cl}(\Omega_{2})]$ coincide in $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$, where $\varepsilon=(3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)$. Since $\Pi$ induces a bijection from $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$ to $B,\text{Gr}(2)]$ by Proposition 3.19, $\Omega_{1}$ and $\Omega_{2}$ determine the same true circle bundle as long as
$\sqrt{2}\left((3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)\right)<\frac{\sqrt{2}}{2}$
or equivalently, if $5\alpha+2\delta<\frac{\sqrt{2}}{4}$.
Corollary 3.22.
Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ are homotopic through $T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$. Then $\{f_{j}\}$ and $\{g_{j}\}$ determine the same true circle bundle.
Proof.
Suppose $\{h_{j}:\pi^{-1}(U_{j})\times I\to\mathbb{S}^{1}\}$ are continuous maps such that $\{h_{j}(\cdot,t)\}\in T_{\alpha}(\pi,\mathcal{U})$ for all $t$, and $h_{j}(x,0)=f_{j}(x)$ and $h_{j}(x,1)=g_{j}(x)$. Proposition 3.21 implies that for any $\alpha<\frac{\sqrt{2}}{20}$, there exists some $\delta>0$ such that $\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\}\in T_{\alpha}(\pi,\mathcal{U})$ determine the same true bundle as long as $d_{T}(\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\})<\delta$. By continuity, it follows that $\{h_{j}(\cdot,t)\}$ determines the same true bundle for all $t\in[0,1]$, and in particular that the true bundles determined by $\{f_{j}\}$ and $\{g_{j}\}$ coincide.
Next, we provide a construction for producing a family $\{\widetilde{f}_{j}\}\in T_{0}(\pi,\mathcal{U})$ from a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ when $\alpha$ is sufficiently small.
Construction 3.23.
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. Let $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ be any witness, and let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$, and for each $j$, define $\widetilde{f}_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}$ by
$\widetilde{f}_{j}(x)=\exp\left(\sum_{k\in J}\rho_{k}(\pi(x))\log\left(\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\right)\right)$ (17)
Note that if $\pi(x)\notin U_{j}\cap U_{k}$ then $\rho_{k}(\pi(x))=0$, so the maps are well-defined.
Proposition 3.24.
Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Then the maps $\{\widetilde{f}_{j}\}$ described in Construction 3.23 satisfy $\widetilde{f}_{j}(x)=\widetilde{\Omega}_{jk}(\pi(x))\widetilde{f}_{k}(x)$ for all $x\in\pi^{-1}(U_{j}\cap U_{k})$ and $U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$.
Proof.
Let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$, and assume $\Omega$ decomposes as $\Omega_{jk}(b)=\exp(\Theta_{jk}(b))r(\omega_{jk})$ for some $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$, where $\exp:\mathbb{R}\to SO(2)$ is as defined in (9) and $r:\mathbb{Z}_{2}\to O(2)$ is given by $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Then, for any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
$\tilde{\Omega}_{jk}(b)\tilde{f}_{k}(x)\quad$ $=\quad\tilde{\Omega}_{jk}(b)\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{kl}f_{l}(x)\right)\right)$
$=\quad\exp(\Theta_{jk}(b))r(\omega_{jk})\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
$=\quad\exp\left(\Theta_{jk}(b)+\omega_{jk}\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
$=\quad\exp\left(\sum_{l}\rho_{l}(b)\left(\Theta_{jk}(b)+\omega_{jk}\log\left(\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)\right)$
$=\quad\exp\left(\sum_{l}\rho_{l}(\pi(x))\log\left(\tilde{\Omega}_{jk}(b)\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
$=\quad\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{jl}(b)f_{l}(x)\right)\right)$
$=\quad\tilde{f}_{j}(x)$
Proposition 3.25.
Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{Z}(\Omega,\widetilde{\Omega})<\delta$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. If $\delta+\alpha\sqrt{2}<2$, then for any partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and $x\in U_{j}$ for some $j\in J$, $\widetilde{f}_{j}(x)$ as defined by (17) is equal to the unique weighted Karcher mean of $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ with weights $\{\rho_{k}(\pi(x))\}_{k\in K_{x}}$, where $K_{x}=\{k\in J:x\in U_{j}\cap U_{k}\}$. Furthermore $d_{T}(\{f_{j}\},\{\widetilde{f}_{j}\})<\alpha+\frac{\delta}{\sqrt{2}}$, and the maps $\{\widetilde{f}_{j}\}_{j\in J}$ obtained by different choices of partition of unity are homotopic through $T_{0}(\pi,\mathcal{U})$.
Proof.
For any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
$|f_{j}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$ $\leq|f_{j}(x)-\Omega_{jk}f_{k}(x)|+|\Omega_{jk}f_{k}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$
$<\alpha+\frac{|\Omega_{jk}-\widetilde{\Omega}_{jk}(\pi(x))|_{F}}{\sqrt{2}}$
$\leq\alpha+\frac{\delta}{\sqrt{2}}$
Thus, by Proposition A.11, $\widetilde{f}_{j}(x)$ as defined in (17) is the unique weighted Karcher mean of the set $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ as long as $\alpha+\frac{\delta}{\sqrt{2}}<\sqrt{2}$. Clearly we have $d_{T}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})\leq\alpha+\frac{\delta}{\sqrt{2}}$ in this case.
Now, suppose $\{\rho^{\prime}_{j}\}_{j\in J}$ is another partition of unity subordinate to $\mathcal{U}$. Each partition of unity can be viewed as the component functions of a map $\rho,\rho^{\prime}:B\to\Delta^{n-1}$. Since $\Delta^{n-1}$ is convex, we can form the straight line homotopy $h_{t}:B\to\Delta^{n-1}$ from $\rho$ to $\rho^{\prime}$. Replacing the components of $\rho$ with those of $h_{t}$ in (17) for each $t\in[0,1]$ gives a continuously varying family in $T_{0}(\pi,\mathcal{U})$, so $\{f_{j}\}$ and $\{\widetilde{f}_{j}\}$ are homotopic through $T_{0}(\pi,\mathcal{U})$.
Corollary 3.26.
For $\alpha<\frac{\sqrt{2}}{28}$, Construction 3.23 defines a map $\Pi_{T}:T_{\alpha}(\pi,\mathcal{U})\to T_{0}(\pi,\mathcal{U})$ such that for any $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$, the families $\{f_{j}\}$ and $\Pi_{T}(\{f_{j}\})$ determine the same true bundle.
Proof.
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{28}<\frac{\sqrt{2}}{20}$. By Proposition 3.4, any witness satisfies $\Omega\in\check{Z}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, and Proposition 3.18 implies $d_{Z}(\Omega,\widetilde{\Omega})<(27\sqrt{2})\alpha$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. The result then follows from Proposition 3.25.
4 Effective Computation Of Characteristic Classes
In this section, we give explicit algorithms for computing the characteristic classes of a circle bundle from a discrete approximate cocycle representative. One can then use Construction 3.5 to compute classes from discrete approximate trivializations. We also introduce a notion of persistence for these classes.
We first note some simplifications which occur in the computation of the twisted Euler class when the transition functions are $O(2)$-valued (rather than $\text{Homeo}(\mathbb{S}^{1})$-valued). Let $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ denote the section of (2) described in Section 2.3, and let $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ be the group isomorphism described there given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$. Relative to the reduced structure group $O(2)$, we have $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Moreover, Proposition A.9 implies $\text{Adj}_{r(\omega)}\Lambda=\Lambda^{r(\omega)}$ for any $\Lambda\in O(2)$ and $\omega=\pm 1\in\mathbb{Z}_{2}$, so the group multiplication law in $\mathbb{S}^{1}\rtimes_{r}\mathbb{Z}_{2}$ simplifies to
$(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\Lambda_{2}^{\omega_{1}},\omega_{1}\omega_{2})$ (18)
Thus, the cocycle condition for $(\mathbb{S}^{1}\rtimes\mathbb{Z}_{2})$-valued cocycles is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
$\Lambda_{jk}\Lambda_{kl}^{\omega_{jk}}=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$ (19)
It follows that in this context, the connecting homomorphism $\Delta_{\omega}$ described in Proposition 2.10 reduces to
$\Delta_{\omega}([\Omega])_{jkl}=\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}$ (20)
where $\Omega=(\varphi^{-1})_{*}(\Lambda,\Theta)$. For later use, we define $\text{abs}:O(2)\to SO(2)$ by $\text{abs}(\Omega)=\Omega r(\det(\Omega))$ – note $\varphi^{-1}(\Omega)=(\text{abs}(\Omega),\det(\Omega))$. In fact, it is easy to see that if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\geq 0$ and $\omega=\det_{*}(\Omega)$, then $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$.
Below we give an explicit algorithm for computing characteristic classes from a discrete approximate representative. To prove stability, we first prove a generalization of a technical result from [SP23]. Let $1\to F\to G\xrightarrow{\varphi}H\to 1$ be a central extension of Lie groups with H compact and F discrete. Fix a bi-invariant Riemannian metric on $H$ and use it to metrize $H$ and $G$ with the geodesic distance and $F$ with the distance inherited from $G$. Suppose $F$, $G$ and $H$ have left $\Gamma$-actions by isometries for some abelian group $\Gamma$, and that the maps in the central extension are $\Gamma$-equivariant. Let $\mathcal{U}$ be an open cover of a topological space $B$ such that each set and non-empty binary intersection is locally path connected and simply connected. Let $\text{sys}(H)$ denote the systole of $H$.
Proposition 4.1.
For $\varepsilon\leq\frac{\text{sys}(H)}{8}$ and any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$, we have a well-defined map $\Delta_{\omega}:\check{Z}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ given by $\Delta_{\omega}(\Lambda)=[\langle\delta^{1}_{\omega}\Theta\rangle]$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ is any lift of $\Lambda$ respect to $\varphi_{*}$. Furthermore, $\Delta_{\omega}$ descends to a map $\Delta_{\omega}:\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ which is stable in the sense that if $\Lambda],[\Lambda^{\prime}]\in\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$ satisfy $d_{H}([\Lambda],[\Lambda^{\prime}])<\frac{\text{sys}(H)}{8}$, then $\Delta_{\omega}([\Lambda])=\Delta_{\omega}([\Lambda^{\prime}])$.
Remark 4.2.
If $G$ is a metric group and $\Gamma$ is an abelian group which acts on $G$ by isometries, the $d_{Z}$ metric on $\check{C}^{1}(\mathcal{U};\mathcal{C}_{G})$ naturally extends to a metric on $\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$. We then define $\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G,\omega})$ and $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G,\omega})$ for all $\varepsilon\geq 0$ in the obvious way.
Proof.
The proof is essentially the same as that of Theorem 6.6 in [SP23]. The fundamental difference is that we are doing cohomology with local coefficients, so the coboundary maps are slightly different. On the other hand, because $\Gamma$ acts by isometries, the same arguments apply. In particular, for any $\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$, one can use the same steps shown in [SP23] to verify the 2-cochain given by $\beta_{jkl}=\langle(\delta_{\omega}^{1}\Theta)_{jkl}\rangle$ is a genuine cocycle, and the stability arguments are identical.
An immediate result of the proposition above is the following:
Corollary 4.3.
For $\varepsilon\leq 1$, we have a well-defined map $\text{Cl}:\check{H}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{O(2)})\to\coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{H}^{2}(B;\mathbb{Z}_{\omega})$ given by $\Omega\longmapsto([\det_{*}\Omega],[\Delta_{\omega}(\text{abs}_{*}\Omega)])$. This map is 1-stable in the sense that $\text{Cl}(\Omega)=\text{Cl}(\Omega^{\prime})$ whenever $d_{H}(\Omega,\Omega^{\prime})<1$. Moreover, if $\varepsilon<\frac{1}{9}$, then $\text{Cl}(\Omega)$ returns the first Stiefel-Whitney class and twisted Euler class of the true circle bundle represented by $\Pi_{Z}(\Omega)$.
Proof.
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\leq 1$. Lemma 1.3 of [SP23] implies $\omega=\det_{*}(\Omega)$ is a cocycle and that $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any other $\Omega^{\prime}\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Omega^{\prime})<2$. If $\varepsilon<1/9$, then by Proposition 3.18, there exists a true cocycle $\Omega^{\prime}$ such that $d_{Z}(\Omega,\Omega^{\prime})<1$, so $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any such $\Omega^{\prime}$. We conclude that the first Stiefel-Whitney class of any such $\Omega^{\prime}$ is the same.
Now, recall central extension (9). Let $SO(2)_{g}$ denote the Lie group $SO(2)$ endowed with the geodesic distance metric (via the natural identification of $SO(2)$ with $\mathbb{S}^{1}$). The induced metrics on $\mathbb{R}$ and $\mathbb{Z}$ are then the usual ones, and all three groups have $\mathbb{Z}_{2}$-actions by isometries such that the inclusion and $\exp$ maps are equivariant. Thus, Proposition 4.1 gives a well-defined map $\Delta_{\omega}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ for all $\varepsilon\leq\frac{sys(SO(2)_{g})}{8}$, where $\omega=\det_{*}(\Omega)$. By Lemma A.18 from [SP23], $\frac{sys(SO(2)_{g})}{8}=\frac{2\pi\sqrt{2}}{8}=\frac{\pi\sqrt{2}}{4}$ with respect to geodesic distance. To obtain a result in terms of the Frobenius norm, we note Corollary A.17 from [SP23] implies we have an inclusion $\check{Z}_{1}^{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})\subset\check{Z}^{1}_{\frac{\pi\sqrt{2}}{4}}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})$. Thus, $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$, so $\Delta_{\omega}(\text{abs}_{*}(\Omega))\in\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ (and therefore $\text{Cl}(\Omega)$) is well-defined.
[…]
Remark 4.4.
Because the isomorphism class of a circle bundle is completely determined by its characteristic classes, Proposition 4.1 implies that if $\varepsilon<\frac{1}{9}$, then any true cocycle $\widetilde{\Omega}$ such that $d_{Z}(\Omega,\widetilde{\Omega})<1$ must satisfy $\text{Cl}(\widetilde{\Omega})=\text{Cl}(\Omega)$.
Below is an explicit algorithm for computing characteristic class representatives.
Input: A cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ subordinate to a finite open cover $\mathcal{U}$.
Output: A pair of cochains $\omega\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}\in\check{C}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
  1. 1.
    Compute $\omega_{jk}=\det(\Omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
  2. 2.
    For each $(jk)\in\mathcal{N}(\mathcal{U})$, compute $\Lambda_{jk}=\Omega_{jk}r(\omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, where $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Then, choose a lift $\Theta_{jk}\in\mathbb{R}$ of each $\Lambda_{jk}$ via the $\exp$ map defined in (9).
  3. 3.
    Compute $\tilde{e}_{jkl}=\langle\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}\rangle$ for each $(jkl)\in\mathcal{N}(\mathcal{U})$.
Corollary 4.5.
If $\mathcal{U}$ is a finite good open cover and $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\underline{O(2)})$ for some $\varepsilon<\frac{1}{2}$, the outputs $\omega$ and $\tilde{e}$ of the algorithm above are cocycles. If $\varepsilon<\frac{1}{9}$, then $\omega$ and $\tilde{e}$ represent the characteristic classes of $\Pi_{Z}(\Omega)$.
In the case where $B$ is a closed connected 2-manifold, the following describes how to explicitly compute the twisted Euler number associated with a twisted Euler class:
Input: A finite good open cover $\mathcal{U}$ of a closed connected surface $B$ and cocycles $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, $\tilde{e}\in\check{Z}^{1}(\mathcal{U};\mathbb{Z}_{\omega})$.
Output: The twisted Euler number $\tilde{\eta}$ associated to $\tilde{e}$ (up to choice of sign).
  1. 1.
    For $p=2,3$, construct the matrix representation $D_{\omega,p}$ for the twisted boundary operator $\partial_{\omega,p}:\check{C}_{p+1}(\mathcal{U};\mathbb{Z}_{\omega})\to\check{C}_{p}(\mathcal{U};\mathbb{Z}_{\omega})$ with respect to some total orderings on the simplices in each dimension (e.g., lexicographical). Explicitly, the non-zero entries of $D_{\omega,p}$ are given by
    $(D_{\omega,1})_{jkl}^{kl}=\omega_{jk},\quad(D_{\omega,1})_{jkl}^{jl}=-1,\quad(D_{\omega,1})_{jkl}^{jk}=1\\$ (21)
    for each $(jkl)\in\mathcal{N}(\mathcal{U})$ and
    $(D_{\omega,2})_{jklm}^{klm}=\omega_{jk},\quad(D_{\omega,2})_{jklm}^{jlm}=-1,\quad(D_{\omega,2})_{jklm}^{jkm}=1,\quad(D_{\omega,2})_{jklm}^{jkl}=-1$
    for each $(jklm)\in\mathcal{N}(\mathcal{U})$.
  2. 2.
    Compute the Smith normal form $LD_{\omega,2}R=S$ of $D_{\omega,2}$ over $\mathbb{Z}$ and let $k$ denote the number of zero columns in $S$). If $k=1$, set $\mu_{B,\omega}$ denote the last column of $R$ and proceed to Step 4. Otherwise, let $K$ be the submatrix of $R$ containing the last $k$ columns, and let $B$ equal the last $k$ rows of the product $R^{-1}D_{\omega,3}$.
  3. 3.
    Compute the Smith normal form $L^{\prime}BR^{\prime}=S^{\prime}$ of $B$. Let $v$ denote the last column of $R^{\prime}$, and set $\mu_{B,\omega}=Kv$.
  4. 4.
    Compute $\tilde{\eta}=\sum_{(jkl)\in\mathcal{N}(\mathcal{U})}\tilde{e}_{jkl}\cdot(\mu_{B,\omega})_{jkl}$.
Remark 4.6.
Although Čech theory is typically formulated in terms of cohomology, there is a natural dual notion of Čech homology with the property that $\check{H}_{*}(\mathcal{U};\mathcal{L})\cong H_{*}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of a space $B$ and any local system $\mathcal{L}$ of abelian groups (in particular, given a good open cover $\mathcal{U}$, we can define $\check{C}_{k}(\mathcal{U};\mathcal{L})=\text{Hom}_{\mathbb{Z}}(\check{C}^{k}(\mathcal{U};\mathcal{L}),\mathbb{Z})$ and consider the associated homology groups – these groups then coincide with the corresponding singular homology). We can therefore interpret $D_{\omega}$ in the construction above as the matrix representation of a twisted boundary operator whose kernel is generated by the associated twisted fundamental class.
Proposition 4.7.
The algorithm above computes the twisted Euler number associated to a twisted Euler class (up to choice of sign).
Proof.
Since $B$ is a compact connected surface, we have $\text{rank}(H_{2}(B;\mathbb{Z}_{\omega}))=1$, and any generator coincides with the fundamental class up to sign. If $\mathcal{U}$ has no quadruple intersections, then $H_{2}(B;\mathbb{Z}_{\omega})=Z_{2}(B;\mathbb{Z}_{\omega})=\text{ker}\partial_{\omega,2}$, so the last column $r$ of the matrix $R$ above represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$ (specifically, the vector $r$ corresponds to the simplicial cycle $\sum_{U_{jkl}\in\mathcal{N}(\mathcal{U})^{(2)}}r_{jkl}\cdot U_{jkl}$).
If $k>1$, then the columns of $K$ form a basis for $\ker D_{\omega,2}$. Moreover, $\text{col}(D_{\omega,3})\subseteq\text{col}(K)\subseteq\text{col}(R)$, so only the last $k$ rows of $R^{-1}D_{\omega,3}$ are non-zero. The matrix $B$ expresses the columns of $D_{3,\omega}$ as linear combinations of the columns of $K$, so the kernel of $B$ consists of vectors which represent non-trivial 2-cycles. Since $\text{rank}(H_{2}(B;\mathbb{Z}_{\omega}))=1$ by assumption, we conclude $\ker B$ is spanned by the last column $v$ of $R^{\prime}$ and $\mu_{B,\omega}=Kv$ represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$.
Finally, pairing $\mu_{B}$ with $\widetilde{e}$ gives the twisted Euler number associated with $\widetilde{e}$ up to sign. If desired, one can calibrate the fundamental class representative $\mu_{B,\omega}$ obtained from the algorithm above by imposing a sign choice for the pairing of $\mu_{B,\omega}$ with a fixed Euler class representative.
5 Persistence Of Characteristic Classes
Discrete approximate trivializations constructed from a real-world dataset are unlikely to satisfy the fairly restrictive criteria required for our theoretical guarantees to hold. Due to practical considerations such as non-uniform sampling density or the presence of outliers for which the feature map is ill-defined, we may have reason to believe that some of the local measurements incorporated into our model are more reliable than others. In other cases, we may want to make a deliberate choice to ignore or modify certain parts of the data so that the restricted discrete approximate cocycle representative becomes trivial. Given a family of discrete approximate local trivializations $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and a witness $\Omega$, we therefore introduce a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ which measures the alignment quality of each pair $f_{j}|_{\pi^{-1}(U_{j}\cap U_{k})}$ and $f_{k}|_{\pi^{-1}(U_{j}\cap U_{k})}$ by the corresponding $\Omega_{jk}$. We then construct an induced simplex-wise filtration $\mathcal{W}$ on $\mathcal{N}(\mathcal{U})$ and compute the persistence of $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ with respect to $\mathcal{W}$. In Section 6, we show how to construct a coordinatization map which is in some sense compatible with the restriction of $\Omega$ to any subcomplex $W^{r}$ in $\mathcal{W}$.
5.1 Weights Filtration
Definition 5.1.
Suppose $\pi:X\to B$ is a map between metric spaces, $\mathcal{U}$ is an open cover of $B$ and $\Omega\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. For each $U_{jk}\in\mathcal{N}(\mathcal{U})^{(1)}$, define
$w(U_{jk})=\frac{1}{|\pi^{-1}(U_{jk})|}\sum_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-\Omega_{jk}(\pi(x))f_{k}(x)|$ (22)
Extend $w$ to a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ by setting $w(U_{j})=0$ for all $j$ and inductively defining $w$ for higher simplices by $w(U_{\sigma})=\max_{\sigma^{\prime}\subset\sigma}w(U_{\sigma^{\prime}})$. Order the simplices of $\mathcal{N}(\mathcal{U})$ by (increasing) weight, then by (increasing) dimension, then lexicographically. This ordering induces a simplex-wise filtration $\mathcal{W}^{\Omega}$ of $\mathcal{N}(\mathcal{U})$, called the weights filtration of $\mathcal{N}(\mathcal{U})$ associated with $\Omega$, since all 0-simplices appear first, and any higher simplex appears later than each of its faces.
Remark 5.2.
Though we chose (22) as a measurement of the alignment quality of $f_{j}|_{\pi^{-1}(U_{jk})}$ and $f_{k}|_{\pi^{-1}(U_{jk})}$ via $\Omega_{jk}$, there are plenty of other reasonable choices one could make depending on context. For instance, one could use max error rather than average error. To ensure the weight function induces a total ordering of $\mathcal{N}(\mathcal{U})$, we assume $w$ is injective on 1-simplices (one can always obtain this setup by applying arbitrarily small perturbations).
Definition 5.3.
Suppose $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$. Let $\mathcal{W}^{\Omega}=\{W_{r}\}_{r=1}^{|\mathcal{N}(\mathcal{U})|}$ denote the associated weights filtration. Given a cochain $\Lambda\in\check{C}^{p}(\mathcal{U};\mathcal{F})$ for some presheaf of groups $\mathcal{F}$ and $p\geq 1$, we define the cobirth of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$, where $r$ is the largest index such that $\Lambda|_{W_{r}}$ is a cocycle. Similarly, we define the codeath of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$ for $r$ the largest index such that $\Lambda|_{W_{r}}$ is a coboundary.
(a) Full nerve
(b) $\omega$ cobirth
(c) $\omega$ codeath
Figure 2: Various stages of the weights filtration on the nerve of an open cover. Each edge is labeled with its associated weight and the value assigned by a $\mathbb{Z}_{2}$-simplicial cochain $\omega$.
Construction 5.4.
Suppose $G$ is a PID and $G_{\omega}$ is a local system with stalk $G$. If $\mathcal{U}$ is finite, one can efficiently compute the cobirth and codeath of $\Lambda$ as follows: let $D$ denote the matrix representation of the coboundary operator $\delta_{\omega}:\check{C}^{*}(\mathcal{U};G_{\omega})\to\check{C}^{*}(\mathcal{U};G_{\omega})$ when the simplices are arranged in increasing order according to the weights filtration. Define $\hat{\Lambda}\in\mathbb{R}^{|\mathcal{N}(\mathcal{U})|}$ such that $\hat{\Lambda}_{r}=\Lambda(U_{\sigma_{r}})$ whenever $\sigma_{r}$ is a $p$-simplex and $\hat{\Lambda}_{r}=0$ otherwise. It is easy to see that the cobirth of $\Lambda$ is 1 less than the smallest index $r_{b}$ such that $(D\hat{\Lambda})_{b_{p}}$ is nonzero (if none exists, the cobirth is $|\mathcal{N}(\mathcal{U})|$).
To determine the codeath, write $S=LDR$, where $L$ and $R$ are invertible and $S$ is the Smith normal form of $D$. For $p>0$, let $r_{p}$ denote the largest index of a row of $S$ corresponding to a $p$-simplex and containing a non-zero entry. The codeath of $\Lambda$ is 1 less than the largest index $d_{b}>r_{p}$ such that $(L\hat{\Lambda})_{d_{b}}$ is nonzero (if none exists, then the codeath equals the cobirth).
Given a cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$, we can use Construction 5.4 to compute the cobirth index $b_{sw}$ and codeath index $d_{sw}$ of $\omega=\det_{*}(\Omega)\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ (here the local system is trivial). Then, for any $r\geq b_{sw}$, we can meaningfully ask whether $\tilde{e}(\Omega)|_{W^{r}}=\Delta_{\omega}(\Omega)|_{W^{r}}\in C^{1}(W^{r};\mathbb{Z}_{\omega})$ is a cocycle, so we compute the persistence of $\tilde{e}(\Omega)$ restricted to $W^{b_{sw}}$.
5.2 Interpretation
We would like to interpret each filtration stage $W^{r}\in\mathcal{W}$ as the nerve of a modified open cover $\mathcal{U}^{r}$ of a new base space $B^{r}$. For example, removing an edge $(jk)$ from $\mathcal{N}(\mathcal{U})$ effectively separates sets $U_{j}$ and $U_{k}$. For the purposes of coordinatization, there are various choices one could make about how to handle data in $\pi^{-1}(U_{j}\cap U_{k})$, and the appropriate choice depends on context. One option would be to simply remove these points from the dataset – this would be a reasonable choice if the intersection is small and consists mostly of outliers with ill-defined projections. In other cases, one may want to keep the data while modifying $B$ and $\mathcal{U}$. For instance, if the cobirth of a class representative is $\infty$ (or all the weights are fairly small and close together), one may still want to introduce cuts in the base space in order to artificially create a trivial bundle which admits a global coordinate system. With this in mind, we offer a recursive approach for constructing a family $\{B^{r}\}$ of spaces and corresponding open covers $\{\mathcal{U}^{r}\}$ such that $W^{r}=\mathcal{N}(\mathcal{U}^{r})$.
Construction 5.5.
Let $\mathcal{W}^{\Omega}$ denote the weights filtration associated with some $\Omega\in C^{k}(\mathcal{N}(\mathcal{U});\underline{O(2)})$. Suppose $W^{r}\in\mathcal{W}^{\Omega}$ for $r>1$, and we have a space $B^{r}$ and open cover $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j\in J}$ such that $B_{r}=\bigcup_{j\in J}U^{r}_{j}$ and $W^{r}=\mathcal{N}(\mathcal{U}^{r})$. Let $\sigma_{r}\in\mathcal{N}(\mathcal{U})$ denote the simplex which appears in $W^{r}$ but not $W^{r-1}$. Choose some $j\in\sigma_{r}$ (for instance, the last $j$ lexicographically) and define $U^{r-1}_{j}=U_{j}^{r}\cap(\overline{U^{r}_{\sigma_{r}}})^{c}$ and $B^{r-1}=B^{r}-(U^{r}_{j}\cap\partial U^{r}_{\sigma_{r}})$ (for all $k\neq j$, define $U^{r-1}_{k}=U^{r}_{k}$). It is easy to check that $U^{r-1}_{j}\subset B^{r-1}$ is open, $B^{r-1}=\bigcup_{j}U^{r-1}_{j}$ and $W^{r-1}=\mathcal{N}(\mathcal{U}^{r-1})$. In terms of the original space $B$, we have
$B^{r}=B-\bigcup_{s>r}(U_{j_{s}}^{s}\cap\partial U_{\sigma_{s}^{s}})$
6 Coordinatization
Suppose $\mathcal{U}=\{U_{j}\}_{j=1}^{n}$ is a finite open cover of a metric space $B$ and $\Omega\in\check{Z}_{\varepsilon}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{\varphi_{j}\}_{j=1}^{n}\in T_{\alpha}(\pi,\mathcal{U})$ for some map $\pi:X\to B$ ($X$ a finite metric space). Let $\mathcal{W}=\{W^{r}\}$ denote the associated weights filtration of $\mathcal{N}(\mathcal{U})$. For each $r$, we show how to construct a coordinatization map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ for $d\leq 2n$ which is in some sense compatible with $\Omega|_{W^{r}}$. This gives a menu of options for models for the dataset. In the special case when $\Omega|_{W^{r}}$ is an approximate coboundary, we construct a ’global trivialization’ map $F^{r}:X\to B\times\mathbb{S}^{1}$.
  1. 1.
    Step 1: (Choose A Filtration Stage) Choose a filtration stage $W^{r}\in\mathcal{W}$ of $\mathcal{N}(\mathcal{U})$. Use Construction 5.5 to construct the sets $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j=1}^{n}$, then choose a partition of unity $\{\rho^{r}_{j}\}_{j=1}^{n}$ subordinate to $\mathcal{U}^{r}$.
  2. 2.
    Step 2: (Dimensionality Reduction) For each $j$, define $\widehat{\Phi}^{r}_{j}:U^{r}_{j}\to V(2,2n)$ by
    $\widehat{\Phi}^{r}_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho^{r}_{0}(b)}\Omega_{0j}\\ \vdots\\ \sqrt{\rho^{r}_{n}(b)}\Omega_{nj}\\ \end{array}\right)$ (23)
    Choose a value $d\leq 2n$ and apply Principal Stiefel Coordinates to the point cloud $X_{V}=\bigcup_{j}\Phi_{j}^{r}(U_{j}^{r})\subset V(2,2n)$ to obtain a map $\text{psc}:X_{V}\to V(2,d)$.
  3. 3.
    Step 3: (Classifying Map) For each $j$, let $\widetilde{\Phi}_{j}^{r}=\text{psc}\circ\widehat{\Phi}_{j}^{r}:U_{j}^{r}\to V(2,d)$. Define $\tilde{f}^{r}:B^{r}\to\mathbb{R}^{d\times d}$ by
    $\widetilde{f}^{r}(b)=\sum_{j=1}^{n}\rho^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)^{T}$ (24)
    and let $f^{r}=\Pi\circ\tilde{f}^{r}:B^{r}\to Gr(2,d)$.
  4. 4.
    Step 4: (Cocycle Projection) For each $j$, define $\Phi^{r}_{j}:U^{r}_{j}\to V(2,d)$ by $\Phi^{r}_{j}(b)=\widehat{\Pi}(f^{r}(b),\widetilde{\Phi}^{r}_{j}(b))$, where $\widehat{\Pi}$ is the map from Construction 3.15. Then, define $\Omega^{r}\in\check{Z}^{1}(\mathcal{U}^{r};\mathcal{C}_{O(2)})$ by $\Omega^{r}_{jk}(b)=(\Phi^{r}_{j}(b))^{T}\Phi^{r}_{k}(b)$.
  5. 5.
    Step 5: (Bundle Map) Define $F^{r}_{j}:\pi^{-1}(U^{r}_{j})\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ by
    $F_{j}^{r}(x)=\Phi^{r}_{j}(\pi(x))\exp\left(\sum_{k=0}^{n}\rho_{k}^{r}(\pi(x))\log\Omega_{jk}^{r}(\pi(x))\varphi_{k}(x)\right)$ (25)
    The $F^{r}_{j}$’s agree on overlaps, yielding a globally defined map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$.
6.1 Coordinatization Pipeline For Trivial Bundles
In the special case when a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U}^{r})$ is witnessed by a discrete approximate coboundary $\Omega$, we construct a global trivialization map $\varphi:X\to B\times\mathbb{S}^{1}$. In particular, suppose $\omega=\text{det}_{*}(\Omega)\in\check{Z}^{1}(\mathcal{U}^{r};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}=\Delta_{\omega}(\Omega)\in\check{Z}^{2}(\mathcal{U}^{r};\mathbb{Z}_{\omega})$ are coboundaries.
  1. 1.
    Construct a solution to the linear system $\delta^{0}\phi=\omega$ over $\mathbb{Z}_{2}$. Define $\hat{\Omega}=\phi\cdot\Omega\in\check{Z}^{1}(\mathcal{U}^{r};\underline{SO(2)})$, and let $\hat{f}_{j}=\phi_{j}\circ f_{j}$ for all $j$.
  2. 2.
    Define $e_{jkl}=\phi_{j}\tilde{e}_{jkl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U}^{r})^{(2)}$ and construct a solution to the linear system $\delta^{1}\beta=e$ over $\mathbb{Z}$.
  3. 3.
    Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and any lift $\Theta\in\check{C}^{1}(\mathcal{U}^{r};\underline{\mathbb{R}})$ of $\hat{\Omega}$ via $\text{exp}_{*}$. Define $\mu\in\check{C}^{1}(\mathcal{U}^{r};\mathcal{C}_{\mathbb{R}})$ by
    $\mu_{j}(b)=\sum_{k=1}^{n}\rho_{k}(b)(\Theta_{kj}-\beta_{kj})$ (26)
    and let $\bar{f}_{j}=\exp(\mu_{j})\circ\hat{f}_{j}$ for all $j$.
  4. 4.
    Define $F:X\to B\times\mathbb{S}^{1}$ by
    $F(x)=\left(\pi(x),\ \exp\left(\sum_{j=1}^{n}\rho_{j}(\pi(x))\log\tilde{f}_{j}(x)\right)\right)$ (27)
Remark 6.1.
Since trivial bundles are flat, we can often use Singer’s method [WS13] in practice to obtain a locally-constant true coboundary $\widetilde{\Omega}$ which is close to an approximate one. Though we have no theoretical guarantees on stability, the fact that the result is locally-constant guarantees that the resulting trivialization map does not have unnecessary variation.
7 Implementations
Here we demonstrate the practical viability and utility of the algorithms we have described in this paper through a series of implementations on real and synthetic datasets. In each case, we used the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations. A summary of the results of each experiment can be found in Table 2. Note that the witness quality $\alpha$ in each case was well outside the range required to meet our theoretical hypotheses, which suggests that our approach is stable for a larger class of objects than the current results apply to. Lastly, further details about our implementation on optical flow data can be found in an upcoming companion paper.
Experiment Base Space Ambient Dim Sample Size SW1 $\widetilde{\eta}$
Optical Flow Patches $\mathbb{RP}^{1}$ 18 25,000 0 0
Prism Densities $\mathbb{RP}^{2}$ 15,408 25,000 $\neq 0$ $\pm 3$
Star Densities $\mathbb{S}^{2}$ 15,408 25,000 0 $\pm 10$
Table 1: A summary of the setups for the three experiments.
Experiment $\varepsilon_{\text{trivs}}$ $\delta$ $\alpha$ $\varepsilon_{\text{cocycle}}$ $d_{Z}(\Omega,\Pi_{Z}(\Omega))$
Optical Flow Patches 0.976 0 0.976 0 0
Prism Densities 0.582 0.272 0.799 0.250 0.248
Star Densities 1.120 0.620 2.949 0.501 0.471
Table 2: Summary of experimental results.
Experiment $(\mathcal{U})$ $w_{\text{max}}$ $b_{\omega}$ $d_{\omega}$ $b_{\widetilde{e}}$ $d_{\widetilde{e}}$ $(W^{r})^{(1)}$ $(W^{r})^{(2)}$
Optical Flow Patches 12 0.976 0.976 0.976 0.976 0.976 12/12 0/0
Prism Densities 73 0.987 0.987 0.231 0.987 0.231 70/216 4/73
Star Densities 146 1.939 1.939 1.939 1.939 1.912 431/432 286/288
Table 3: Summary of persistence information for the computed characteristic class representatives in each experiment.
7.1 Optical Flow Patches
Optical flow refers to the apparent motion of objects between consecutive frames of a video [HS81]. This perceived motion is characterized by an assignment of a vector $v\in\mathbb{R}^{2}$ to each pixel in each frame. The vector at pixel $(x,y)$ in frame $t$ points to the location in frame $t+1$ where pixel $(x,y)$ appears to move. The estimation of optical flow from a video is central to many computer vision tasks, including motion tracking, object segmentation, and video compression [MG15], [Zha+21]. The Sintel dataset [But+12], derived from the open-source animated short film “Sintel” by the Blender Foundation, was created as a high-quality benchmark for testing and training optical flow models.
Figure 3: A frame from the Sintel video with sample (scaled) optical flow vectors shown
In [Ada+20], Adams et al. studied the nonlinear statistics of 3 x 3 high-contrast optical flow patches from [But+12]. They proposed that a dense core subset $X$ of (normalized) high-contrast patches is concentrated around a 2-manifold in $\mathbb{R}^{18}$ homeomorphic to $\mathbb{T}^{2}$. To support their claim, they introduced a feature map $\pi_{\text{flow}}:X\to\mathbb{RP}^{1}$ which roughly measures a patch’s predominant axis of flow (see Figure 4). They then offered evidence that the data in $X$ over any small interval in $\mathbb{RP}^{1}$ has circular topology, clearly supporting the hypothesis that an underlying circle bundle structure is present.
Figure 4: A sample of mean-centered and normalized high-contrast optical flow patches labeled by predominant direction (represented as an angle in $0,\pi)$)
By Proposition 2.13, there are exactly two non-isomorphic circle bundles over $\mathbb{RP}^{1}\cong\mathbb{S}^{1}$ which can be distinguished computationally by their Stiefel-Whitney classes in $H^{1}(\mathbb{RP}^{1};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$. The total spaces of the two bundles are the torus and the Klein bottle. To confirm the torus model proposed by [Ada+20], we used the same pre-processing to obtain a sample of 50,000 normalized high-contrast optical flow patches from the Sintel dataset and refine to a sample $X_{\text{flow}}$ of 25,000 patches from the dense core subset $X(300,50)$. We then computed the predominant axis of flow $\pi_{\text{flow}}(x)\in\mathbb{RP}^{1}$ for each patch $x\in X_{\text{flow}}$. Figure 5 shows the persistence diagram for a random subsample of 1000 points from $X_{\text{flow}}$. For comparison, we also computed persistence diagrams for a random 1000-point sampling from $X(300,30)$ and a noisy synthetic sampling of the proposed torus model. Observe that the synthetic dataset exhibits the Betti signature of $\mathbb{T}^{2}$ ($\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$) over a large range of scales, while the most persistent Betti signature for the samples of $X(300,50)$ and $X(300,30)$ is $\beta_{0}=1$, $\beta_{1}=1$, $\beta_{2}=0$.
Figure 5:
On the other hand, as observed in [Ada+20], the local picture of $X_{\text{flow}}$ with respect to $\pi_{\text{flow}}$ tells a different story. We covered $\mathbb{RP}^{1}$ with 12 open intervals of the form $U_{j}=(\frac{j\pi}{12}-\frac{\pi}{16},\frac{j\pi}{12}+\frac{\pi}{16})$ (note that $\frac{\pi}{16}$ is the maximal radius for which the nerve of $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ has no triple intersections). Figure 6 shows the first two principal components and the persistence diagrams for the data in each $\pi_{\text{flow}}^{-1}(U_{j})$, strongly suggesting the presence of an underlying circle bundle structure.
Figure 6: PCA and persistence diagrams for the sets $\{\pi^{-1}(U_{j})\}_{j=1}^{12}$
We computed a family $\{f_{j}\}_{j=1}^{12}$ of discrete approximate local trivializations for $X_{\text{flow}}$ subordinate to $\mathcal{U}$ using the sparse circular coordinates algorithm, then obtained a witness $\Omega\in\check{C}^{1}(\mathcal{N}(\mathcal{U});O(2))$ using the procedure outlined in Construction 3.5. Figure 7 shows a visualization of $\mathcal{N}(\mathcal{U})$; the weight $w(\Omega_{jk})$ of each edge $(jk)$ is shown in black, and the values assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)$ are shown in red. A 0-cochain $\phi\in C^{0}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ satisfying $\delta^{0}\phi=\omega$ is shown in blue, confirming a topological torus structure underlying the data.
Figure 7: (Left) Correlations on intersections between the discrete approximate local trivializations $\{f_{j}\}_{j=1}^{12}$ computed for $X_{\text{flow}}$. (Right) The nerve of the open cover $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ of $\mathbb{RP}^{1}$. Each edge is labeled by its associated weight. The components of the Stiefel-Whitney class representative $\omega=\text{det}_{*}(\Omega)$ are shown in blue, and a potential $\phi$ for $\omega$ is shown in red.
Once we verified the global triviality of the underlying bundle structure, we used the coordinatization pipeline described in Section 6 to obtain a global coordinate system for $X_{\text{flow}}$. To our surprise, we discovered that a significant portion of the patches had ambiguous predominant flow directions despite being high-contrast (see Figure 8). Such patches are not accounted for in the torus model proposed by [Ada+20]. Upon further investigation, we found that the fibers of the predominant direction map $\pi$ are geometrically more like cylinders than circles, and that this extra degree of freedom locally parametrizes variations in directional strength among patches in $X_{\text{flow}}$. We present the details of our unexpected findings and a refined model for the dataset in an upcoming paper. The generalized model we describe also accounts for other families of high-contrast patches not featured in the original torus model which are particularly relevant for motion tracking analysis.
Figure 8: A sample of high-contrast optical flow patches for which the predominant direction is essentially arbitrary
Figure 9 shows a selection of coordinatized patches from $X_{\text{flow}}$ which are highly directional (see the upcoming paper for our formal definition of directional strength). Notice that patches near the top and bottom of the diagram roughly coincide, reflecting the circular topology of the fibers. Patches near the left and right edges also roughly coincide, which reflects the global torus structure.
Figure 9: (Left) The optical flow torus model proposed by [Ada+20]. (Right) A sample of coordinatized patches from $X_{\text{flow}}$ with high directional strength, arranged by predominant flow direction (x) and assigned fiber coordinate (y).
From a modeling perspective, the results of our analysis of $X_{\text{flow}}$ demonstrate how the interplay between local and global measurements depends crucially on the geometry of the chosen feature map. On the other hand, a careful fiberwise analysis helped us to both confirm an existing model and identify subtle, informative additional structure in the data which had previously been missed. See our upcoming paper for additional details.
7.2 3D Densities
Many applications of computer vision and imaging involve data in the form of density or intensity functions on some region of space or time. The ambient dimension for such data is inevitably very large, and distances computed using the standard Euclidean metric often fail to reflect one’s intuition about how densities should compare (for example, two densities which differ by a small translation may be very far apart with respect to Euclidean distance). One would ideally opt for an Earth-mover-type distance such as the Wasserstein metric [RTG00], but these metrics are prohibitively expensive to compute, so an effective local-to-global approach to topological modeling which circumvents the need to construct a full distance matrix is highly desirable.
Consider the space $\mathcal{D}$ of real-valued functions on the unit ball $\mathbb{B}^{3}$. The space $\mathcal{D}$ has a natural right $SO(3)$-action defined by $(\rho\cdot A)(x)=\rho(A^{T}x)$ for all $\rho\in\mathcal{D}$, $A\in SO(3)$ and $x\in\mathbb{B}^{3}$. In particular, the orbit of any fixed $\rho_{0}\in\mathcal{D}$ is diffeomorphic to a quotient of $SO(3)$ by the corresponding stabilizer subgroup. For our demonstration, we sampled the $SO(3)$-orbit of a density $\rho_{\text{prism}}\in\mathcal{D}$ concentrated around the boundary of a triangular prism $P\subset\mathbb{B}^{3}$ (see Figure 10). Explicitly, $\rho_{\text{prism}}:\mathbb{B}^{3}\to\mathbb{R}$ is given by
$\rho_{\text{prism}}(p)=\exp\left(-\frac{d(p,\partial P)^{2}}{2\sigma^{2}}\right)$ (28)
where $d(p,\partial P)$ denotes the (Euclidean) distance from $p$ to the boundary of $P$, and $\sigma=0.05$. Note that $\rho_{\text{prism}}$ has a 3-fold rotational symmetry in the $yz$-plane as well as a 2-fold rotational symmetry along the $x$-axis. The orbit $\mathcal{O}_{\text{prism}}$ is therefore homeomorphic to the quotient $L(6,1)/\mathbb{Z}_{2}$, where $L(6,1)$ denotes the lens space $\mathbb{S}^{3}/\mathbb{Z}_{6}\cong SO(3)/\mathbb{Z}_{3}$.
Figure 10: A sample of discrete 3D prism densities from $X_{\text{prism}}$. Each density is represented by its 2D projection along the $x$-axis as well as the corresponding rotation of the prism $P$ used to generate $\rho_{\text{prism}}$. Note that any two rotated prisms which are related by an element of $\text{stab}(\rho_{\text{prism}})$ result in indistinguishable 3D density functions.
Consider the feature map $\pi:\mathcal{O}_{\text{prism}}\to\mathbb{RP}^{2}$ defined as follows: given any $\rho\in\mathcal{O}_{\text{prism}}$, let $M_{\rho}$ be the covariance matrix given by
$M_{\rho}=\int_{\mathbb{B}^{3}}\rho(x)(x-\mu_{\rho})(x-\mu_{\rho})^{T}dx$
where $\mu_{\rho}=\int_{\mathbb{B}^{3}}x\ \rho(x)dx$ is the center-of-mass. Then, let $\pi(\rho)\in\mathbb{RP}^{2}$ be the axis corresponding to the smallest eigenvalue of $M_{\rho}$. One can interpret $\pi(\rho)$ as the axis along which $\rho$ has the smallest variance. In particular, for any $\rho\in\mathcal{O}_{\text{prism}}$, $\pi(\rho)$ is precisely the axis of symmetry (explicitly, $\pi(\rho_{\text{prism}}\cdot A)=A^{T}e_{1}$). In Section A.2, we show $(\mathcal{O}_{\text{prism}},\pi,\mathbb{RP}^{2})$ is a non-orientable circle bundle with twisted Euler number $\pm 3$.
We generated a synthetic dataset by applying 25,000 random rotations to $\rho_{\text{prism}}$ and sampling the resulting densities on a lattice $L$ (in particular, we restricted a 32 x 32 x 32 lattice on $1,1]^{3}$ to $\mathbb{B}^{3}$). The sampled density values were then flattened to vectors $X_{\text{prism}}\subset\mathbb{R}^{32^{3}}$. Figure 11 shows the PCA cumulative explained variance $X_{\text{prism}}$, clearly illustrating that the embedding of the underlying 3-manifold is highly non-linear. Figure 12 shows a PCA visualization of a ’fat fiber’ of the dataset over a small neighborhood in $\mathbb{RP}^{2}$.
Figure 11:
Figure 12: A PCA visualization of the data in a ’fat fiber’ of $X_{\text{prism}}$ (right) over a small neighborhood in $\mathbb{RP}^{2}$ (left), portrayed here by a hemisphere with boundary identifications.
One can estimate the covariance matrix $M_{\rho}$ of a density $\rho$ from its discretization $v_{\rho}$ as
$\widetilde{M}_{\rho}=\sum_{l\in L}v_{\rho}(l)(l-\widetilde{\mu}_{\rho})(l-\widetilde{\mu}_{\rho})^{T}$
where $\tilde{\mu}_{\rho}=\sum_{l\in L}l\ v_{\rho}(l)$. Setting $\pi(v_{\rho})\in\mathbb{RP}^{2}$ equal to the axis associated with the minimum eigenvalue of $\widetilde{M}_{\rho}$ yields a well-defined map $\pi:X_{\text{prism}}\to\mathbb{RP}^{2}$.
To classify and coordinatize the resulting discrete approximate circle bundle, we constructed a cover $\mathcal{U}=\{U_{j}\}_{j=1}^{73}$ of $\mathbb{RP}^{2}$ and used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations, then used Construction 3.5 to obtain a witness $\Omega\in C^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ and compute characteristic class representatives. As suggested by Figure 12, we found that the standard Euclidean metric was sufficient to model the small portion of $X_{\text{prism}}$ in each $\pi^{-1}(U_{j})$.
Figure 13(a) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$. The node representing each $U_{j}\in\mathcal{U}$ is labeled by the rotated prism used to generate a sample $x_{j}\in X_{\text{prism}}$ whose base projection $\pi(x_{j})\in\mathbb{RP}^{2}$ was near the center of $U_{j}$. Note in particular that the prisms at nodes which are connected by an edge have similar axes of symmetry. The edge representing each 1-simplex $U_{jk}\in\mathcal{N}(\mathcal{U})$ is colored according to the value assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$, red for 1-1). We verified that $\omega$ was not a coboundary, confirming that one cannot consistently choose an orientation for the data (since $H^{1}(\mathbb{RP}^{2};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$, verifying non-triviality is sufficient to uniquely determine the orientation class). With respect to $\omega$, the twisted Euler number of $\Omega$ was found to be 3-3, in agreement with our theoretical model.
Using the algorithm described in Section 5, we computed the cobirth and codeath of each characteristic class representative (see Table 3). Each representative was found to be a cocycle on the entire nerve of $\mathcal{U}$, but the two cocycles only became coboundaries after 146 of the 216 1-simplices in $\mathcal{N}(\mathcal{U})$ were removed. One can interpret this result as a reflection of the compatibly non-orientable structures of the bundle and the base space. More specifically, the restriction of the underlying bundle $(\mathcal{O}_{\text{prism}},\pi,\mathbb{RP}^{2})$ to any non-nullhomotopic loop (e.g. an equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$) is non-orientable, hence isomorphic to the Klein bottle. Thus, the restriction of an orientation class representative $\omega\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ to any subcomplex of $\mathcal{N}(\mathcal{U})$ which includes a non-nullhomotopic loop must still be non-trivial.
(a) Full Nerve
(b) Maximal Subcomplex
Figure 13: The 1-skeleton of the nerve $\mathcal{N}(\mathcal{U})$ of the open cover of $\mathbb{RP}^{2}$ used to analyze $X_{\text{prism}}$. Edges are colored according to the values of the Stiefel-Whitney class representative (blue for 1, red for -1). The figure on the right shows the 1-skeleton of the maximal subcomplex of $\mathcal{N}(\mathcal{U})$ on which the characteristic class representatives restricted to coboundaries; the nodes which correspond to open sets containing the base projections of points in $\widetilde{X}_{\text{prism}}$ are highlighted.
To illustrate, we used the coordinatization algorithm described in Section 6 to compute global coordinates for $X_{\text{prism}}$ compatible with the maximal stage $W\subset\mathcal{N}(\mathcal{U})$ of the weights filtration on which the characteristic class representatives became coboundaries. We isolated the set of densities $\widetilde{X}_{\text{prism}}\subset X_{\text{prism}}$ whose base point projections satisfied $|z|<0.1$, then projected these base points to the equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$. Figure 13(b) shows the 1-skeleton of $W$; the nodes which correspond to open sets containing the base projections of points in $\widetilde{X}_{\text{prism}}$ are highlighted. Figure 14 shows a sample of $\widetilde{X}_{\text{prism}}$ arranged according to assigned base angles in $\mathbb{RP}^{1}$ and assigned fiber angles in $\mathbb{S}^{1}$. Observe that the base angle assigned to each density roughly captures its axis of symmetry, and that this axis steadily traverses a rotation by $\pi$ as the base angle varies from 0 to $\pi$. The fiber angles assigned to densities in any narrow vertical strip roughly parametrize a rotation by $\frac{2\pi}{3}$ about a common axis of symmetry. On the other hand, careful inspection shows that the direction of rotation abruptly flips near the dotted vertical line, reflecting a location in the nerve where a cut was made before coordinatization. This also confirms the Klein bottle structure of the restricted bundle underlying $\widetilde{X}_{\text{prism}}$.
(a) 2D Density Projections
(b) Underlying Prism Meshes
Figure 14: A collection of coordinatized samples from $\widetilde{X}_{\text{prism}}$ arranged according to base projection angle (x) and assigned fiber coordinate (y). The dotted vertical line indicates an orientation flip in the coordinate system.
Finally, we used the pipeline described in Section 6 to construct a point cloud embedding $F:X_{\text{prism}}\to V(2)$ which respects the global topology implied by the characteristic classes (recall these were well-defined on the entire nerve $\mathcal{N}(\mathcal{U})$). In this sense, our pipeline serves a similar function as Isomap, but with the preservation of topological (rather than geometric) features in mind. The mean projection errors for different projection dimensions are shown in Figure 15 .
Figure 15:
7.3 Geometric Signal Unwrapping
Our final implementation provides an example of how the techniques and algorithms developed here can be integrated with existing topological modeling tools to form a larger local-to-global TDA framework. We consider a second dataset of discretized densities, sampled from the $SO(3)$-orbit of the density $\rho_{\text{star}}\in\mathcal{D}$ shown in Figure 16. Note that $\rho_{\text{star}}$ has a single 5-fold rotational symmetry in the $yz$-plane, so $\mathcal{O}_{\text{star}}$ is homeomorphic to the lens space $SO(3)/\mathbb{Z}_{5}\cong L(10,1)$. Applying the same projection map as we used for $\mathcal{O}_{\text{prism}}$ yields a bundle $(\mathcal{O}_{\text{star}},\pi,\mathbb{RP}^{2})$ with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$.
Figure 16: A sample of 3D densities from $X_{\text{star}}$. Each density is represented by its 2D projection along the $x$-axis as well as the corresponding rotated triangle mesh used to generate the density.
Figure 17: A PCA visualization of the data in a ’fat fiber’ of $X_{\text{star}}$ (right) over a small neighborhood in $\mathbb{RP}^{2}$ (left), portrayed here by a hemisphere with boundary identifications.
[…]
As in the previous experiment, we sampled 25,000 densities from $\mathcal{O}_{\text{star}}$ and formed a dataset $X_{\text{star}}\subset\mathbb{R}^{32^{3}}$ from the discretizations. We used a Mapper-like pipeline to cluster the data $X_{j}=X_{j}^{+}\coprod X_{j}^{-}$ over each open set $U_{j}$ in the same cover of $\mathbb{RP}^{2}$ (the sign labels for each pair of local clusters $X_{j}^{\pm}$ were assigned arbitrarily). Figure 18(a) shows a graph $G$ whose nodes are the local clusters $\{X_{j}^{\pm}\}_{j=1}^{73}$ and whose edges represent intersections between clusters. Edges between local clusters with the same sign label are shown in blue, and intersections between local clusters with opposite sign labels are shown in red. Figure 18(b) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$ with edges colored according to the values assigned by the induced connectivity cocycle $\nu\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$ and red for 1-1). Note that each edge in Figure 18(b) corresponds to a pair of edges in Figure 18(a) of the same color. Figure 19 gives a more detailed visualization of $G$; each local cluster $X_{j}^{\pm}$ is labeled by the rotated star pyramid used to generate a sample density it contains. For readability, edges connecting nodes with opposite sign labels are not shown explicitly, but their presence is indicated by light gray lines. Lastly, observe that $G$ can be canonically identified with the 1-skeleton of the nerve of a good open cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$.
We verified that $\nu$ was a non-trivial cocycle, confirming that $X_{\text{star}}$ could not be globally separated into two connected components and that $\pi|_{X_{\text{star}}}$ could instead be lifted to a map $\tilde{\pi}:X_{\text{star}}\to\mathbb{S}^{2}$. To explicitly construct $\tilde{\pi}$, we selected a random $v_{\rho}\in X_{1}^{+}$ and chose an arbitrary lift $\tilde{\pi}(v_{\rho})\in\mathbb{S}^{2}$ of $\pi(v_{\rho})\in\mathbb{RP}^{2}$. For every other $v_{\rho^{\prime}}\in X_{1}^{+}$, we then defined $\tilde{\pi}(v_{\rho^{\prime}})$ to be the lift‘ of $\pi(v_{\rho^{\prime}})$ whose inner product with $\tilde{\pi}(v_{\rho})$ is positive. Similarly, for each $v_{\rho^{\prime}}\in X_{1}^{-}$, we defined $\tilde{\pi}(v_{\rho^{\prime}})$ so that $\tilde{\pi}(v_{\rho^{\prime}})\cdot\tilde{\pi}(v_{\rho})<0$. We then propagated this choice of lift across the entire dataset using a breadth-first search of $G$. Finally, we used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations for $X_{\text{star}}$ over $\mathbb{S}^{2}$, then used the algorithms described in Section 4 to obtain representatives of the orientation and twisted Euler classes. The results are summarized in Table 2. Note the Stiefel-Whitney class representative we obtained was trivial, and we obtained a twisted Euler number of $10$, as expected.
(a) Local Cluster Graph $G$
(b) Connectivity Cocycle $\nu$
Figure 18: Visualizations of (a) the graph $G$ which represents the connectivity relations among local clusters in $X_{\text{star}}$. (b) The 1-skeleton of the nerve $\mathcal{N}(\mathcal{U})$ of the open cover of $\mathbb{RP}^{2}$ used to analyze $X_{\text{star}}$ (right). Each 1-simplex of $\mathcal{N}(\mathcal{U})$ is colored according to the assigned value of the connectivity class representative $\nu$ (blue for 1, red for -1), and the pair of corresponding edges in $G$ are assigned this same color.
Figure 19: A detailed visualization of the local cluster graph $G$ which double covers the 1-skeleton of $\mathcal{N}(\mathcal{U})$. Nodes corresponding to local clusters labeled with a ++ are shown on the left, and the corresponding clusters labeled by a - are shown on the right. Edges which connect local clusters with the same sign label are shown in blue. For readability, edges which connect clusters with opposite sign labels are not shown explicitly, but are indicated with gray lines. Note the diagram above can equivalently be interpreted as representing the nerve of the cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$ which we ultimately used to model $X_{\text{star}}$ as a discrete approximate circle bundle.
The process of lifting a bundle with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$ to a genuine circle bundle over a suitable covering space can be viewed as a form of geometric signal unwrapping. In other fields, classical signal unwrapping techniques (e.g., in interferometry, MRI, or phase retrieval) are used to resolve ambiguities in angular measurements (see, for example, [GP98]); our construction provides a topological analog in which the discrete fiber ambiguity is resolved via a covering, yielding a continuous, globally well-defined structure.
8 Conclusion and Future Work
We have introduced a discrete, approximate analog of circle bundles, motivated by data-theoretic applications. We established conditions under which such an object can be identified with a true circle bundle and described algorithms for recovering the classification of the underlying bundle. We also proposed a coordinatization scheme that mimics the construction of a bundle map to a universal bundle, which can be tuned to capture different global topologies using a persistence-based approach on characteristic classes.
A natural direction for future research is to establish a recovery theorem providing conditions under which a finite metric space $X$ and a feature map $\pi:X\to B$ uniquely determine a true circle bundle, independent of the choice of discrete, approximate local trivializations.
Beyond circle bundles, the theory and algorithms presented here could be extended in several ways. One could consider discrete, approximate analogs of bundles with other fibers, such as disjoint unions of circles $\coprod_{j}\mathbb{S}^{1}_{j}$, higher-dimensional spheres $\mathbb{S}^{n}$, or tori $\mathbb{T}^{n}$. Another promising direction is the study of more general topological objects, such as fibrations or stratified bundles, which allow for singularities in the base or fiber. These generalizations open the door to applications involving more complex or singular data structures.
A Technical Details
Here we present some background definitions and results which are used throughout the paper.
A.1 The Group $\text{Homeo}(\mathbb{S}^{1})$
In this section, we prove some technical results about $\text{Homeo}(\mathbb{S}^{1})$. We denote the connected components of orientation-preserving and -reversing homeomorphisms by $\text{Homeo}_{+}(\mathbb{S}^{1})$ and $\text{Homeo}_{-}(\mathbb{S}^{1})$, respectively. A model for the universal cover $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$ of each component is given by
$\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}=\{f\in\text{Homeo}(\mathbb{R}):f(x+1)=f(x)\pm 1\}$ (29)
We then define $\widetilde{\text{Homeo}(\mathbb{S}^{1})}=\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}\coprod\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$.
We denote the group of isometries of $\mathbb{S}^{1}$ by $\text{Iso}(\mathbb{S}^{1})<\text{Homeo}(\mathbb{S}^{1})$ and denote the connected components by $\text{Iso}_{\pm}(\mathbb{S}^{1})$. We have a natural identification of $\text{Iso}_{+}(\mathbb{S}^{1})$ with $\mathbb{S}^{1}$ itself via the map $R^{+}:\mathbb{S}^{1}\to\text{Iso}_{+}(\mathbb{S}^{1})$ which sends $e^{i\theta}$ the counter-clockwise rotation $R^{+}_{\theta}$ (equivalently, $R^{+}_{\theta}$ denotes left-translation in $\mathbb{S}^{1}$ by $e^{i\theta}$). Similarly, every orientation-reversing isometry is expressible as $R_{\theta}^{-}$ for some $\theta$, where $R_{\theta}^{-}(z)=e^{i\theta}\bar{z}$ for all $z\in\mathbb{S}^{1}$. We can model the universal covers $\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})$ as subsets of $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$:
$\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})=\{T^{\pm}_{a}(x)=\pm x+a\ :\ a\in\mathbb{R}\}$ (30)
Proposition A.1.
Suppose $\tilde{f}\in\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}$ is $L$-Lipschitz. Then, $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})\leq\frac{L-1}{L}$.
Proof.
Assume $\tilde{f}\in\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}$ (the proof for $\tilde{f}\in\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$ is similar). For any $x\in(0,1)$, we have $\tilde{f}(x)-\tilde{f}(0)\leq Lx$ and $\tilde{f}(1)-\tilde{f}(x)\leq L(1-x)$. On the other hand, we also have $\tilde{f}(0)<\tilde{f}(x)<\tilde{f}(1)$ for any $x\in(0,1)$, so the point $(x,\tilde{f}(x))$ is constrained to lie in a polygonal region. In particular, the quantity $|y-(x+\tilde{f}(0))|$ is maximized at the corner points $(\frac{1}{L},\frac{L-1}{L})$ and $(1-\frac{1}{L},\frac{L-1}{L})$. It follows that $d_{\infty}(\tilde{f},T^{+}_{\tilde{f}(0)})\leq\frac{L}{L-1}$.
Proposition A.2.
[…]
Proof.
Assume $f$ is orientation-preserving (the proof when $f$ reverses orientation is similar), and let $\tilde{f}$ be any lift of $f$. Let $w,z\in\mathbb{S}^{1}$, and let $\tilde{w},\tilde{z}\in\mathbb{R}$ be lifts of $w$ and $z$ such that $d_{\mathbb{R}}(\tilde{w},\tilde{z})\leq\frac{1}{2}$ (such a pair of lifts always exists). We have
[…]
Proposition A.3.
Suppose $\tilde{f},\tilde{g}\in\widetilde{\text{Homeo}(\mathbb{S}^{1})}$ are lifts of $f,g\in\text{Homeo}(\mathbb{S}^{1})$, respectively. If $d_{\infty}(\tilde{f},\tilde{g})<\varepsilon$, then $d_{\infty}(f,g)<2\pi\varepsilon$.
Proof.
Choose any $z\in\mathbb{S}^{1}$ and let $\tilde{z}\in\mathbb{R}$ be any lift. Then,
$d_{\mathbb{S}^{1}}(f(z),g(z))\$$)\displaystyle=\ d_{\mathbb{S}^{1}}(\exp\circ\tilde{f}(\tilde{z}),\ \exp\circ\tilde{g}(\tilde{z}$))
$)\displaystyle\leq\ 2\pi d_{\mathbb{R}}(\tilde{f}(\tilde{z}),\ \tilde{g}(\tilde{z}$))
$<2\pi\varepsilon$
Proposition A.4.
If $f\in\text{Homeo}_{\pm}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,R_{\log f(1)}^{\pm})\leq 2\pi(\frac{L-1}{L})$.
Proof.
By Proposition A.2, any lift $\tilde{f}$ of $f$ is $L$-Lipschitz, so by Proposition A.1, any lift satisfies $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})<\frac{L-1}{L}$. Finally, using Proposition A.3, we obtain $d_{\infty}(f,R^{\pm}_{\log f(0)})=d_{\infty}(f,\exp\circ T_{\tilde{f}(0)})<2\pi(\frac{L-1}{L})$.
A.2 Circle Bundles Over $\mathbb{S}^{2}$ and $\mathbb{RP}^{2}$
In this section, we provide the details of the constructions and classifications of the two non-trivial circle bundle models described in the Implementations section. Since both models are associated with quotients of $SO(3)$, we begin by describing principal $\mathbb{S}^{1}$-bundle structures for $SO(3)$ and its universal cover $\mathbb{S}^{3}$.
For any $v\in\mathbb{S}^{2}$ and $\theta\in\mathbb{R}$ let $R_{v,\theta}\in SO(3)$ denote the counterclockwise rotation by $\theta$ about $v$. Fixing a $v_{0}\in\mathbb{S}^{2}$, we obtain a right $\mathbb{S}^{1}$-action defined by $A\cdot e^{i\theta}=AR_{v_{0},\theta}$. One can easily verify that this action is free and that the orbits are precisely the fibers of the map $\pi:SO(3)\to\mathbb{S}^{2}$ defined by $\pi(A)=Av_{0}$. We conclude that $(SO(3),\pi,\mathbb{S}^{2})$ is a principal $\mathbb{S}^{1}$-bundle. Moreover, the isomorphism class of this bundle does not depend on the choice of $v_{0}$. To see this, observe that for any $v_{0},v_{1}\in\mathbb{S}^{2}$ and $B\in SO(3)$ such that $Bv_{0}=v_{1}$, the map $F:SO(3)\to SO(3)$ by $F(A)=AB^{-1}$ is an equivariant bundle isomorphism. We therefore take $v_{0}=e_{1}$ without loss of generality in what follows.
The group $SO(3)$ admits a (universal) double cover $\text{Spin}(3)\cong\mathbb{S}^{3}\subset\mathbb{H}$ via the map $\text{Adj}:\mathbb{S}^{3}\to SO(3)$ defined by $\text{Adj}_{q}(v)=qvq^{-1}$. Here we are identifying $\mathbb{S}^{2}$ with the pure imaginary unit quaternions via $(x,y,z)\leftrightarrow xi+yj+zk$. Thinking of $\mathbb{S}^{1}$ as a subgroup of the unit quaternions, we have a free right $\mathbb{S}^{1}$-action on $\mathbb{S}^{3}$ by group multiplication. Moreover, the continuous surjection $\widetilde{\pi}:\mathbb{S}^{3}\to\mathbb{S}^{2}$ defined by $\widetilde{\pi}(q)=\text{Adj}_{q}(e_{1})$ is invariant under this action, so $(\mathbb{S}^{3},\widetilde{\pi},\mathbb{S}^{2})$ is a principal $\mathbb{S}^{1}$ bundle (called the Hopf bundle). In particular, we have the following commutative diagram:
[…]
Note, however, that the two principal circle bundles are not isomorphic. Indeed, Adj is not a homeomorphism, nor is it $\mathbb{S}^{1}$-equivariant (in fact, one has $\text{Adj}_{q\cdot e^{i\theta}}=\text{Adj}_{q}\cdot e^{i(2\theta)}$). The Hopf bundle famously has Euler number -1 with respect to the standard orientation on $\mathbb{S}^{2}$ (for a proof, see [Hus94]). The following proposition immediately implies that the Euler number of the principal bundle $(SO(3),\pi,\mathbb{S}^{2})$ is 2-2, and will also be useful for classifying the circle bundles from our models in Section 7:
Proposition A.5.
Let $\xi=(E,\pi,B)$ be a principal $\mathbb{S}^{1}$-bundle, and for any $p\in\mathbb{Z}$, let $\varphi_{p}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ be the group homomorphism defined by $\varphi_{p}(z)=z^{p}$. Then $(\varphi_{p})_{*}\xi$ is a principal $\mathbb{S}^{1}$-bundle over the same base space with Euler number $e((\varphi_{p})_{*}\xi)=pe(\xi)$ and total space $E\times_{\varphi_{p}}\mathbb{S}^{1}\cong E/\mathbb{Z}_{p}$.
Proof.
Let $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{S}^{1}})$ be a cocycle representative of $\xi$ subordinate to some open cover $\mathcal{U}$ of $B$. By the results in Section 2.1, $\varphi_{p}^{*}\xi$ is a principal $\mathbb{S}^{1}$-bundle which is represented by $\varphi_{*}\Omega$. If $\theta\in\check{C}^{1}(\mathcal{U};\mathbb{R})$ is any lift of $\widetilde{\Omega}$ (i.e., $\Omega_{jk}(b)=e^{i\theta_{jk}(b)}$ for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$), then
$(\varphi_{*}\Omega)_{jk}(b)=\varphi\circ\Omega_{jk}(b)=\varphi(e^{i\theta_{jk}(b)})=e^{i(p\theta_{jk}(b))}$
for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $p\theta$ is a lift of $\varphi_{*}\Omega$. The components of the Euler class $e(\varphi_{*}\xi)$ are then given by
$e(\varphi_{*}\xi)_{jkl}$ $=p\theta_{jk}(b)+p\theta_{kl}(b)-p\theta_{jl}(b)$
$=p(\theta_{jk}(b)+\theta_{kl}(b)-\theta_{jl}(b))$
$=pe(\xi)_{jkl}$
We conclude $e(\varphi_{*}\xi)=pe(\xi)$.
Corollary A.6.
The principal $\mathbb{S}^{1}$-bundle $(SO(3),\pi,\mathbb{S}^{2})$ described above has Euler number 2-2 with respect to the standard orientation on $\mathbb{S}^{2}$.
Proof.
Let $\xi_{H}=(\mathbb{S}^{3},\widetilde{\pi},\mathbb{S}^{2})$ and $\xi_{SO(3)}=(SO(3),\widetilde{\pi},\mathbb{S}^{2})$ be the principal $\mathbb{S}^{1}$-bundles shown in (31), and consider the group homomorphism $\varphi:\mathbb{S}^{1}\to\mathbb{S}^{1}$ defined by $\varphi(z)=z^{2}$. The total space of $\varphi_{*}\xi_{H}$ is $\mathbb{S}^{3}/\mathbb{Z}_{2}\cong SO(3)$, and the adjoint map precisely satisfies $\text{Adj}_{q\cdot z}=\text{Adj}_{q}\cdot\varphi(z)$. We can therefore identify $\xi_{SO(3)}$ with $\varphi_{*}\xi_{H}$, so by Proposition A.5, $e(\xi_{SO(3)})=2e(\xi_{H})=-2$.
Since principal $\mathbb{S}^{1}$-bundles over $\mathbb{S}^{2}$ are completely determined by their Euler numbers, Proposition A.5 actually implies every such bundle can be realized as a quotient of the Hopf bundle. The quotient 3-manifold $\mathbb{S}^{3}/\mathbb{Z}_{p}$ is often denoted by $L(p,1)$ and called a lens space. Note that for any $p\in\mathbb{Z}$, we have $L(2p,1)\cong SO(3)/\mathbb{Z}_{p}$ and $(\varphi_{2p})_{*}\xi_{\text{Hopf}}\cong(\varphi_{p})_{*}\xi_{SO(3)}$.
The following proposition suggests how to construct the circle bundles over $\mathbb{RP}^{2}$:
Proposition A.7.
Suppose $\xi=(E,\pi,B)$ is a principal $G$-bundle, and $E$ and $B$ have free, proper right actions by a group $H$ such that:
  1. 1.
    $\pi:E\to B$ is $H$-equivariant,
  2. 2.
    There is a representation $\varphi:H\to\text{Aut}(G)$ such that $(e\cdot g)\cdot h=(e\cdot h)\cdot\varphi_{h}(g)$ for all $e\in E$, $g\in G$ and $h\in H$.
Then, the map $\pi^{\prime}:E/H\to B/H$ defined by $\pi^{\prime}([e])=[\pi(e)]$ is a fiber bundle with fiber $G$ and structure group $G\rtimes_{\varphi}H$.
Proof.
We give a sketch – the details are fairly straightforward and are left to the reader.
First, observe that the hypotheses imply $E$ admits a free right $(G\rtimes_{\varphi}H)$-action $e\cdot(g,h)=(e\cdot h)\cdot g$, and that the orbits of this action are precisely the fibers of the map $\Pi:E\to B/H$ defined by $\Pi(e)=[\pi(e)]$. That $\Pi$ is locally trivial follows directly from this property of $\pi$, so $(E,\Pi,B/H)$ is a principal $(G\rtimes_{\varphi}H)$-bundle. Define a left $(G\rtimes_{\varphi}H)$-action on $G$ by $(g,h)\cdot g^{\prime}=\varphi_{h}(g^{\prime})g^{-1}$ and form the associated bundle $(E\times_{G\rtimes_{\varphi}H}G,\Pi^{\prime},B/H)$ with fiber $G$. One can check that the assignment $(e,g)\longmapsto[e\cdot g]$ is a homeomorphism which fits into the commutative diagram below:
[…]
where $\pi^{\prime}:E/H\to B/H$ is the map defined in the proposition. We conclude that $(E/H,\pi^{\prime},B/H)$ is a fiber bundle with fiber $G$ and structure group $G\rtimes_{\varphi}H$.
Corollary A.8.
For any integer $p$, let $(L(2p,1),\pi_{2p},\mathbb{S}^{2})$ be the principal $\mathbb{S}^{1}$-bundle over $\mathbb{S}^{2}$ with Euler number 2p-2p, and consider the free right $\mathbb{Z}_{2}$-actions on $\mathbb{S}^{2}$ and $L(2p,1)$ defined respectively by $v\cdot(-1)=-v$ and $q]\cdot(-1)=[qj]$. The map $\pi_{2p}^{\prime}:L(2p,1)/\mathbb{Z}_{2}\to\mathbb{RP}^{2}$ defined by $\pi_{2p}^{\prime}([e])=[\pi_{2p}(e)]$ is a non-orientable circle bundle with (twisted) Euler number $\pm p$.
Now, suppose $X$ is a Hausdorff space which admits a continuous right $SO(3)$-action. The orbit $\mathcal{O}_{x}\subset X$ of any $x\in X$ is homeomorphic to the quotient of $SO(3)$ by the stabilizer subgroup $\text{stab}(x)=\{g\in SO(3):x\cdot g=x\}$. In particular, the map $T_{x}:SO(3)/\text{stab}(x)\to\mathcal{O}_{x}$ defined by $T_{x}([A])=x\cdot A$ is a homeomorphism.
If $\text{stab}(x)$ is trivial, then $\mathcal{O}_{x}\cong SO(3)$, and $\mathcal{O}_{x}$ inherits a principal $\mathbb{S}^{1}$-structure defined by the right $\mathbb{S}^{1}$-action $y\cdot e^{i\theta}=y\cdot R_{e_{1},\theta}$. The induced projection map $\widetilde{\pi}:\mathcal{O}_{x}\to\mathbb{S}^{2}$ is defined by $\widetilde{\pi}(x\cdot A)=\pi(A)=Ae_{1}$ for all $A\in SO(3)$, and $T_{x}$ is an equivariant bundle isomorphism.
Now suppose $x\in X$ has a $p$-fold rotational symmetry in the sense that $\text{stab}(x)\cong\mathbb{Z}_{p}$ for some positive integer $p$. The orbit $\mathcal{O}_{x}$ is then homeomorphic to the lens space $L(2p,1)\cong SO(3)/\mathbb{Z}_{p}$. Just as before, $\mathcal{O}_{x}$ inherits a principal $\mathbb{S}^{1}$-structure such that $T_{x}$ is an equivariant bundle isomorphism. We conclude $(\mathcal{O}_{x},\widetilde{\pi},\mathbb{S}^{2})$ is a principal $\mathbb{S}^{1}$-bundle with Euler number 2p-2p.
Lastly, suppose that in addition to a $p$-fold rotational symmetry in the plane perpendicular to $e_{1}$, $x\in X$ also has an additional 2-fold rotational symmetry with respect to this plane (imagine a prism). Then $\text{stab}(x)$ is the non-abelian group generated by $\{R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\}$, where $R_{e_{1},\theta}R_{e_{2},\pi}=R_{e_{2},\pi}R_{e_{1},-\theta}$. In particular, observe that $R_{e_{2},\pi}$ induces a free $\mathbb{Z}_{2}$-action on $SO(3)/\langle R_{e_{1},\frac{2\pi}{p}}\rangle$ given by $A]\cdot(-1)=[AR_{e_{2},\pi}]$. Moreover, one can easily verify that $\pi([A]\cdot(-1))=-\pi([A])$ and $([A]\cdot R_{e_{1},\theta})\cdot(-1)=([A]\cdot(-1))\cdot R_{e_{1},-\theta}$ for all $A\in SO(3)$. Thus, by Proposition A.7, the induced projection map $\pi^{\prime}:SO(3)/\langle R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\rangle\to\mathbb{RP}^{2}$ given by $\pi^{\prime}([A])=[\pi(A)]$ is a non-orientable circle bundle. In fact, using the adjoint map to identify $\mathbb{S}^{3}/\mathbb{Z}_{2}$ with $SO(3)$, this bundle is canonically isomorphic to the circle bundle $\pi^{\prime}_{2p}:L(2p,1)/\mathbb{Z}_{2}\to\mathbb{RP}^{2}$ described in Corollary A.8 (note in particular that $\text{Adj}_{j}=R_{e_{2},\pi}$).
A.3 Miscellaneous
Proposition A.9.
Suppose $r,R\in O(2)$ with $\det(r)=-1$ and $\det(R)=1$ ($r$ is a reflection and $R$ is a rotation). Then, $rRr=R^{-1}$.
Proof.
Let $\text{Adj}_{r}(R)=rRr^{-1}$. Note that $SO(2)\subset O(2)\subset GL(2,\mathbb{R})$, so $r$ and $R$ are invertible (in particular, $r=r^{-1}$). Moreover, $\text{det}(\text{Adj}_{r}(R))=1$, so $\text{Adj}_{r}$ is a map from $SO(2)$ to itself. In fact, $\text{Adj}_{r}$ is clearly an involution, so we must have $\text{Adj}_{r}\in\text{Aut}(SO(2))\cong\mathbb{Z}_{2}$. On the other hand, $\text{Adj}_{r}$ cannot be the identity since this would imply $rR=Rr$, which we know is false by the semidirect decomposition of $O(2)$. We conclude that $\text{Adj}_{r}$ is the inversion map on $SO(2)$, so $rRr=R^{-1}$.
Proposition A.10.
Suppose $A\in M_{n\times n}$ and $X\subset\mathbb{S}^{n-1}\subset\mathbb{R}^{n}$ such that $|Ax|<\varepsilon$ for all $x\in X$. If $d_{H}(X,\mathbb{S}^{n-1})<\delta$ (with respect to the Euclidean metric), then $|A|_{F}<\frac{\varepsilon\sqrt{n}}{1-\delta}$.
Proof.
Choose any $v\in\mathbb{S}^{n-1}$, and let $x\in X$ such that $|v-x|<\delta$. Then,
[…]
Thus,
$|A|_{F}\ <\ \sqrt{n}\left(\frac{|A|_{F}}{\sqrt{n}}\delta\ +\ \varepsilon\right)$
or equivalently, $|A|_{F}<\frac{\varepsilon\sqrt{n}}{1-\delta}$.
Proposition A.11.
Suppose $V=\{v_{j}\}_{j=0}^{n}\subset\mathbb{S}^{1}\subset\mathbb{R}^{2}$ such that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$. Then $V$ has a unique weighted Karcher mean with respect to the geodesic distance on $\mathbb{S}^{1}$ (and any weights).
Proof.
The fact that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$ implies $d_{\mathbb{S}^{1}}(v_{j},v_{0})<\frac{\pi}{2}$ for all $j$, so by the triangle inequality, the diameter of $V$ (with respect to $d_{\mathbb{S}^{1}}$) is less than $\pi$. Since $\mathbb{S}^{1}$ is geodesically convex on any subset of diameter less than $\pi$, $V$ has a unique weighted Karcher mean with respect to any weights.

### v3

@@
-Discrete Approximate Circle Bundles
+Brad Turow, Jose A. Perea
 Abstract
@@
 To illustrate the practical utility and viability of our algorithms, we present applications to both real and synthetic datasets from computer vision (e.g., modeling optical flow).
+The paper is accompanied by an open-source software package, with full documentation and tutorials, enabling reproducible implementation of the proposed algorithms and experiments, including those used to generate the figures in this paper
 1 Background And Problem Statement
 1.1 Motivation And Context
-In recent years, it has been observed that high-dimensional datasets arising in fields such as computer vision, computational chemistry, and motion tracking often lie near highly non-linear, low-dimensional manifolds [TSL00, RS00, BN03].
-These manifolds can exhibit complex geometric and topological structure, motivating the development of a range of techniques for identifying and leveraging such structures in tasks including dimensionality reduction, statistical analysis, pattern recognition, and data visualization [CL06, Sap01].
-By modeling the intrinsic manifold geometry, these methods have been shown to improve compression, denoising, and feature extraction in domains where traditional linear techniques fail to capture the underlying structure of the data.
+It has been observed in recent years that high-dimensional datasets arising in fields such as computer vision, computational chemistry, and motion tracking often lie near highly non-linear, low-dimensional manifolds [29, 21, 2].
+These manifolds can exhibit complex geometric and topological structure, motivating the development of a range of techniques for identifying and leveraging such structures in tasks including dimensionality reduction, statistical analysis, pattern recognition, and data visualization [6, 23].
+By modeling the intrinsic manifold geometry, these methods have been shown to improve compression, denoising, and feature extraction in domains where traditional linear techniques fail to capture the underlying structure of the data [15].
+In practice, complex non-linear structure in large, high-dimensional datasets is often difficult to observe directly; it is therefore essential to have techniques for inferring global structure from purely local computations on small portions of the data.
 In this paper, we focus on a particular class of datasets which can be modeled as circle bundles.
 Roughly speaking, a circle bundle is a continuously-varying family of (topological) circles parameterized by another space, called the base space of the bundle.
-The individual circles are called fibers of the bundle, and the union of the fibers is the total space.
-The parametrization is given by a projection map $\pi:E\to B$ from the total space to the base space – $E$ is locally homeomorphic to the product $B\times\mathbb{S}^{1}$, but globally the structure may be twisted.
+The individual circles are called the fibers of the bundle, and the union of the fibers is the total space.
+The parametrization is given by a projection map $\pi:E\to B$ from the total space to the base space; in particular $E$ is locally homeomorphic to a product $U\times\mathbb{S}^{1}$, for $U\subseteq B$ open, but globally the structure may be twisted.
 For instance, the torus and the Klein bottle can each be characterized as the total space of a circle bundle with base space $\mathbb{S}^{1}$, but the two are not homeomorphic (see Figure 1).
-Similarly, the unit tangent bundle on the 2-sphere is a circle bundle with base $\mathbb{S}^{2}$, but one can show that there is no bundle isomorphism between this bundle and the product space $\mathbb{S}^{2}\times\mathbb{S}^{1}$.
+Similarly, the 3-manifold $SO(3)$ admits a non-trivial circle bundle structure over $\mathbb{S}^{2}$.
+Indeed, consider the map $\pi:SO(3)\to\mathbb{S}^{2}$ which assigns an (oriented) axis of rotation to each element of $SO(3)$; the fibers of this map are copies of $SO(2)\cong\mathbb{S}^{1}$, but one can show that $SO(3)\ncong\mathbb{S}^{2}\times\mathbb{S}^{1}$.
 […]
@@
 On the machine learning side, various datasets related to computer vision and graphics have been shown to have circle bundle structures.
-For instance, in [Ada+20], Adams et al.
+For instance, in [1], Adams et al.
 use local measurements and a feature map to support a torus model for high-contrast optical flow data.
-Similarly, in [Car+08], the authors propose that a certain space of high-contrast natural image patches is concentrated around a Klein bottle embedded in high dimension.
-Beyond these examples, datasets arising from 3D objects with rotational symmetries — such as projection images of macromolecules in cryo-electron microscopy — naturally give rise to circle bundle structures, where each fiber corresponds to rotations of the object about a fixed axis [SS14].
-The possible circle bundle structures over a metric space $B$ are completely described by a pair of discrete invariants called characteristic classes.
+Similarly, in [5], it is shown that frequently-occurring high-contrast patches from natural images concentrate around a Klein bottle embedded in a high-dimensional Euclidean space.
+It is worth noting that in each of these contexts, a direct persistent homology computation fails to satisfactorily confirm the topology of the proposed underlying model.
+Figure 2 shows an illustration of how this failure can occur, and shows how the local-to-global methodologies we propose in this paper successfully recover the underlying structure.
+Refer to caption (a) Persistence diagrams of the original toroidal dataset $X$ Refer to caption (b) $X$ colored by the circular coordinate derived from its most persistent 1-dimensional cohomology class [20] Refer to caption (c) Fiberwise angle resulting from trivializing the circle bundle over $\mathbb{S}^{1}$ given by (b) Refer to caption
+Figure 2: A noisy sample $X$ from a torus whose fiber radius continuously oscillates.
+The persistence diagrams show only a single significant class in dimension 1 (recall that the non-zero Betti numbers for the torus are $\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$).
+Figure (b) shows the dataset colored according to circular coordinates computed with DREiMac [19].
+Figure (c) shows the same dataset colored according to the fiber coordinates computed using the software package accompanying this paper.
+Beyond these examples, datasets arising from 3D objects with rotational symmetries — such as projection images of macromolecules in cryo-electron microscopy — naturally give rise to circle bundle structures, where each fiber corresponds to rotations of the object about a fixed axis [26].
+Here the underlying model is typically a 3-manifold such as $SO(3)$, and a direct persistence computation to capture the topology is simply intractable.
+On the other hand, one can infer the underlying manifold from purely local computations by leveraging the bundle structure, as we demonstrate with a synthetic dataset of 3D density functions in Section 7.3.
+The different possible circle bundle structures over a topological space $B$ are completely described, up to isomorphism, by a pair of discrete invariants called characteristic classes.
 In particular, a non-vanishing class indicates the bundle is non-trivial, so there is no hope of constructing a coordinatization for the whole space as a product.
 Additionally, characteristic classes are invariant under continuous perturbations, so one might hope to be able to stably recover them from a noisy finite sampling.
-Another key feature of characteristic classes is that they can be computed from local measurements alone, which makes them an ideal tool for working with complex, high-dimensional data not amenable to more direct techniques.
+Another key feature of characteristic classes is that they can be computed from local measurements alone, which makes them suitable for working with complex, high-dimensional data not amenable to more direct techniques.
 Lastly, the characteristic classes of a bundle can be used to construct a so-called bundle map from the total space to a more familiar space which respects topological structure.
-This observation has recently been leveraged in works such as [Lee+25] and [SP23] to develop non-linear dimensionality reduction techniques based on vector bundles.
+This observation has recently been leveraged in works such as [14] and [24] to develop non-linear dimensionality reduction techniques based on vector bundles.
 In this paper, we extend this approach to circle bundles.
@@
   1. 1.
-    We define notions of discrete approximate circle bundles and local trivializations and show how these relate to other well-studied objects such as discrete approximate Čech cocycles.
-Building on results of [Tin22] and [SP23], we give conditions under which a discrete approximate circle bundle can be meaningfully identified with a true class of circle bundles.
+    We define notions of discrete approximate circle bundles (Definition 3.8) in terms of approximate local trivializations (Definition 3.6), and show how these relate to other well-studied objects such as discrete approximate Čech cocycles (Proposition 3.15).
+Building on results of [30] and [24], we give conditions under which a discrete approximate circle bundle can be meaningfully identified with a true class of circle bundles (Theorem 3.42).
   2. 2.
-    We show that every true circle bundle is uniquely characterized by an orientation class and a twisted Euler class, and provide an algorithm for computing these classes from a family of discrete approximate local trivializations.
-We further introduce a metric on each space of discrete approximate local trivializations and show that two trivializations which are sufficiently close produce the same characteristic classes.
-In particular, we prove that if a discrete approximate local trivialization is sufficiently close to a true one, the characteristic classes obtained from our algorithm are the same.
+    We show that every true circle bundle is uniquely characterized up to isomorphism by an orientation class and a twisted Euler class (Theorem 2.12), and provide an algorithm for computing these classes from a system of discrete approximate local trivializations (Algorithm 1).
+We further introduce a metric on each space of discrete approximate local trivializations and show that two trivializations which are sufficiently close produce the same characteristic classes (Corollary 4.3).
+In particular, we prove that if a discrete approximate local trivialization is sufficiently close to a true one, the characteristic classes obtained from our algorithm are the same (Corollary 4.5).
   3. 3.
-    Given a discrete approximate circle bundle over a base space $B$ and an open cover $\mathcal{U}$ for $B$, we introduce a weights filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and a corresponding notion of persistence for characteristic class representatives.
-We then propose a scheme for coordinatization and dimensionality reduction which produces a map from the dataset to the total space of the associated bundle $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, where $V(2,d)$ denotes the Stiefel manifold of 2-planes in $\mathbb{R}^{d}$.
-Our pipeline can be tuned to be compatible with any stage of the weights filtration on $\mathcal{N}(\mathcal{U})$, yielding a point cloud with a different global topology at each stage.
+    Given a discrete approximate circle bundle over a base space $B$ and an open cover $\mathcal{U}$ for $B$, we introduce a weights filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and a corresponding notion of persistence for characteristic class representatives (Section 5.1).
+We then propose a scheme for coordinatization and dimensionality reduction which produces a map from the dataset to the total space of the associated bundle $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, where $V(2,d)$ denotes the Stiefel manifold of 2-planes in $\mathbb{R}^{d}$ (Section 6).
+Our pipeline is designed to be compatible with any stage of the weights filtration on $\mathcal{N}(\mathcal{U})$, yielding multi-scale representations and coordinatizations of the data consistent with its multi-scale global topology.
   4. 4.
-    We provide implementations of our algorithms on several real and synthetic datasets related to imaging and optical flow.
-In Section 7.1, we use these tools to confirm the torus model proposed in [Ada+20].
-A more detailed analysis of our findings will appear in an upcoming companion paper.
+    We demonstrate implementations of our algorithms on several real and synthetic datasets related to imaging and optical flow.
+In Section 7.1, we use these tools to confirm the torus model proposed in [1].
+A more detailed analysis of our findings can be found in our companion paper [32].
+  5. 5.
+    We provide an open-source software package implementing the algorithms introduced in this paper, available at https://github.com/bradturow/Circle_Bundles, with full documentation and tutorials hosted at https://circle-bundles.readthedocs.io
 1.3 Prior Work
-A number of approaches to local-to-global inference for analyzing geometrically and topologically complex datasets have recently emerged.
+Recently, there has been an emergence of new approaches in local-to-global inference for analyzing geometrically and topologically complex datasets.
 The basic principle in each case is to model small pieces of the dataset and analyze how these parts glue together to determine the global shape.
 Given a dataset $X$, a common technique is to choose a feature map $\pi:X\to S$ to some metric space $S$ (e.g., $\mathbb{R}$) and study how the level sets vary.
-The Mapper pipeline takes this approach to represent a point cloud as a graph with nodes representing local clusters.
-Similarly, in [SP22] the authors describe a pipeline called FibeRed which uses local PCA together with a feature map $\pi:X\to S$ to obtain a low-dimensional representation of $X$ which is faithful to both local geometry and global topology.
+The Mapper pipeline takes this approach to represent a point cloud as a graph with nodes representing local clusters [27].
+Similarly, in [25] the authors describe a pipeline called FibeRed which uses linear dimensionality reduction — e.g., via PCA or ISOMAP — on the fibers of a feature map $\pi:X\to S$ to a Riemannian manifold $S$, to obtain a low-dimensional representation of $X$ which captures both local geometry and global topology.
 Another aspect of local-to-global analysis is the synchronization of local coordinate systems.
-Cohomological obstructions to synchronization are explored in works such as [GBM19] using the language of principal $G$-bundles.
-We note, however, that the theory provided there makes the highly restrictive assumption that the underlying bundle is flat.
-In [SP23], the authors introduce a theory of discrete approximate Euclidean vector bundles which handles non-flat cases in the context of $O(d)$-synchronization.
-Building on the language and technology developed in [Tin22], they show that under good conditions, one can uniquely and stably identify discrete approximate Euclidean vector bundles with true ones.
+Cohomological obstructions to synchronization are explored in works such as [9] using the language of principal $G$-bundles.
+We note, however, that the theory provided there makes the restrictive assumption that the underlying bundle is flat (which is not the case in most of the applications we consider).
+In [24], the authors introduce a theory of discrete approximate Euclidean vector bundles which handles non-flat cases in the context of $O(d)$-synchronization.
+Building on the language and techniques developed in [30], they show that under appropriate conditions, one can uniquely and stably identify discrete approximate Euclidean vector bundles with true ones.
 They also provide stable algorithms for computing characteristic classes of vector bundles from discrete approximate cocycle representatives.
@@
 The rest of the paper is structured as follows:
-Section 3: Mathematical Background.
-We review the essential mathematical foundations required for our results, including the general theory of fiber bundles, the classification of circle bundles and cohomology with local coefficients.
+Section 2: Mathematical Preliminaries.
+We review the essential mathematical elements required for our results, including the general theory of fiber bundles, the classification of circle bundles and cohomology with local coefficients.
 This section establishes notation and key concepts that will be used throughout the paper.
 Section 3: Discrete Approximate Circle Bundles.
-We introduce the notion of discrete approximate circle bundles and explore their relationship with discrete approximate Čech cocycles and classifying maps as defined in [SP23].
+We introduce our proposed notion of discrete approximate circle bundles and explore its relationship with discrete approximate Čech cocycles and classifying maps as defined in [24].
 We also discuss how these discrete constructions relate to classical circle bundles.
-Section 5: Algorithms for Characteristic Classes. We present algorithms to compute characteristic classes from discrete approximate local trivializations, and prove that our algorithms are stable.
+Section 4: Algorithms for Characteristic Classes.
+We present algorithms to compute characteristic classes from discrete approximate local trivializations, and prove that our algorithms are robust to input perturbations (i.e., stable).
 Section 5: Weights Filtration and Persistence.
-Given a discrete approximate circle bundle over a metric space $B$ and an open cover $\mathcal{U}$ of $B$, we introduce a simplexwise filtration on $\mathcal{N}(\mathcal{U})$ and describe a type of persistent cohomology for characteristic class representatives with respect to this filtration.
+Given a discrete approximate circle bundle over a metric space $B$ and an open cover $\mathcal{U}$ of $B$, we introduce a filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and describe a type of persistent cohomology for characteristic class representatives with respect to this filtration.
 Section 6: Coordinatization Pipeline.
-We provide a pipeline for coordinatization and dimensionality reduction which integrates the Principal Stiefel Coordinates algorithm [Lee+25].
-The result is a map from the dataset into a Stiefel manifold which is, in some sense which we explain, compatible with a chosen stage of the associated weights filtration.
+We provide a pipeline for coordinatization and dimensionality reduction of circle bundle data, which integrates the Principal Stiefel Coordinates algorithm [14].
+The result is a function from the dataset into a Stiefel manifold which is, in some sense which we explain, compatible with a chosen stage of the associated weights filtration.
 Section 7: Implementations and Experiments.
-We demonstrate our methods on both synthetic and real datasets, applying the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations and showcasing the resulting computations of characteristic classes and coordinatization.
+We demonstrate our methods on both synthetic and real datasets, applying the sparse circular coordinates algorithm [20] to obtain discrete approximate local trivializations and showcasing the resulting computations of characteristic classes and coordinatization.
 Section 8: Future Work. We conclude with a discussion of potential extensions, open questions, and directions for further exploration in both theory and applications.
 2 Preliminaries
-2.1 Theory Of Fiber Bundles
-In this section, we briefly review the key aspects of the theory of fiber bundles which are relevant for us. Our exposition emphasizes a computational perspective. For additional details, see [Hus94].
+2.1 Fiber Bundles
+In this section, we briefly review the key aspects of the theory of fiber bundles which are relevant for us. Our exposition emphasizes a computational perspective. For additional details, see [12].
 Definition 2.1.
@@
 Any such map $\varphi$ is called a local trivialization of $\xi$.
-An open cover $\mathcal{U}$ of $B$ is called a trivializing cover for $\xi$ if there exists a local trivialization of $\xi$ over each $U_{j}\in\mathcal{U}$.
+An open cover $\mathcal{U}$ of $B$ is called a trivializing cover for $\xi$ if there exists a local trivialization of $\xi$ over each $U\in\mathcal{U}$.
 Two bundles $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are isomorphic if there exists a homeomorphism $f:E\to E^{\prime}$ such that $\pi^{\prime}\circ f=\pi$.
-Let $\text{Bun}_{F}(B)$ denote the set of isomorphism classes.
-Transition Maps: Given a fiber bundle $\xi=(E,\pi,B)$ with fiber $F$, an open cover $\mathcal{U}$ of $B$ and a family $\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times F\}_{j\in J}$ of local trivializations subordinate to $\mathcal{U}$, there exist unique continuous maps $\Omega_{jk}:U_{j}\cap U_{k}\to\text{Homeo}(F)$ such that $\varphi_{j}\circ\varphi_{k}^{-1}(b,v)=(b,\Omega_{jk}(b)(v))$ for any $(b,v)\in(U_{j}\cap U_{k})\times F$.
-The maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ are called the transition maps associated with the family of local trivializations.
+Let $\text{Bun}_{F}(B)$ denote the set of isomorphism classes of fiber bundles over $B$ with fiber $F$.
+Definition 2.2.
+Let $G$ be a topological group.
+A principal $G$-bundle is a fiber bundle $\xi=(E,\pi,B)$ with fiber $G$, satisfying the following properties: $E$ has a free right $G$-action so that $\pi(eg)=\pi(e)$ for all $(e,g)\in E\times G$ — i.e., the action is fiber-preserving — and which is transitive when restricted to each fiber $\pi^{-1}(b)$, $b\in B$.
+Moreover the local trivializations $\varphi:\pi^{-1}(U_{b})\to U_{b}\times G$ can be chosen to be $G$-equivariant, where the right $G$-action on $U_{b}\times G$ is group multiplication in $G$.
+Two principal $G$-bundles $(E,\pi,B)$ and $(E^{\prime},\pi^{\prime},B)$ are isomorphic (as principal $G$-bundles) if there exists a $G$-equivariant homeomorphism $f:E\to E^{\prime}$ so that $\pi^{\prime}\circ f=\pi$.
+Let $\text{Prin}_{G}(B)$ denote the set of isomorphism classes of principal $G$-bundles over $B$.
+Transition Maps: Given a fiber bundle $\xi=(E,\pi,B)$ with fiber $F$, a trivializing cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a family $\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times F\}_{j\in J}$ of local trivializations subordinate to $\mathcal{U}$, there exist unique continuous maps $\Omega_{jk}:U_{j}\cap U_{k}\to\text{Homeo}(F)$ such that $\varphi_{j}\circ\varphi_{k}^{-1}(b,v)=(b,\Omega_{jk}(b)(v))$ for any $(b,v)\in(U_{j}\cap U_{k})\times F$.
+Here $\text{Homeo}(F)$ denotes the set of homeomorphisms from $F$ to $F$, endowed with the compact-open topology.
+The maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$, indexed by the 1-simplices of the nerve of $\mathcal{U}$, are called the transition maps associated with the family of local trivializations.
 Note that the transition functions associated with any family of local trivializations must satisfy the cocycle condition:
-$\Omega_{jk}(b)\circ\Omega_{kl}(b)=\Omega_{jl}(b)\hskip 28.45274pt(b\in U_{jkl}\;,\;(jkl)\in\mathcal{N}(\mathcal{U}))$ (1)
+$\Omega_{jk}(b)\circ\Omega_{kl}(b)=\Omega_{jl}(b)\hskip 28.45274pt(b\in U_{j}\cap U_{k}\cap U_{l}\;,\;(jkl)\in\mathcal{N}(\mathcal{U}))$ (1)
 If we require the transition maps to be $G$-valued for some subgroup $G\leq\text{Homeo}(F)$, we say that $G$ is the structure group of the bundle.
@@
 Part A: Let $B$ and $F$ be topological spaces, and let $G$ be a topological group with a continuous left action on $F$.
-Given an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a collection $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ satisfying the cocycle condition, there exists a fiber bundle $\pi:E\to B$ with fiber $F$ and structure group $G$ that is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$.
-Part B: Suppose $\xi^{\prime}=(E^{\prime},\pi^{\prime},B^{\prime})$ is another fiber bundle with fiber $F$ and structure group $G$ which is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}^{\prime}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$.
-If the action of $G$ on $F$ is faithful, then $\xi\cong\xi^{\prime}$ if and only if there exist maps $\{\mu_{j}:U_{j}\to G\}_{j\in J}$ such that $\Omega_{jk}^{\prime}=\mu_{j}^{-1}\Omega_{jk}\mu_{k}$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
+Given an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a collection $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ of continuous maps satisfying the cocycle condition (1), there exists a fiber bundle $\pi:E\to B$ with fiber $F$ and structure group $G$ that is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ [12, Chapter 5, Theorem 3.2].
+Part B: Suppose $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ is another fiber bundle with fiber $F$ and structure group $G$ (endowed with the same left-action) which is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}^{\prime}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$.
+If the action of $G$ on $F$ is faithful, then $\xi\cong\xi^{\prime}$ if and only if there exist maps $\{\mu_{j}:U_{j}\to G\}_{j\in J}$ such that $\Omega_{jk}^{\prime}=\mu_{j}^{-1}\Omega_{jk}\mu_{k}$ for all $(jk)\in\mathcal{N}(\mathcal{U})$ [12, Chapter 5, Theorem 2.7].
 Note that if $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are fiber bundles with fiber $F$ over a common paracompact space $B$ with trivializing covers $\mathcal{U}$ and $\mathcal{U}^{\prime}$, then any common refinement $\mathcal{V}$ of $\mathcal{U}$ and $\mathcal{U}^{\prime}$ is a trivializing cover for $\xi$ and $\xi^{\prime}$, so we can compare the two bundles using this refinement (and the result does not depend on the choice of refinement).
-Associated Bundles: Let $\xi=(E,\pi,B)$ be a fiber bundle with fiber $F$ and structure group $G$, and let $F^{\prime}$ be a left $G$-space.
-The associated bundle to $\xi$ with fiber $F^{\prime}$ is the bundle $\xi^{\prime}=(E\times_{G}F^{\prime},\pi^{\prime},B)$ whose existence is guaranteed by the fiber bundle construction lemma (using the transition maps from any trivialization of $\xi$).
-The special case of $F^{\prime}=G$ with left action by group multiplication is called the principal $G-\textbf{bundle}$ associated with $\xi$.
-Note that for any space $B$, the associated bundle construction gives a bijective correspondence between fiber bundles with fiber $F$ and principal $\text{Homeo}(F)$-bundles over $B$.
+Associated Bundles: Let $\xi=(E,\pi,B)$ be a fiber bundle with fiber $F$ and structure group $G$.
+Let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be a trivializing cover for $\xi$ and let $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N(U)}}$ be the transition maps of a local trivializing of $\xi$ over $\mathcal{U}$.
+If $F^{\prime}$ is a left $G$-space, then Part A of the Fiber Bundle Construction Lemma defines a fiber bundle $\xi^{\prime}$ over $B$ with fiber $F^{\prime}$.
+The association $\xi\mapsto\xi^{\prime}$ is well-defined up to isomorphism of fiber bundles, and we call $\xi^{\prime}$ the associated bundle to $\xi$ with fiber $F^{\prime}$.
+The special case $F^{\prime}=G$ with left action by group multiplication is called the principal $G-\textbf{bundle}$ associated to $\xi$.
+When $\xi=(E,\pi,B)$ is a principal $G$-bundle and $F^{\prime}$ is a left $G$-space, one can construct $\xi^{\prime}$ directly without appealing to transition functions or isomorphism classes as follows: The multiplication $\cdot:(E\times F^{\prime})\times G\to E\times F^{\prime}$ given by $(e,v^{\prime})\cdot g=(eg\,,\,g^{-1}*v^{\prime})$ defines a right $G$-action on $E\times F^{\prime}$, and the orbit space $E\times_{G}F^{\prime}$ has a map $\pi^{\prime}:E\times_{G}F^{\prime}\to B$ given by $\pi^{\prime}([e,v^{\prime}])=\pi(e)$.
+One can show that $\xi^{\prime}=(E\times_{G}F^{\prime},\pi^{\prime},B)$ is a fiber bundle with fiber $F^{\prime}$.
+Finally, we note that for any space $B$, the associated bundle construction gives a bijective correspondence between $\text{Bun}_{F}(B)$ and $\text{Prin}_{\text{Homeo}(F)}(B)$, where $F$ is understood to be endowed with the tautological $\text{Homeo}(F)$-action by evaluation.
 If $\xi=(E,\pi,B)$ is a principal $G$-bundle and $\varphi:G\to H$ is a continuous group homomorphism, then $H$ has a continuous left $G$-action given by $g\cdot h=\varphi(g)h$.
@@
 This perspective can be articulated using the language of Čech cohomology, which we briefly describe below.
-Let $\mathcal{U}$ be an open cover of a topological space $B$ and let $\mathcal{F}$ be a presheaf of abelian groups on $B$.
+See [18, Chapter IX] for definitions.
+Let $\mathcal{U}=\{U_{\alpha}\}_{\alpha\in\Lambda}$ be an open cover of a topological space $B$ and let $\mathcal{F}$ be a presheaf of abelian groups on $B$.
 For $k\geq 0$, define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}:\check{C}^{k}(\mathcal{U};\mathcal{F})\to\check{C}^{k+1}(\mathcal{U};\mathcal{F})$ by
-$\check{C}^{k}(\mathcal{U};\mathcal{F})=\prod_{\alpha_{0}<\cdots<\alpha_{k}}\mathcal{F}(U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k}})$
-$(\delta^{k}\Omega)_{\alpha_{0}\cdots\alpha_{k+1}}=\sum_{j=0}^{k+1}(-1)^{j}\Omega_{\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}}|_{U_{\alpha_{0}}\cap\cdots\cap U_{\alpha_{k+1}}}$
-Then $(\check{C}^{*}(\mathcal{U};\mathcal{F}),\delta^{*})$ is a cochain complex, and the associated cohomology is called Čech cohomology subordinate to $\mathcal{U}$ with coefficients in $\mathcal{F}$.
-Note that if $\mathcal{F}$ is a presheaf of non-abelian groups, we can still define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}$ as above, replacing negation by group inversion in the coboundary formula.
-In general, however, we do not have $\ker\delta^{k+1}\subseteq\text{im }\delta^{k}$ so $\check{H}^{k}(\mathcal{U};\mathcal{F})$ is not well-defined for $k>0$.
-For $k=1$, we still have a well-defined cohomology set $\check{H}^{1}(\mathcal{U};\mathcal{F})$, though the classes do not have a natural group structure.
+$\check{C}^{k}(\mathcal{U};\mathcal{F})=\prod_{(\alpha_{0},\ldots,\alpha_{k})}\mathcal{F}(U_{\alpha_{0}\cdots\alpha_{k}})$
+$(\delta^{k}\Omega)_{\alpha_{0}\cdots\alpha_{k+1}}=\sum_{j=0}^{k+1}(-1)^{j}\Omega_{\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}}|_{U_{\alpha_{0}\cdots\alpha_{k+1}}}$
+where $U_{\alpha_{0}\cdots\alpha_{k}}=U_{\alpha_{0}}\cap\ldots\cap U_{\alpha_{k}}$ and $\alpha_{0}\cdots\widehat{\alpha_{j}}\cdots\alpha_{k+1}$ is the result of removing $\alpha_{j}$ from the original tuple.
+Then $(\check{C}^{*}(\mathcal{U};\mathcal{F}),\delta^{*})$ is a cochain complex, and the associated cohomology groups $\check{H}^{*}(\mathcal{U};\mathcal{F})$ are the Čech cohomology subordinate to $\mathcal{U}$ with coefficients in $\mathcal{F}$.
+Note that if $\mathcal{F}$ is a presheaf of non-abelian groups, we can still define $\check{C}^{k}(\mathcal{U};\mathcal{F})$ and $\delta^{k}$ as above, replacing the minus sign by group inversion in the coboundary formula.
+In general, however, we do not have $\text{im }\delta^{k}\subseteq\ker\delta^{k+1}$ so $\check{H}^{k}(\mathcal{U};\mathcal{F})$ is not well-defined for $k>0$.
+For $k=1$, we still have a well-defined cohomology set $\check{H}^{1}(\mathcal{U};\mathcal{F})$, though it does not have a natural group structure.
 Let $\mathcal{F}$ be a presheaf of groups over $B$, and let $\mathcal{U}$ be an open cover of $B$.
 If $\mathcal{V}$ is a refinement of $\mathcal{U}$, one has an induced map $\check{H}^{*}(\mathcal{U};\mathcal{F})\to\check{H}^{*}(\mathcal{V};\mathcal{F})$ in cohomology.
-The open covers of $B$ form a directed set with respect to refinements, so the corresponding Čech cohomology groups form a direct system of abelian groups.
-The direct limit is denoted $\check{H}^{*}(B;\mathcal{F})$ and called the Čech cohomology of $B$.
-If $B$ is paracompact, then for any topological group $G$, the fiber bundle construction lemma implies a one-to-one correspondence between $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(F)})$ and $\text{Bun}_{F}(B)$, where $\mathcal{C}_{\text{Homeo}(F)}$ denotes the sheaf of continuous $\text{Homeo}(F)$-valued functions on $B$.
-We state the following additional properties of Čech cohomology without proof. For further details, see [Bre97].
+The open covers of $B$ form a partially ordered set with respect to refinements, so the corresponding Čech cohomology groups (or sets) form a directed system (of groups or sets, respectively).
+The direct limit is denoted $\check{H}^{*}(B;\mathcal{F})$ and called the Čech cohomology of $B$ with coefficients in $\mathcal{F}$.
+If $B$ is paracompact, then the Fiber Bundle Construction lemma implies a one-to-one correspondence between $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(F)})$ and $\text{Bun}_{F}(B)$, where $\mathcal{C}_{G}$ denotes the sheaf of continuous $G$-valued functions on $B$ for a topological group $G$.
+That is, for $U\subseteq B$ open, $\mathcal{C}_{G}(U)=\text{Maps}(U,G)$ with multiplication in $G$.
+We state the following additional properties of Čech cohomology without proof. For further details, see [3].
 Proposition 2.3.
@@
   2. 2.
-    If $B$ has the homotopy type of a CW complex and $G$ is abelian, then $\check{H}^{*}(B;\underline{G})\cong H^{*}(B;G)$, where $\underline{G}$ denotes the sheaf of locally constant $G$-valued functions on $B$.
+    If $B$ has the homotopy type of a CW complex and $G$ is an abelian group, then there is an isomorphism between Čech and cellular cohomology, $\check{H}^{*}(B;\underline{G})\cong H^{*}(B;G)$, where $\underline{G}$ denotes the sheaf of locally constant $G$-valued functions on $B$.
   3. 3.
@@
   4. 4.
-    If $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced cochain map $\varphi_{*}:\check{Z}^{*}(B;\mathcal{C}_{G})\to\check{Z}^{*}(B;\mathcal{C}_{H})$ given by $(\varphi_{*}\Omega)_{\alpha_{0}\cdots\alpha_{k}}=\varphi(\Omega_{\alpha_{0}\cdots\alpha_{k}})$.
-Moreover, if $\varphi$ is a homotopy equivalence, then $\varphi_{*}$ induces an isomorphism in cohomology.
+    If $\mathcal{F}$ is a sheaf of abelian groups over a paracompact space $B$, then there is an isomorphism between sheaf and Čech cohomology, $H^{k}(B;\mathcal{F})\cong\check{H}^{k}(B;\mathcal{F})$, for all $k\geq 0$.
   5. 5.
-    If $F\to G\to H$ is a central extension of topological groups, there is an induced short exact sequence $\mathcal{C}_{F}\to\mathcal{C}_{G}\to\mathcal{C}_{H}$ of sheaves, which in turn gives rise to a short exact sequence in Čech cohomology.
+    If $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced cochain map $\varphi^{\#}:\check{C}^{*}(\mathcal{U};\mathcal{C}_{G})\to\check{C}^{*}(\mathcal{U};\mathcal{C}_{H})$ given by $(\varphi^{\#}\Omega)_{\alpha_{0}\cdots\alpha_{k}}=\varphi(\Omega_{\alpha_{0}\cdots\alpha_{k}})$.
+Moreover, if $\varphi$ is a homotopy equivalence, then $\varphi^{\#}$ induces an isomorphism in cohomology.
+  6. 6.
+    If $F\to G\to H$ is a central extension of topological groups, there is an induced short exact sequence $\mathcal{C}_{F}\to\mathcal{C}_{G}\to\mathcal{C}_{H}$ of sheaves, which in turn gives rise to an exact sequence in Čech cohomology.
 Moreover, if $F$ is abelian, there is an induced connecting homomorphism $\Delta:\check{H}^{1}(B;\mathcal{C}_{H})\to\check{H}^{2}(B;\mathcal{C}_{F})$ which yields a long exact sequence.
@@
 and $\pi^{\prime}:f^{*}E\to B^{\prime}$ is given by $\pi^{\prime}(b^{\prime},e)=b^{\prime}$.
-Note that if $\xi$ has fiber $F$ then so does $f^{*}\xi$.
+Note that if $\xi$ has fiber $F$, then so does $f^{*}\xi$.
 Moreover, one can show that homotopic maps yield isomorphic pullback bundles.
@@
 The construction is essentially unique in the sense that any two universal $G$-bundles are isomorphic and any two classifying spaces for $G$ are (weakly) homotopy equivalent.
-For any paracompact space $B$, there is a bijective correspondence between $[B,BG]$ and $\text{Prin}_{G}(B)$ given by $f]\mapsto[f^{*}\xi_{G}]$, where $\xi_{G}$ is a universal $G$-bundle.
+For any paracompact space $B$, there is a natural bijective correspondence between $[B,BG]$, the set of homotopy classes of maps from $B$ to $BG$, and $\text{Prin}_{G}(B)$.
+The correspondence is given by $f]\mapsto[f^{*}\xi_{G}]$, where $\xi_{G}$ is a universal $G$-bundle.
 Given any $\xi]\in\text{Prin}_{G}(B)$, a map $f:B\to BG$ such that $\xi\cong f^{*}\xi_{G}$ is called a classifying map for $\xi$.
-Lastly, we note that the classifying space construction is functorial in the following sense: if $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced map $B\varphi:BG\to BH$ with the property that if $f:B\to BG$ classifies a principal $G$-bundle $\xi$, then $B\varphi\circ f:B\to BH$ classifies $\varphi^{*}\xi$.
+Lastly, we note that the classifying space construction can be made functorial in the following sense: if $\varphi:G\to H$ is a continuous homomorphism of topological groups, there is an induced map $B\varphi:BG\to BH$ with the property that if $f:B\to BG$ classifies a principal $G$-bundle $\xi$, then $B\varphi\circ f:B\to BH$ classifies $\varphi^{*}\xi$.
 2.2 Cohomology With Local Coefficients
-We provide a brief overview of cohomology with local coefficients which we use in our classification of circle bundles. For additional details, see [Eil47], [Hat02].
+We provide a brief overview of cohomology with local coefficients which we use in the classification of circle bundles. For additional details, see [8, Chapter V] or [10, Section 3.I].
 Definition 2.4.
-Given a topological space $B$ and an abelian group $\Gamma$, a local system for $B$ with coefficients in $\Gamma$ is a locally constant sheaf $\mathcal{L}$ with stalk $\Gamma$.
-We will primarily be interested in the case where $\Gamma$ is a topological group and $B$ a path-connected paracompact space (in particular, $B$ has a universal cover $p:\widetilde{B}\to B$).
-In this setting, a local system on $B$ admits a variety of important equivalent characterizations:
+Given a topological space $B$ and an abelian group $G$, a local system for $B$ with coefficients in $G$ is a locally constant sheaf $\mathcal{L}$ with stalk $G$.
+Note that if $B$ is path-connected and has a universal cover $p:\widetilde{B}\to B$, then a local system on $B$ can equivalently be defined as a sheaf $\mathcal{L}$ with stalk $G$ such that the pullback $p^{*}\mathcal{L}$ is a constant sheaf on $\widetilde{B}$.
+To see this, note that if $\mathcal{L}$ is a locally-constant sheaf, then $p^{*}\mathcal{L}$ is also locally-constant, but a locally-constant sheaf on a simply-connected space is constant.
+Conversely, if $\mathcal{L}$ is a sheaf on $B$ and $p^{*}\mathcal{L}$ is constant, then clearly $\mathcal{L}$ is locally-constant.
+We will primarily be interested in the case where $G$ is a discrete group and $B$ a path-connected paracompact space.
+In this setting, a local system on $B$ admits several equivalent characterizations:
 Proposition 2.5.
-Suppose $G$ is a topological group and $B$ is a path-connected space with universal cover $p:\widetilde{B}\to B$.
-A local system on $B$ with coefficients in $G$ can be equivalently defined as any one of the following:
+Suppose $G$ is a discrete group and $B$ is a path-connected, locally path-connected and semilocally simply connected space.
+Endow the automorphism group $\text{Aut}(G)$ with the discrete topology, and let $\underline{\text{Aut}(G)}$ denote the associated locally-constant sheaf on $B$.
+There are bijective correspondences between:
   1. 1.
-    A sheaf $\mathcal{L}$ with stalk $G$ such that $p^{*}\mathcal{L}$ is a constant sheaf on $\widetilde{B}$,
+    Isomorphism classes of local systems on $B$ with stalk $G$,
   2. 2.
-    A monodromy representation $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$,
+    Isomorphism classes of flat fiber bundles on $B$ with fiber $G$ and structure group $\text{Aut}(G)$,
   3. 3.
-    A flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$,
+    Conjugacy classes of homomorphisms $\rho:\pi_{1}(B)\to\text{Aut}(G)$ (called monodromy representations),
   4. 4.
-    A Čech cocycle $\omega\in\check{Z}^{1}(B;\underline{\text{Aut}(G)})$.
+    Cohomology classes $\omega]$ in the Čech cohomology set $\check{H}^{1}(B;\underline{\text{Aut}(G)})$.
 Proof.
-If $\mathcal{L}$ is a locally-constant sheaf, then $p^{*}\mathcal{L}$ is also locally constant.
-Since $\widetilde{B}$ is simply-connected, $\mathcal{L}$ must be constant.
-Conversely, if $\mathcal{L}$ is a sheaf on $B$ and $p^{*}\mathcal{L}$ is constant, then clearly $\mathcal{L}$ is locally-constant.
-1 $\to$ 2: Since the unit interval $I=[0,1]$ is contractible, the sheaf $\gamma^{*}\mathcal{L}$ on $I$ is constant for any loop $\gamma:I\to B$ based at $b_{0}$.
-We therefore have an induced automorphism of $G$ by identifying $(\gamma^{*}\mathcal{L})_{0}$ with $(\gamma^{*}\mathcal{L})_{1}$, and the automorphisms induced by homotopic loops coincide.
-We therefore have a map $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$, and one can easily check that this map is a representation.
-2 $\to$ 3: Note that $p:\widetilde{B}\to B$ is a principal $\pi_{1}(B,b_{0})$-bundle, where $\pi_{1}(B,b_{0})$ acts on $\widetilde{B}$ by deck transformations.
-In particular, this bundle is flat since $\pi_{1}(B,b_{0})$ is a discrete group.
-A monodromy representation gives a left $\pi_{1}(B,b_{0})$-action on $G$ by automorphisms, so we can form the associated bundle $\widetilde{B}\times_{\rho}G\to B$.
+1 $\leftrightarrow$ 2: Suppose $\mathcal{L}$ is a local system over $B$ with stalk $G$.
+Choose an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ for $B$ and isomorphisms $\{\phi_{j}:\mathcal{L}|_{U_{j}}\to\underline{G}\}_{j\in J}$.
+Each map $\phi_{j}\circ\phi_{k}^{-1}:\underline{G}\to\underline{G}$ is locally-constant (hence constant) on $\underline{G}$, thus determines an element $g_{jk}\in\text{Aut}(G)$.
+The $g_{jk}$’s satisfy the cocycle condition, so the collection of all these maps can be interpreted as a Čech cocycle $g\in\check{Z}^{1}(\mathcal{U};\underline{\text{Aut}(G)})$.
+By the Fiber Bundle Construction Lemma, $g$ determines a flat fiber bundle with fiber $G$ and structure group $\text{Aut}(G)$ (and furthermore, the isomorphism class of the bundle does not depend on the choice of trivializing cover).
+Conversely, given a flat fiber bundle $\xi$ with fiber $G$ and structure group $\text{Aut}(G)$, the sheaf of local sections of $\xi$ is a local system on $B$ with stalk $G$.
+One can check that these constructions are inverses up to isomorphism.
+2 $\leftrightarrow$ 3: Suppose we have a representation $\rho:\pi_{1}(B)\to\text{Aut}(G)$.
+The universal cover $p:\widetilde{B}\to B$ of $B$ is a principal $\pi_{1}(B)$-bundle, where $\pi_{1}(B)$ acts on $\widetilde{B}$ from the left by deck transformations (and taking inverses yields the corresponding right action).
+In particular, this bundle is flat since $\pi_{1}(B)$ is a discrete group.
+Since $\rho$ gives a left $\pi_{1}(B)$-action on $G$ by automorphisms, so we can form the associated bundle $\widetilde{B}\times_{\rho}G\to B$.
+Observe that if $\rho^{\prime}:\pi_{1}(B)\to\text{Aut}(G)$ is another representation in the same conjugacy class as $\rho$, then by the Fiber Bundle Construction Lemma, the resulting bundles are isomorphic.
+Conversely, let $\mathcal{L}$ be a local system on $B$ with stalk $G$.
+Choose any $b_{0}\in B$ and isomorphism $\mathcal{L}_{b_{0}}\cong G$, as well as a loop $\gamma:I\to B$ based at $b_{0}$.
+Since $I=[0,1]$ is contractible, the sheaf $\gamma^{*}\mathcal{L}$ is constant, and the restriction maps induce isomorphisms $(\gamma^{*}\mathcal{L})_{0}\to(\gamma^{*}\mathcal{L})(I)\to(\gamma^{*}\mathcal{L})_{1}$.
+We therefore obtain an induced map $T_{\gamma}\in\text{Aut}(G)$ via our identification of $(\gamma^{*}\mathcal{L})_{1}=(\gamma^{*}\mathcal{L})_{0}=\mathcal{L}_{b_{0}}$ with $G$.
+One can check that the automorphisms induced by homotopic loops coincide, and that the assignment $\gamma]\longmapsto T_{\gamma}$ defines a representation $\rho:\pi_{1}(B,b_{0})\to\text{Aut}(G)$.
+Note that if $b_{1}\in B$, there exists a path $\widetilde{\gamma}:[0,1]\to B$ from $b_{0}$ to $b_{1}$, and we have $\pi_{1}(B,b_{1})\cong\pi_{1}(B,b_{0})$ via the assignment $\gamma]\longmapsto[\widetilde{\gamma}*\gamma*\widetilde{\gamma}^{-1}]$.
+Similarly, different identifications $\mathcal{L}_{b_{0}}\cong G$ yield homomorphisms which differ by conjugation.
+Thus, an isomorphism class of local systems with stalk $G$ determines a representation $\rho:\pi_{1}(B)\to\text{Aut}(G)$ up to conjugacy.
 2 $\leftrightarrow$ 4: This equivalence is immediately implied by the Fiber Bundle Construction Lemma.
-3 $\to$ 1: If $\xi$ is a flat fiber bundle over $B$ with fiber $G$ and structure group $\text{Aut}(G)$, the sheaf of locally-constant sections of $\xi$ is a locally-constant sheaf on $B$ with stalk $L$.
-Remark 2.5.
-The equivalences above all descend to the level of isomorphism classes. For example, two locally-constant sheaves on $B$ are isomorphic if and only if they induce equivalent monodromy representations.
 We will primarily be interested in local systems with stalk $\mathbb{Z}$:
@@
 A local system with stalk $\mathbb{Z}$ is called an orientation sheaf.
-Given a cocycle $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, we will denote the orientation sheaf associated with $\omega$ by $\mathbb{Z}_{\omega}$.
-In particular, if $B$ is a smooth manifold, the orientation sheaf associated with $w_{1}(TB)$ is called the orientation sheaf of the manifold.
-By Proposition 2.4, we can think of an orientation sheaf over $B$ as the sheaf of locally-constant sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}\mathbb{Z}$, where $B_{\omega}$ is a double cover of $B$.
+Given a path-connected, semilocally simply connected space $B$, an open cover $\mathcal{U}$ of $B$ and a cocycle $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ (where $\mathbb{Z}_{2}=\{1,-1\}\cong\mathrm{Aut}(\mathbb{Z}))$, we will denote the orientation sheaf corresponding to $\omega$ by $\mathbb{Z}_{\omega}$.
+Remark 2.7.
+If $B$ is a smooth manifold, the orientation sheaf corresponding to any representative of the Stiefel-Whitney class $w_{1}(TB)\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ of its tangent bundle, is called an orientation sheaf for the manifold.
+By Proposition 2.5, $B$ has an orientation sheaf $\mathcal{O}$, unique up to isomorphism.
+Let $\mathcal{U}$ be an open cover of a path-connected, locally path-connected, semilocally simply connected space $B$, and let $\mathbb{Z}_{\omega}$ denote the orientation sheaf on $B$ associated with a cocycle $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$.
+Proposition 2.5 implies we can think of $\mathbb{Z}_{\omega}$ as the sheaf of locally-constant sections of an associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}\mathbb{Z}$, where $B_{\omega}$ is a double cover of $B$ and $B_{\omega}\to B$ is a principal $\mathbb{Z}_{2}$-bundle.
 More generally, given a topological group $G$ with a $\mathbb{Z}_{2}$-action, we will be interested in sheaves of sections of associated bundles of the form $B_{\omega}\times_{\mathbb{Z}_{2}}G$:
@@
 Given a topological group $G$ with a $\mathbb{Z}_{2}$-action, let $\mathcal{P}_{G,\omega}$ be the presheaf over $B$ defined by
-$\mathcal{P}_{G,\omega}(V)=\{\text{families }\{f_{j}\in\mathcal{C}_{G}(U_{j}\cap V)\}:\ f_{j}|_{U_{j}\cap U_{k}\cap V}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}\cap V}\}$
+$\mathcal{P}_{G,\omega}(V)=\{\text{families }\{f_{j}\in\mathcal{C}_{G}(U_{j}\cap V)\}:\ f_{j}|_{U_{jk}\cap V}=\omega_{jk}\cdot f_{k}|_{U_{jk}\cap V}\}$
 The sheaf $\mathcal{C}_{G,\omega}$ of continuous $G$-valued functions on $B$ twisted by $\omega$ is the sheafification of $\mathcal{P}_{G,\omega}$.
-Note that $\mathcal{C}_{G,\omega}$ is equivalently defined as the sheaf of sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}G$, where $p_{\omega}:B_{\omega}\to B$ denotes the orientation double cover represented by $\omega$.
-If $G$ is a discrete group, we will write $G_{\omega}$ instead of $\mathcal{C}_{G,\omega}$.
+Note that $\mathcal{C}_{G,\omega}$ is equivalently defined as the sheaf of sections of the associated bundle $B_{\omega}\times_{\mathbb{Z}_{2}}G$.
+Notation: If $G$ is a discrete group, we will write $G_{\omega}$ instead of $\mathcal{C}_{G,\omega}$.
 Proposition 2.9.
@@
 Proof.
-Let $\{f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$ such that $f_{j}|_{U_{j}\cap U_{k}}=\omega_{jk}\cdot f_{k}|_{U_{j}\cap U_{k}}$ whenever $U_{j}\cap U_{k}$ is nonempty, and let $\{f^{\prime}_{j}=\phi_{j}\cdot f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$.
+Let $\{f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$ such that $f_{j}|_{U_{jk}}=\omega_{jk}\cdot f_{k}|_{U_{jk}}$ whenever $U_{j}\cap U_{k}$ is nonempty, and let $\{f^{\prime}_{j}=\phi_{j}\cdot f_{j}\in\mathcal{C}_{G}(U_{j})\}_{j\in J}$.
 Since $\omega=\omega^{\prime}\cdot\delta^{0}\phi$, we have
-$\tilde{f}_{j}|_{U_{j}\cap U_{k}}=\phi_{j}\cdot f_{j}|_{U_{j}\cap U_{k}}=(\phi_{j}\omega_{jk})\cdot f_{k}|_{U_{j}\cap U_{k}}=\phi_{k}\cdot f_{k}|_{U_{j}\cap U_{k}}=\tilde{f}_{k}|_{U_{j}\cap U_{k}}$
+$\tilde{f}_{j}|_{U_{jk}}=\phi_{j}\cdot f_{j}|_{U_{jk}}=(\phi_{j}\omega_{jk})\cdot f_{k}|_{U_{jk}}=\phi_{k}\cdot f_{k}|_{U_{jk}}=\tilde{f}_{k}|_{U_{jk}}$
 whenever $U_{j}\cap U_{k}$ is nonempty.
@@
 The details are not important for us, but we note that when $B$ is paracompact, the singular and Čech cohomology groups with coefficients in $\mathcal{L}$ are isomorphic.
-We also have the following generalization of Poincare duality:
+We also have the following generalization of Poincaré duality [31, Theorem 2.4]:
 Proposition 2.10.
-Let $B$ be a closed connected $n$-manifold, and let $\mathcal{L}$ be a local system of abelian groups.
-Then $H^{k}(B;\mathcal{L})\cong H_{n-k}(B;\mathcal{L}\otimes\mathcal{O})$ for all $k$, where $\mathcal{O}$ denotes the orientation sheaf of $B$.
+Let $M$ be a closed connected $n$-manifold, and let $\mathcal{L}$ be a local system of abelian groups.
+Then $H^{k}(M;\mathcal{L})\cong H_{n-k}(M;\mathcal{L}\otimes\mathcal{O})$ for all $k$, where $\mathcal{O}$ denotes the orientation sheaf of $M$.
 2.3 Classification Of Circle Bundles
@@
 $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})=\{f\in\text{Homeo}_{+}(\mathbb{R}):\ f(x+1)=f(x)+1\ \text{ for all }x\}$ (4)
-The projection map $\widetilde{\exp}:\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is defined by $\widetilde{\exp}(\tilde{f})=\text{exp}\circ\tilde{f}\circ\text{exp}^{-1}$, where $\exp:\mathbb{R}\to\mathbb{S}^{1}$ is given by $\exp(\theta)=e^{2\pi i\theta}$.
-Notice that sequence (2) splits, and the section $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ induced by complex conjugation gives rise to a semidirect decomposition $\text{Homeo}(\mathbb{S}^{1})\cong\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}$.
-In more detail, we obtain a right $\mathbb{Z}_{2}$-action on $\text{Homeo}_{+}(\mathbb{S}^{1})$ via the assignment $\pm 1\longmapsto\text{Adj}_{r(\pm 1)}$ which induces a group multiplication law $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\text{Adj}_{r(\omega_{1})}(\Lambda_{2}),\omega_{1}\omega_{2})$.
-The map $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$ is then a group isomorphism.
-Thus, for any open cover $\mathcal{U}$ of a paracompact space $B$, $\varphi$ induces an isomorphism on cohomology, where the cocycle condition is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
-$\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$
-Consequently, for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, there is a natural bijection between cocycles $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ satisfying $\deg_{*}(\Omega)=\omega$ and cochains $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$ satisfying $\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}\Lambda_{kl}=\Lambda_{jl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U})$.
-Alternatively, we can interpret $\omega$ as a local system and $\Lambda$ as a Cech cocycle in $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})$:
+[…]
+Notice that sequence (2) splits, and the section $s:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ induced by complex conjugation, gives rise to a semidirect decomposition $\text{Homeo}(\mathbb{S}^{1})\cong\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{s}\mathbb{Z}_{2}$.
+Explicitly, we obtain a right $\mathbb{Z}_{2}$-action on $\text{Homeo}_{+}(\mathbb{S}^{1})$ via the assignment $\pm 1\longmapsto\text{Adj}_{s(\pm 1)}$ defined by $\text{Adj}_{s(\pm 1)}(\Lambda)=s(\pm 1)\Lambda s(\pm 1)^{-1}$.
+This induces a group multiplication law $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\text{Adj}_{s(\omega_{1})}(\Lambda_{2}),\omega_{1}\omega_{2})$.
+The map $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{s}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ given by $\varphi(\Lambda,\omega)=\Lambda s(\omega)$ is then a group isomorphism.
+Thus, for any open cover $\mathcal{U}$ of a paracompact space $B$, $\varphi$ induces an isomorphism on cohomology, where the cocycle condition $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$ is given explicitly by
+$\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}(\Lambda_{kl})=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$
+Consequently, for any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, there is a natural bijection between cocycles $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ satisfying $\deg_{*}(\Omega)=\omega$ and cochains $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$ satisfying $\Lambda_{jk}\text{Adj}_{r(\omega_{jk})}(\Lambda_{kl})=\Lambda_{jl}$ for all $(jkl)\in\mathcal{N}(\mathcal{U})$.
+Alternatively, we can interpret $\omega$ as a local system and $\Lambda$ as a Čech cocycle in $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})$:
 $\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})=\coprod_{\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\\$ (5)
@@
 Proposition 2.11.
-Suppose $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. The following sequence is short exact:
+Suppose $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ for some open cover $\mathcal{U}$ of a paracompact space $B$. The following sequence of sheaves is short exact:
 $1\to\mathbb{Z}_{\omega}\to\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}\xrightarrow{\widetilde{\exp}_{*}}\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega}\to 1$ (6)
-Moreover, the sheaf $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so the induced connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ is an isomorphism.
+Moreover, the sheaf $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so the induced connecting homomorphism
+$\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
+is an isomorphism.
 Proof.
@@
 At the level of stalks, sequence (6) reduces to exact sequence (3), hence it is also exact as a sequence of sheaves of abelian groups.
-By Proposition 2.2, we therefore obtain a long exact sequence in Cech cohomology, and in particular, a connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
-Moreover, $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ is contractible, so $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, so Proposition 2.2 implies $\Delta_{\omega}$ is an isomorphism by exactness.
-The proposition above immediately leads to the following classification result:
-Proposition 2.11.
-For any open cover $\mathcal{U}$ of a paracompact space $B$, there is a bijective correspondence
+By Proposition 2.3, we therefore obtain a long exact sequence in Čech cohomology, and in particular, a connecting homomorphism $\Delta_{\omega}:\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1}),\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
+Moreover, $\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$ is contractible, so $\mathcal{C}_{\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1}),\omega}$ is soft, and Proposition 2.3 implies $\Delta_{\omega}$ is an isomorphism by exactness.
+The proposition above immediately leads to the main classification result of this section:
+Theorem 2.12.
+For any open cover $\mathcal{U}$ of a paracompact space $B$, there is a natural bijective correspondence
 $\check{H}^{1}(\mathcal{U};\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})\ \cong\ \coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}}_{2})}\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
-via the assignment $\Omega]\longmapsto\left(w_{1}(\Omega),\tilde{e}(\Omega)\right)$ given by $w_{1}(\Omega)=[\deg_{*}(\Omega)]$ and $\tilde{e}(\Omega)=\Delta_{\omega}([\Omega])$.
+via the assignment $\Omega]\longmapsto\left(\mathsf{w}_{1}(\Omega),\tilde{e}(\Omega)\right)$ given by $\mathsf{w}_{1}(\Omega)=[\deg_{*}(\Omega)]$ and $\tilde{e}(\Omega)=\Delta_{\omega}([\Omega])$.
 We therefore have a complete classification of circle bundles over paracompact spaces in terms of a pair of discrete invariants.
-As we will later show, these two invariants are closely connected to the first Stiefel-Whitney class and Euler class for real plane bundles, so we will refer to $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ respectively as the Stiefel-Whitney class (or orientation class) and the twisted Euler class of the circle bundle represented by $\Omega$.
+As we will later show, these two invariants are closely connected to the first Stiefel-Whitney class and Euler class for real plane bundles, so we will refer to $\mathsf{w}_{1}(\Omega)$ and $\tilde{e}(\Omega)$ respectively as the Stiefel-Whitney class (or orientation class) and the twisted Euler class of the circle bundle represented by $\Omega$.
 Remark 2.13.
-The notation in Proposition 2.11 is slightly abusive, since each cocycle $\omega$ within a cohomology class determines a different local system $\mathbb{Z}_{\omega}$.
+The notation in Theorem 2.12 is slightly abusive, since each cocycle $\omega$ within a cohomology class determines a different local system $\mathbb{Z}_{\omega}$.
 On the other hand, by Proposition 2.9, the local systems determined by any two cohomologous cocycles are isomorphic.
 In particular, if $\omega$ is a coboundary, then $\mathbb{Z}_{\omega}\cong\underline{\mathbb{Z}}$.
-Observe that when $B$ is a locally finite 1-dimensional CW complex (e.g., a graph), the Proposition 2.11 implies that circle bundles over $B$ are completely determined by their Stiefel-Whitney classes.
+Observe that when $B$ is a locally finite 1-dimensional CW complex (e.g., a graph), then Theorem 2.12 implies that circle bundles over $B$ are completely classified by the first Stiefel-Whitney class.
 In the case when $B$ is a compact connected surface, the classification is also especially simple and explicit:
 Proposition 2.14.
-Let $B$ be a closed connected 2-manifold, and let $\omega\in H^{1}(B;\mathbb{Z}_{2})$ determine a double cover $p_{\omega}:B_{\omega}\to B$. Then
+Let $B$ be a closed connected 2-manifold, and let $\omega]\in H^{1}(B;\mathbb{Z}_{2})$. Then
 […]
@@
 Proof.
-Twisted Poincare duality asserts that for any compact connected $n$-manifold $M$ and locally constant sheaf $\mathcal{L}$ of Abelian groups on $M$ (with fiber $G$), one has $H^{*}(M;\mathcal{L})\cong H_{n-*}(M;\mathcal{L}\otimes\mathcal{O})$, where $\mathcal{O}$ denotes the orientation sheaf of $M$.
+Twisted Poincare duality 2.10 asserts that for any compact connected $n$-manifold $M$ and locally constant sheaf $\mathcal{L}$ of Abelian groups on $M$ (with fiber $G$), one has $H^{*}(M;\mathcal{L})\cong H_{n-*}(M;\mathcal{L}\otimes\mathcal{O})$, where $\mathcal{O}$ denotes the orientation sheaf of $M$.
 Moreover, $H_{0}(M;\mathcal{L})$ is isomorphic to the coinvariants of the monodromy action $\omega_{\mathcal{L}}:\pi_{1}(M)\to\text{Aut}(G)$ induced by $\mathcal{L}$, i.e.,
 $H_{0}(M;\mathcal{L})\cong G/\langle\omega_{\mathcal{L}}(\gamma)\cdot a-a\ |\ a\in G,\gamma\in\pi_{1}(M)\rangle$
-Thus, for any compact connected surface $B$ and orientation character $\omega\in H^{1}(B;\mathbb{Z}_{2})$, one has
+Thus, for any compact connected surface $B$ and orientation character $\omega]\in H^{1}(B;\mathbb{Z}_{2})$, one has
 […]
-Since $\mathbb{Z}_{\omega}\otimes\mathcal{O}\cong\underline{\mathbb{Z}}$ if and only if $\omega+w_{1}(B)=0$ (or, equivalently, $\omega=w_{1}(B)^{-1}=w_{1}(B)$), the proposition follows.
-In light of the proposition above, for any compact connected surface $B$ and $\omega\in Z^{1}(B;\mathbb{Z}_{2})$, one can define a twisted fundamental class $B]_{\omega}$ which generates $\check{H}_{0}(B;\mathbb{Z}_{\omega})$ – this choice is unique when $\omega\neq w_{1}(B)$ and unique up to sign when $\omega=w_{1}(B)$.
+Since $\mathbb{Z}_{\omega}\otimes\mathcal{O}\cong\underline{\mathbb{Z}}$ if and only if $\omega]+\mathsf{w}_{1}(B)=0$ (or, equivalently, $\omega]=\mathsf{w}_{1}(B)^{-1}=\mathsf{w}_{1}(B)$), the proposition follows.
+In light of the proposition above, for any compact connected surface $B$ and $\omega]\in H^{1}(B;\mathbb{Z}_{2})$, one can define a twisted fundamental class $B]_{\omega}$ which generates $H_{2}(B;\mathbb{Z}_{\omega})$ – this choice is unique when $\omega]\neq\mathsf{w}_{1}(B)$ and unique up to sign when $\omega]=\mathsf{w}_{1}(B)$.
 The pairing $\langle\tilde{e},[B]_{\omega}\rangle$ then gives an explicit bijection between $\check{H}^{2}(B;\mathbb{Z}_{\omega})$ and its image in $\mathbb{Z}$.
@@
 2.4 Reduction Of Structure Group
-The structure group $\text{Homeo}(\mathbb{S}^{1})$ is rather unwieldly – in particular, it is disconnected, non-abelian and infinite-dimensional, making it challenging to work with from a computational perspective.
+The group $\text{Homeo}(\mathbb{S}^{1})$ is disconnected, non-abelian and infinite-dimensional, making it challenging to work with from a computational perspective.
 In this section, we identify reductions of this structure group to simpler ones.
@@
   2. 2.
-    The cohomology class $\Omega_{\xi}\in\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ lifts to a class in $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$,
+    The cohomology class $\Omega_{\xi}]\in\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}(\mathbb{S}^{1})})$ lifts to a class in $\check{H}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})$,
   3. 3.
-    The Stiefel-Whitney class $w_{1}(\Omega_{\xi})\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ vanishes.
+    The orientation class $\mathsf{w}_{1}([\Omega_{\xi}])\in\check{H}^{1}(B;\underline{\mathbb{Z}_{2}})$ vanishes.
 The next proposition implies another useful characterization of orientable circle bundles:
 Proposition 2.16.
-The inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ is a homotopy equivalence.
+The inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ given by multiplication in $\mathbb{S}^{1}$ is a homotopy equivalence.
 Proof.
 Consider the map $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined by $\tau(f)=f(1)$.
-Observe that $\tau$ is a continuous surjection, and the fiber of $\tau$ over each $z\in\mathbb{S}^{1}$ contains the orientation-preserving homeomorphisms which map $1$ to $z$.
-One can easily check that each $\tau^{-1}(z)$ is contractible, so $\tau$ is a homotopy inverse for $\iota$.
+Observe that $\tau$ is a continuous surjection, and the fiber of $\tau$ over each $z\in\mathbb{S}^{1}$ is comprised of the orientation-preserving homeomorphisms which map $1$ to $z$.
+One can check that each $\tau^{-1}(z)$ is contractible, and a theorem of Smale [28] then implies that $\tau$ is a homotopy inverse for $\iota$.
 An immediate consequence of the proposition above is that one can always reduce the structure group of an orientable circle bundle $\xi$ to $\mathbb{S}^{1}$ itself, which means every orientable circle bundle admits a principal $\mathbb{S}^{1}$-structure.
@@
 This is consistent with our earlier observation that the (twisted) Euler number of a circle bundle is only well-defined up to sign until an orientation class representative for the bundle is specified.
-More generally, Proposition 2.15 implies that the inclusion of $\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}=\text{Iso}(\mathbb{S}^{1})$ into $\text{Homeo}(\mathbb{S}^{1})$ is a homotopy equivalence.
+More generally, Proposition 2.16 implies that the inclusion of $\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}\cong\text{Iso}(\mathbb{S}^{1})$ into $\text{Homeo}(\mathbb{S}^{1})$ is a homotopy equivalence (where $\text{Iso}(\mathbb{S}^{1})$ denotes isometries of $\mathbb{S}^{1}$ with respect to geodesic distance).
 It follows that over paracompact spaces, any circle bundle, no matter how wild, admits a family of trivializations with $\text{Iso}(\mathbb{S}^{1})$-valued transition maps.
-While the inclusion $\iota:\mathbb{S}^{1}\to\text{Homeo}_{+}(\mathbb{S}^{1})$ induces a one-to-one correspondence between oriented circle bundles and principal $\mathbb{S}^{1}$-bundles, we note that the retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ defined in the proof of Proposition 2.15 is not a group homomorphism, hence does not induce a cochain map $\check{C}^{1}(B;\mathcal{C}_{\text{Homeo}_{+}(\mathbb{S}^{1})})\to\check{C}^{1}(B;\mathcal{C}_{\mathbb{S}^{1}})$ which descends to an isomorphism on cohomology.
-In particular, the following proposition shows that no retract can induce such a map:
-Proposition 2.16.
-There does not exist a retract $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ which is also a group homomorphism.
-Proof.
-One can show $\text{Homeo}_{+}(\mathbb{S}^{1})$ is a perfect group in the sense that every element is expressible as a product of commutators (see [Ghy01] Proposition 5.11).
-On the other hand, if $\tau:\text{Homeo}_{+}(\mathbb{S}^{1})\to\mathbb{S}^{1}$ is a group homomorphism, we must have $\tau([f,g])=[\tau(f),\tau(g)]$ for all $g,h\in\text{Homeo}_{+}(\mathbb{S}^{1})$.
-Since $\mathbb{S}^{1}$ is Abelian and $\text{Homeo}_{+}(\mathbb{S}^{1})$ is perfect, it follows immediately that $\tau(f)=1$ for all $f$.
-We conclude that $\tau$ cannot be a retract onto $\mathbb{S}^{1}$.
-Thus, given an arbitrary family of trivializations for a circle bundle, it may be difficult to find explicit homotopies so that the new transitions maps take values in $\text{Iso}(\mathbb{S}^{1})$.
-On the other hand, if we define the supremum distance on $\text{Homeo}(\mathbb{S}^{1})$ by
-$d_{\infty}(f,g)=\sup_{z\in\mathbb{S}^{1}}d_{\mathbb{S}^{1}}(f(z),g(z))\\$ (7)
-where $d_{\mathbb{S}^{1}}$ denotes the geodesic distance metric on $\mathbb{S}^{1}$, the proposition below states that well-behaved homeomorphisms of $\mathbb{S}^{1}$ are close to isometries.
-Since the proof requires some technical digressions, we defer it to the appendix.
-Proposition 2.17.
-If $f\in\text{Homeo}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\leq 2\pi(\frac{L-1}{L})$.
-Note in particular that $d_{\infty}(f,\text{Iso}(\mathbb{S}^{1}))\to 0$ as $L\to 1$.
 From a computational perspective, the structure group $\text{Iso}(\mathbb{S}^{1})\cong\mathbb{S}^{1}\rtimes\mathbb{Z}_{2}$ is vastly preferable to $\text{Homeo}(\mathbb{S}^{1})$.
-In particular, $\text{Iso}(\mathbb{S}^{1})$ is a 1-dimensional Lie group (with two connected components), and it has a natural identification with the matrix group $O(2)$ which gives rise to the following commutative diagram of split extensions:
+In particular, $\text{Iso}(\mathbb{S}^{1})$ is a 1-dimensional Lie group (with two connected components), and it has a natural identification with the group $O(2)$ of $2\times 2$ orthogonal matrices, which gives rise to the following commutative diagram of split extensions:
 […]
@@
 $\mathbb{Z}$$\mathbb{R}$${SO(2)}{\mathbb{Z}}$$\widetilde{\text{Homeo}}_{+}(\mathbb{S}^{1})$$\text{Homeo}_{+}(\mathbb{S}^{1})$$\scriptstyle{\exp}$$\scriptstyle{\iota}$$\scriptstyle{\iota}$$\scriptstyle{\widetilde{\exp}}$ (8)
-where $\exp:\mathbb{R}\to SO(2)$ is given by $\exp(\theta)=\left(\begin{array}[]{cc}\cos(2\pi\theta)&-\sin(2\pi\theta)\\ \sin(2\pi\theta)&\cos(2\pi\theta)\\ \end{array}\right)$.
-Passing to Cech cohomology, it is clear that the Stiefel-Whitney and twisted Euler classes we defined to classify circle bundles correspond precisely to their classical counterparts for vector bundles.
-However, it is useful to know how these classes arise and can be explicitly computed from arbitrary $\text{Homeo}(\mathbb{S}^{1})$-valued transition data.
-2.5 The Universal Circle Bundle And Dimensionality Reduction
-A classifying map can be used to model the total space of a fiber bundle as an embedded submanifold of a more familiar space.
-Specifically, if $G$ is a topological group and $\mu_{G}=(EG,p,BG)$ is a universal $G$-bundle, then for any principal $G$-bundle $\xi$ over a paracompact space $B$, we can construct a classifying map $f:B\to BG$ such that such that $f^{*}\mu_{G}=\xi$.
-This map is covered by a bundle map $F:E(\xi)\to EG$ which embeds the total space of $\xi$ into $EG$.
-More generally, if $\xi_{F}$ is the associated bundle of $\xi$ with fiber $F$, then $\xi_{F}$ is isomorphic to the pullback of the associated bundle $EG\times_{G}F\to BG$ along the same classifying map.
+[…]
+2.5 The Tautological Circle Bundle Over $\text{Gr}(2)$ And Dimensionality Reduction
+A classifying map can be used to identify the total space of a fiber bundle with a subspace of a more familiar object.
+Specifically, if $G$ is a topological group and $\mu_{G}=(EG,p,BG)$ is a universal $G$-bundle, then for any principal $G$-bundle $\xi=(E,\pi,B)$ over a paracompact space $B$, we can construct a classifying map $f:B\to BG$ with $f^{*}\mu_{G}\cong\xi$.
+This map is covered by a bundle map $\tilde{f}:E\to EG$ in the sense that $p\circ\tilde{f}=f\circ\pi$, and serves to map the total space of $\xi$ into $EG$.
+More generally, if $\xi^{\prime}=(E\times_{G}F,\pi^{\prime},B)$ is the associated bundle of $\xi$ with fiber a left $G$-space $F$, then $\xi^{\prime}$ is isomorphic to the pullback of the associated bundle $EG\times_{G}F\to BG$ along the same classifying map $f$.
+In this sense, we think of $EG\times_{G}F\to BG$ as “universal”, though the total space is generally not contractible.
 In the context of circle bundles, the associated bundle construction gives bijective correspondences
-$\text{Bun}_{\mathbb{S}^{1}}(B)\quad\leftrightarrow\quad\text{Prin}_{O(2)}(B)\quad\leftrightarrow\quad\text{Vect}_{2,\mathbb{R}}(B)$
-for any paracompact space $B$.
-It follows that up to homotopy, every circle bundle $\xi$ over $B$ can be realized (up to isomorphism) as the sphere bundle of a rank-2 real vector bundle $\xi_{\text{vect}}$.
-A model for the classifying space of real rank-2 vector bundles is the infinite Grassmann manifold $\text{Gr}(2)=\text{Gr}_{2}(\mathbb{R}^{\infty})$ consisting of all 2-planes in $\mathbb{R}^{\infty}$, topologized as the direct limit $\text{Gr}_{2}(\mathbb{R}^{2})\subset\text{Gr}_{2}(\mathbb{R}^{3})\subset\cdots$ (see [MS74] for more details).
-Identifying each plane with its projection matrix gives an embedding of $\text{Gr}(2)$ into $\mathbb{R}^{\infty\times\infty}$.
+$\text{Bun}_{\mathbb{S}^{1}}(B)\quad\leftrightarrow\quad\text{Prin}_{O(2)}(B)\quad\leftrightarrow\quad\text{Vect}_{2,\mathbb{R}}(B)$ (9)
+for any paracompact space $B$ (once one fixes the action of $O(2)$ on $\mathbb{S}^{1}$).
+It follows that up to homotopy, every circle bundle $\xi$ over $B$ can be realized (up to isomorphism) as the sphere bundle of a rank-2 real Euclidean vector bundle $\xi_{\text{vect}}$.
+A model for the classifying space of real rank-2 vector bundles is the infinite Grassmann manifold $\text{Gr}(2)=\text{Gr}_{2}(\mathbb{R}^{\infty})$ consisting of all 2-planes in $\mathbb{R}^{\infty}$, topologized as the direct limit $\text{Gr}_{2}(\mathbb{R}^{2})\subset\text{Gr}_{2}(\mathbb{R}^{3})\subset\cdots$ (see [17] for more details).
+Identifying each plane with its orthogonal projection matrix gives an embedding of $\text{Gr}(2)$ into $\mathbb{R}^{\infty\times\infty}$.
 The total space of the universal plane bundle $\mu$ over $\text{Gr}(2)$ can be modeled by the Stiefel manifold $V(2)=V(2,\infty)$:
-$V(2,\infty)=\{A\in\mathbb{R}^{\infty\times 2}:A^{T}A=I\}$
-The columns of each $A\in V(2)$ form an orthonormal basis for a plane in $\text{Gr}(2)$ whose projection matrix is given by $AA^{T}$.
-Thus, the universal plane bundle is then defined by the surjection $\pi_{Gr}:V(2)\to\text{Gr}(2)$ given by $\pi_{Gr}(A)=AA^{T}$.
-Given a circle bundle $\xi$ over a paracompact space $B$, the associated plane bundle $\xi_{\text{vect}}$ can be expressed as the pullback of $\mu$ along a classifying map $f:B\to\text{Gr}(2)$.
+$V(2,\infty)=\{A\in\mathbb{R}^{\infty\times 2}:A^{T}A=I_{2}\}$
+The columns of each $A\in V(2)$ form an orthonormal basis for a 2-plane in $\text{Gr}(2)$ whose orthogonal projection matrix is given by $AA^{T}$.
+Thus, the universal plane bundle is then defined by the surjection
+$\begin{array}[]{rccl}\pi_{Gr}:&V(2)&\longrightarrow&\text{Gr}(2)\\ &A&\mapsto&AA^{T}\end{array}$ (10)
+Given a circle bundle $\xi$ over a paracompact space $B$, the associated plane bundle $\xi_{\text{vect}}$ (9) can be expressed as the pullback of $\mu$ along a classifying map $f:B\to\text{Gr}(2)$.
 It follows that $\xi$ is isomorphic to the pullback of the associated circle bundle $\mu_{\circ}=V(2)\times_{O(2)}\mathbb{S}^{1}\to\text{Gr}(2)$ along $f$.
-In this sense $\mu_{\circ}$ is a universal circle bundle, and $\xi$ is classified by the same map as $\xi_{\text{vect}}$.
-From the perspective of data science and dimensionality reduction, a major advantage of choosing $\mu_{\circ}$ as our model for the universal circle bundle is that we can apply known techniques to obtain low-dimensional representations of point clouds in $V(2)\times_{O(2)}\mathbb{S}^{1}$.
-Our coordinatization pipeline for discrete approximate circle bundles integrates Principal Stiefel Coordinates [Lee+25] for this purpose.
+Thus, $\mu_{\circ}$ is “universal” in the sense described above, and $\xi$ is classified by the same map as $\xi_{\text{vect}}$.
+From the perspective of data science and dimensionality reduction, a major advantage of choosing $\mu_{\circ}$ as our model for the universal circle bundle is that we can apply known dimensionality reduction techniques in $V(2)$ to obtain low-dimensional representations of point clouds in $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, for small $d$.
+Our coordinatization pipeline for discrete approximate circle bundles (Section 6) integrates Principal Stiefel Coordinates [14] for this purpose.
 3 Discrete Approximate Circle Bundles
-In this section, we introduce discrete approximate notions of local trivializations, Cech cohomology and classifying maps for circle bundles.
+In this section, we introduce discrete approximate notions of local trivializations, Čech cohomology and classifying maps for circle bundles.
 The purpose of these objects is to represent true circle bundles with finite data which can be algorithmically, efficiently and stably computed from a point cloud and a feature map.
-3.1 Three Characterizations
+3.1 Discrete Approximate Local Trivializations
 Definition 3.1.
-A discrete $(\varepsilon,\delta)$-approximate circle bundle is a map $\pi:X\to B$ between metric spaces together with an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a family of maps $\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ with the following properties:
+Let $f:X\to Y$ be a function between metric spaces. The distortion of $f$ is defined by
+$\text{dist}(f)=\sup_{x,x^{\prime}\in X}|d_{Y}(f(x),f(x^{\prime}))-d_{X}(x,x^{\prime})|$
+Given functions $f:X\to Y$ and $g:Y\to X$, the codistortion of $f$ and $g$ is defined by
+$\text{codist}(f,g)=\sup_{x\in X,y\in Y}|d_{Y}(f(x),y)-d_{X}(x,g(y))|$
+Definition 3.2.
+Let $\pi:X\to B$ and $\pi^{\prime}:X^{\prime}\to B$ be functions between metric spaces.
+A discrete $(\beta,\varepsilon)$-approximate bundle map between $(X,\pi,B)$ and $(X^{\prime},\pi^{\prime},B)$ is a function $\varphi:X\to X^{\prime}$ with the following properties:
   1. 1.
-    $d_{H}(f_{l}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})<\delta$ (with respect to Euclidean distance) for every $U_{jkl}\in\mathcal{N}(\mathcal{U})$.
+    There exists a function $\psi:X^{\prime}\to X$ such that
+    $\max\{\text{dist}(\varphi),\ \text{dist}(\psi),\ \text{codist}(\varphi,\psi)\}\leq\varepsilon,$
   2. 2.
+    The functions $\varphi$ and $\psi$ satisfy
+    $\max\{\sup_{x\in X}d_{B}(\pi(x),\pi^{\prime}\circ\varphi(x)),\quad\sup_{x^{\prime}\in X^{\prime}}d_{B}(\pi^{\prime}(x^{\prime}),\pi\circ\psi(x^{\prime}))\}\leq\frac{\beta}{2}$
+Any function $\psi$ satisfying the properties above is called an almost-inverse for $\varphi$. Note that $\psi$ is also a discrete $(\beta,\varepsilon)$-approximate bundle map.
+We note several immediate consequences of the definition above for later use:
+Proposition 3.3.
+Suppose $\varphi:(X,\pi,B)\to(X^{\prime},\pi^{\prime},B)$ is a discrete $(\beta,\varepsilon)$-approximate bundle map and $\psi$ is an almost-inverse for $\varphi$.
+  1. 1.
+    Almost-Inverse Property:
+    $\sup_{x\in X}d_{X}(\psi\circ\varphi(x),x)\leq\varepsilon\hskip 14.22636pt\sup_{x^{\prime}\in X^{\prime}}d_{X}(\varphi\circ\psi(x^{\prime}),x^{\prime})\leq\varepsilon$
+  2. 2.
+    Restriction: Let $A=\pi^{-1}(U)\subseteq X$ and $A^{\prime}=(\pi^{\prime})^{-1}(U)\subseteq X^{\prime}$ for some set $U\subseteq B$.
+If $\varphi(A)\subseteq A^{\prime}$ and $\psi(A^{\prime})\subseteq A$, then $\varphi|_{A}$ and $\psi|_{A^{\prime}}$ are almost-inverse discrete $(\beta,\varepsilon)$-approximate bundle maps.
+  3. 3.
+    Composition: If $\varphi^{\prime}:X^{\prime}\to X^{\prime\prime}$ is a discrete $(\beta^{\prime},\varepsilon^{\prime})$-approximate bundle map from $(X^{\prime},\pi^{\prime},B)$ to $(X^{\prime\prime},\pi^{\prime\prime},B)$ with almost-inverse $\psi^{\prime}$, then $\varphi^{\prime}\circ\varphi$ is a discrete $(\beta+\beta^{\prime},\varepsilon+\varepsilon^{\prime})$-approximate bundle map from $(X,\pi,B)$ to $(X^{\prime\prime},\pi^{\prime\prime},B)$ and $\psi\circ\psi^{\prime}$ is an almost-inverse.
+Recall that in the classical setting, a local trivialization for a circle bundle $\pi:X\to B$ is a bundle isomorphism between $(\pi^{-1}(U),\pi|_{\pi^{-1}(U)},U)$ and $(U\times\mathbb{S}^{1},p_{U},U)$ for $U\subseteq B$ open (where $p_{U}$ denotes projection onto the first factor).
+We therefore turn our attention to discrete approximate bundle maps into product spaces of the form $U\times\mathbb{S}^{1}$.
+Definition 3.4.
+[…]
+[…]
+More generally, for any $r>0$, let $d_{\mathbb{S}^{1}}^{r}$ denote geodesic distance scaled by a factor of $r$ – that is, $d_{\mathbb{S}^{1}}^{r}(z,z^{\prime})=rd_{\mathbb{S}^{1}}(z,z^{\prime})$ for all $z,z\in\mathbb{S}^{1}$.
+We will denote the metric space $(\mathbb{S}^{1},d_{\mathbb{S}^{1}}^{r})$ by $\mathbb{S}^{1}_{r}$.
+Finally, given any metric space $B$, we denote the supremum product metric on $B\times\mathbb{S}^{1}_{r}$ by $d_{\infty}^{r}$.
+Remark 3.5.
+Since key results related to discrete approximate $O(2)$ Čech cohomology are most easily stated in terms of the Frobenius norm on $O(2)$ (which is closely related to the Euclidean norm on $\mathbb{S}^{1}$), we note the following identity which we will use frequently:
+$|z-w|=2\sin^{-1}\left(\frac{d_{\mathbb{S}^{1}}(z,w)}{2}\right),\hskip 28.45274ptz,w\in\mathbb{S}^{1}$ (11)
+Definition 3.6.
+[…]
+Remark 3.7.
+[…]
+We now introduce the notion of a discrete approximate circle bundle:
+Definition 3.8.
+A map $\pi:X\to B$ between metric spaces is called a discrete $(r,\beta,\varepsilon)$-approximate circle bundle if there is a good open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ (called a trivializing cover for $(X,\pi,B)$) with the following properties:
+  1. 1.
+    For each $(jkl)\in\mathcal{N}(\mathcal{U})$, with $j,k,l\in J$ not necessarily distinct, there exists some $x\in X$ such that $\bar{B}_{\beta}(\pi(x))\subseteq U_{jkl}$.
+  2. 2.
+    For each $j\in J$, there exists a discrete $(r,\beta,\varepsilon)$-approximate local trivialization $\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}^{1}_{r_{j}}$ for some $r_{j}\geq r$.
+The family $\{\varphi_{j}\}_{j\in J}$ is called a discrete $(r,\beta,\varepsilon)$-approximate trivialization of $\pi$, and the set of all such families subordinate to a fixed trivializing cover $\mathcal{U}$ is denoted by $T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$.
+Note that any $\Phi=\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}_{r_{j}}^{1}\}_{j\in J}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$ has an associated system of local circular coordinate functions $\mathfrak{F}(\Phi)=\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ given by $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$.
+We will see that, as in the classical setting, the relationships between local circular coordinates for a discrete approximate circle bundle encode its global topological structure.
+Definition 3.9.
+Let $\pi:X\to B$ be a function between metric spaces, and let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be an open cover of $B$.
+For $\varepsilon>0$ and $0\leq\delta<1$, a system of discrete $(\varepsilon,\delta)$-approximate local circular coordinates for $\pi$ is a family $\mathfrak{F}=\{f_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}\}_{j\in J}$ with the following properties:
+  1. 1.
     There exists a simplicial cochain $\Omega\in C^{1}(\mathcal{N}(\mathcal{U});O(2))$ such that
     $|f_{j}(x)-\Omega_{jk}f_{k}(x)|<\varepsilon$ (12)
-    whenever $x\in\pi^{-1}(U_{jk})\in\mathcal{N}(\mathcal{U})$.
-The family $\{f_{j}\}_{j\in J}$ is called a discrete $(\varepsilon,\delta)$-approximate local trivialization, and the set of discrete $(\varepsilon,\delta)$-approximate local trivializations for $\pi$ subordinate to an open cover $\mathcal{U}$ is denoted by $T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
-Any $\Omega$ satisfying (10) is said to be a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
-Given a discrete approximate local trivialization, we will refer to the quantity $\alpha\equiv\frac{\varepsilon}{1-\delta}$ as the trivialization quality of the trivialization and write $T_{\alpha}(\pi,\mathcal{U})=\bigcup_{\frac{\varepsilon}{1-\delta}<\alpha}T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for any $\alpha\geq 0$.
-We define the following metric on $T_{\infty}(\pi,\mathcal{U})$:
-$d_{T}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})=\sup_{j\in J}\sup_{x\in\pi^{-1}(U_{j})}|f_{j}(x)-g_{j}(x)|\\$ (11)
-Notice that we require witnesses to take values in $O(2)$ rather than $\text{Homeo}(\mathbb{S}^{1})$.
-Next, we introduce discrete and approximate relaxations of classical Cech cohomology.
-As we have seen, Cech cohomology provides a useful framework for computing obstructions and classifications, so these objects will play a central role in our classification and coordinatization algorithms.
+    whenever $x\in\pi^{-1}(U_{jk})$ and $(jk)\in\mathcal{N}(\mathcal{U})$.
+  2. 2.
+    If $d_{H}$ denotes the Hausdorff distance in $(\mathbb{C},|\cdot|)$, then $d_{H}(f_{l}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})<\delta$ for every $(jkl)\in\mathcal{N}(\mathcal{U})$.
+The set of discrete $(\varepsilon,\delta)$-approximate systems of local angle functions for $\pi$ subordinate to $\mathcal{U}$ is denoted by $\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$.
+Any $\Omega$ satisfying (12) is said to be a witness that $\mathfrak{F}\in\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$.
+Given a system of discrete approximate local circular coordinates, we will refer to the quantity
+$\frac{\varepsilon}{1-\delta}$
+as the roughness of the system and write
+$\Gamma_{\alpha}(\pi,\mathcal{U})=\bigcup\limits_{\frac{\varepsilon}{1-\delta}<\alpha}\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$
+for any $\alpha>0$. We define the following metric on $\Gamma_{\infty}(\pi,\mathcal{U})$:
+$d_{\Gamma}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})\;\;=\;\;\sup_{j\in J}\sup_{x\in\pi^{-1}(U_{j})}|f_{j}(x)-g_{j}(x)|\\$ (13)
+In practice, it will often be more convenient to work with systems of discrete approximate local circular coordinates than discrete approximate local trivializations.
+Proposition 3.11 below shows that discrete approximate local trivializations always give rise to systems of discrete approximate local circular coordinates.
+This result is directly implied by the following lemma, which we use again in Section 3.5 to compare systems of local circular coordinates induced by different choices of discrete approximate local trivializations subordinate to the same open cover.
+The proof is deferred to Appendix A.6.
+Lemma 3.10.
+Let $\pi:X\to B$ be a function between metric spaces.
+Suppose $\{\varphi_{j}:(\pi^{-1}(U_{j}),\pi,U_{j})\to(U_{j}\times\mathbb{S}^{1}_{r_{j}},p_{B},U_{j})\}_{j=1}^{2}$ are discrete $(r,\beta,\varepsilon)$-approximate local trivializations for some $r_{1},r_{2}\geq r$ and open sets $U_{1},U_{2}\subseteq B$ such that $\bar{B}_{\beta}(\pi(x))\subseteq U_{12}=U_{1}\cap U_{2}$ for some $x\in X$.
+Then,
+$|r_{1}-r_{2}|\leq\frac{2}{\pi}(\varepsilon+\beta)$
+and there exists some $\Omega_{12}\in O(2)$ such that
+$|f_{1}(x)-\Omega_{12}f_{2}(x)|\quad\leq\quad\tilde{\tau}(\varphi_{1},\varphi_{2})$
+for all $x\in\pi^{-1}(U_{12})$, where $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$ for $j=1,2$ and
+$\tau(\varphi_{1},\varphi_{2})\ =\ \frac{1}{\min\{r_{1},r_{2}\}}(\text{diam}(U_{12})+10\varepsilon+9\beta),\hskip 28.45274pt\widetilde{\tau}(\varphi_{1},\varphi_{2})\ =\ 2\sin^{-1}\left(\frac{\tau(\varphi_{1},\varphi_{2})}{2}\right)\\$ (14)
+The following result is immediate:
+Proposition 3.11.
+Suppose $\mathcal{U}$ is a trivializing cover for a discrete $(r,\beta,\varepsilon)$-approximate circle bundle $\pi:X\to B$ and $\Phi=\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times\mathbb{S}^{1}_{r_{j}}\}_{j\in J}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$.
+The functions $\{f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}\}_{j\in J}$ form a system of discrete $(\tilde{\tau}(\Phi),2\sin^{-1}(\frac{\varepsilon}{2r}))$-approximate local circular coordinates for $\pi$, where
+$\tau(\Phi)=\sup_{(jk)\in\mathcal{N}(\mathcal{U})}\tau(\varphi_{j},\varphi_{k}),\hskip 28.45274pt\tilde{\tau}(\Phi)=2\sin^{-1}\left(\frac{\tau(\Phi)}{2}\right)$ (15)
+Proof.
+Suppose $(jk)\in\mathcal{N}(\mathcal{U})$. If $j<k$, let $\Omega_{jk}\in O(2)$ be any matrix which satisfies
+$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\quad\leq\quad\widetilde{\tau}(\varphi_{j},\varphi_{k})\quad\leq\quad\widetilde{\tau}(\Phi)$
+for all $x\in\pi^{-1}(U_{jk})$ (such a matrix is guaranteed to exist by Lemma 3.10). If $j>k$, define $\Omega_{jk}=\Omega_{kj}^{-1}$ and observe
+$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\quad=\quad|\Omega_{jk}^{-1}f_{j}(x)-f_{k}(x)|\quad=\quad|\Omega_{kj}f_{j}(x)-f_{k}(x)|\quad\leq\quad\widetilde{\tau}(\varphi_{k},\varphi_{j})\quad\leq\quad\widetilde{\tau}(\Phi)$
+Lastly, define $\Omega_{jj}=\text{Id}$ for all $j\in J$.
+Evidently, the matrices $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ can be interpreted as a Čech cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$, and we have
+$|f_{j}(x)-\Omega_{jk}f_{k}(x)|\ \leq\ \widetilde{\tau}(\Phi)$
+whenever $x\in\pi^{-1}(U_{jk})$ for some $(jk)\in\mathcal{N}(\mathcal{U})$.
+Now, suppose $U_{jkl}\neq\emptyset$, and let $z\in\mathbb{S}^{1}$ be arbitrary. Let $x\in X$ be so that $\bar{B}_{\beta}(\pi(x))\subseteq U_{jkl}$ and let $b=\pi(x)$. Then,
+$d_{\mathbb{S}^{1}}^{r_{j}}(f_{j}(x),z)\quad\leq\quad d_{\infty}^{r_{j}}(\varphi_{j}(x),(b,z))\quad\leq\quad\varepsilon$
+so it follows that $d_{H}(f_{j}(\pi^{-1}(U_{jkl})),\mathbb{S}^{1})\leq\frac{\varepsilon}{r_{j}}\leq\frac{\varepsilon}{r}$ with respect to unscaled geodesic distance on $\mathbb{S}^{1}$.
+Applying identity (11), the Hausdorff bound becomes $2\sin^{-1}(\frac{\varepsilon}{2r})$, so the proposition follows.
+Remark 3.12.
+In light of the proposition above, we define the roughness of a family $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$ by
+$\alpha(\Phi)=\frac{\widetilde{\tau}(\Phi)}{1-2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)}$ (16)
+By the previous result, $\alpha(\Phi)$ is equal to the roughness of $\mathfrak{F}(\Phi)$.
+Note that our definition of discrete approximate local trivializations requires $r_{j}>\frac{16\varepsilon}{\pi}$ for each $j$, so
+$2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)\quad<\quad 2\sin^{-1}\left(\frac{\pi}{32}\right)\quad<\quad\frac{1}{5}$ (17)
+for any discrete approximate circle bundle. It follows that $\alpha(\Phi)<\frac{4}{5}\widetilde{\tau}(\Phi)$ for any family of discrete approximate local trivializations.
+3.2 Discrete Approximate Čech Cohomology
+Here, we introduce discrete and approximate relaxations of classical Čech cohomology.
+As we have seen, Čech cohomology provides a useful framework for computing obstructions and classifications of circle bundles, so these objects will play a central role in our classification and coordinatization algorithms.
 Definition 3.13.
-Given a collection of open sets $\mathcal{U}=\{U_{j}\}_{j\in J}$ and a metric group $G$, the set of $G$-valued $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
-$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})=\{\Omega\in\check{C}^{k}(\mathcal{U};\mathcal{C}_{G}):\ d_{Z}(\delta^{k}\Omega,\mathbbm{1})<\varepsilon\}$ (12)
-where $d_{Z}$ is the supremum metric on $\check{C}^{k}(\mathcal{U};\mathcal{C}_{G})$:
-$d_{Z}(\Omega,\Lambda)=\sup_{U_{\sigma}\in\mathcal{N}(\mathcal{U})^{(k)}}\sup_{b\in U_{\sigma}}d_{G}(\Omega_{\sigma}(b),\Lambda_{\sigma}(b))$ (13)
-The set of $G$-valued discrete $\varepsilon$-approximate Cech cocycles subordinate to $\mathcal{U}$ is defined by
+Given a collection $\mathcal{U}=\{U_{j}\}_{j\in J}$ of open subsets of a space $B$ and a metric group $G$, the set of $G$-valued $\varepsilon$-approximate Čech cocycles subordinate to $\mathcal{U}$ is defined by
+$\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})=\left\{\Omega\in\check{C}^{k}(\mathcal{U};\mathcal{C}_{G}):\ d_{Z}(\delta^{k}\Omega,\mathbbm{1})<\varepsilon\right\}$ (18)
+where $d_{Z}$ is the supremum metric on $\check{C}^{k+1}(\mathcal{U};\mathcal{C}_{G})$:
+$d_{Z}(\Omega,\Lambda)=\sup_{\sigma\in\mathcal{N}(\mathcal{U})^{(k+1)}}\;\sup_{b\in U_{\sigma}}d_{G}(\Omega_{\sigma}(b),\Lambda_{\sigma}(b))$ (19)
+and $\mathbbm{1}\in\check{C}^{k+1}(\mathcal{U};\mathcal{C}_{G})$ is the cochain which is constant and equal to the identity of $G$.
+The set of $G$-valued discrete $\varepsilon$-approximate Čech cocycles subordinate to $\mathcal{U}$ is defined by
 $\check{Z}^{k}_{\varepsilon}(\mathcal{U};\underline{G})=\check{Z}^{k}_{\varepsilon}(\mathcal{U};\mathcal{C}_{G})\cap\check{C}^{k}(\mathcal{U};\underline{G})$ (20)
@@
 We will primarily be interested in the case where $G=O(2)$, since every circle bundle over a paracompact space $B$ is represented by a class in $\check{H}^{1}(B;\mathcal{C}_{O(2)})$.
-Here we endow $O(2)$ with the Frobenius metric, since this is the most natural choice for a number of key results.
+Here we endow $O(2)$ with the Frobenius metric, given its predominance in algorithmic linear algebra and optimization.
+3.3 Approximate Classifying Maps
 Finally, we introduce an approximate version of classifying maps. These will play an important role in our coordinatization pipeline.
 Definition 3.14.
-Given a paracompact space $B$ and $\varepsilon\geq 0$, an $\varepsilon$-approximate classifying map is a map $f:B\to\text{Gr}(2)^{\varepsilon}$.
+Given a space $B$ and $\varepsilon\geq 0$, an $\varepsilon$-approximate classifying map is a map $f:B\to\text{Gr}(2)^{\varepsilon}$, where $\text{Gr}(2)^{0}=\text{Gr}(2)$, and $\text{Gr}(2)^{\varepsilon}\subseteq\mathbb{R}^{\infty\times\infty}$ for $\varepsilon>0$ is the $\varepsilon$-thickening
+$\text{Gr}(2)^{\varepsilon}=\{A\in\mathbb{R}^{\infty\times\infty}:\|A-P\|_{F}<\varepsilon\text{ for some }P\in\text{Gr}(2)\}\\$
 3.4 Relationships Between Notions
 In this section, we describe two key constructions which produce functions
-$\text{wit}:T_{\alpha}(\pi,\mathcal{U})\to\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})\hskip 28.45274pt\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(2)^{\sqrt{2}\varepsilon})\\$
-We first show that discrete approximate local trivializations are witnessed by discrete approximate cocycles:
+$\text{wit}:\Gamma_{\alpha}(\pi,\mathcal{U})\to\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)\hskip 28.45274pt\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}\left(B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right)\\$
+We first show that systems of discrete approximate local angle functions are witnessed by discrete approximate cocycles:
 Proposition 3.15.
-Suppose $\Omega$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$. Then $\Omega\in\check{Z}^{1}_{3\alpha\sqrt{2}}(\mathcal{U};\underline{O(2)})$.
+Suppose $\Omega$ witnesses that $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha>0$. Then $\Omega\in\check{Z}^{1}_{3\alpha\sqrt{2}}\left(\mathcal{U};\underline{O(2)}\right)$.
 Proof.
-Assume $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$, where $\alpha=\frac{\varepsilon}{1-\delta}$.
-Suppose $x\in\pi^{-1}(U_{j}\cap U_{k}\cap U_{l})$.
-Let $v_{\tau\tau^{\prime}}=\Omega_{\tau\tau^{\prime}}f_{\tau^{\prime}}(x)-f_{\tau}(x)$ for $\tau,\tau^{\prime}\in\{j,k,l\}$.
+Assume $\mathfrak{F}=\{f_{j}\}_{j\in J}\in\Gamma_{\varepsilon,\delta}(\pi,\mathcal{U})$, where $\alpha=\frac{\varepsilon}{1-\delta}$.
+Suppose $x\in\pi^{-1}(U_{jkl})$ and let $v_{st}=\Omega_{st}f_{t}(x)-f_{s}(x)$ for $s,t\in\{j,k,l\}$.
 Then,
-$\left|\Omega_{jk}\Omega_{kl}f_{l}(x)-\Omega_{jl}f_{l}(x)\right|=\left|\Omega_{jk}(v_{kl}+f_{k}(x))-\Omega_{jl}f_{l}(x)\right|$
-$\left|\Omega_{jk}v_{kl}+(f_{j}(x)+v_{jk})-(f_{j}(x)-v_{jl})\right|=\left|\Omega_{jk}v_{kl}+v_{jk}-v_{jl}\right|$
-$\leq|v_{kl}|+|v_{jk}|+|v_{jl}|<3\varepsilon\\$
-The result then follows immediately from Proposition A.10.
-Next, we show how to algorithmically compute a witness for a discrete approximate circle bundle.
-Construction 3.5.
-Let $X=\{x_{s}\}_{s=1}^{n}$ be a finite set, and let $f,g:S\to\mathbb{S}^{1}$.
-Compute the values $\{\alpha_{s}=\cos^{-1}(f(x_{s}))\}^{n}_{s=1}$ and $\{\beta_{s}=\cos^{-1}(g(x_{s}))\}^{n}_{s=1}$, then define $\gamma_{s}=\alpha_{s}-\beta_{s}$ and $\tilde{\gamma}_{s}=\alpha_{s}+\beta_{s}$ for all $s$.
-Find the respective midpoints $\theta$ and $\widetilde{\theta}$ of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ and the shortest arc $\widetilde{C}$ containing $\{\widetilde{\gamma}_{s}\}_{s=1}^{n}$.
-Compute the values
-$d_{\text{rot}}=\text{max}_{s}|f(x_{s})-R_{\theta}g(x_{s})|,\hskip 28.45274ptd_{\text{ref}}=\text{max}_{s}|f(x_{s})-R_{\widetilde{\theta}}g(x_{s})|$
-If $d_{\text{rot}}<d_{\text{ref}}$, set $\Omega=R_{\theta}$; otherwise, set $\Omega=R_{\widetilde{\theta}}r$, where $R_{\theta}=\left(\begin{array}[]{cc}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\\ \end{array}\right)$ and $r=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$.
+$\left|\Omega_{jk}\Omega_{kl}f_{l}(x)-\Omega_{jl}f_{l}(x)\right|\quad=\quad\left|\Omega_{jk}(v_{kl}+f_{k}(x))-\Omega_{jl}f_{l}(x)\right|$
+$\quad\left|\Omega_{jk}v_{kl}+(f_{j}(x)+v_{jk})-(f_{j}(x)-v_{jl})\right|\quad=\quad\left|\Omega_{jk}v_{kl}+v_{jk}-v_{jl}\right|$
+$\leq\quad|v_{kl}|+|v_{jk}|+|v_{jl}|\quad<\quad 3\varepsilon\\$
+The result then follows from applying Proposition A.12 to each matrix $(\Omega_{jk}\Omega_{kl}-\Omega_{jl})$.
+Below we describe an explicit construction which produces a witness for a system of discrete approximate local angle functions.
+Construction 3.16.
+Let $X$ be any nonempty set, and let $f,g:X\to\mathbb{S}^{1}$.
+For each $x\in X$, define $e^{i\alpha_{x}}=f(x)$ and $e^{i\beta_{x}}=g(x)$, then define $\gamma_{x}=\alpha_{x}-\beta_{x}$ and $\tilde{\gamma}_{x}=\alpha_{x}+\beta_{x}$.
+Let $e^{i\theta}$ denote the midpoint of the shortest arc $C$ in $\mathbb{S}^{1}$ containing $\{e^{i\gamma_{x}}\}_{x\in X}$, and let $e^{i\widetilde{\theta}}$ denote the midpoint of the shortest arc $\widetilde{C}$ containing $\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X}$.
+Define
+$d_{\text{rot}}=\sup_{x\in X}|f(x)-R_{\theta}g(x)|,\hskip 28.45274ptd_{\text{ref}}=\sup_{x\in X}|f(x)-R_{\widetilde{\theta}}g(x)|$
+If $d_{\text{rot}}<d_{\text{ref}}$, set $\Omega^{*}=R_{\theta}$; otherwise, set $\Omega^{*}=R_{\widetilde{\theta}}r$, where $R_{\theta}=\left(\begin{array}[]{cc}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\\ \end{array}\right)$ and $r=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$.
 Proposition 3.17.
-If $n>1$ and $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ in Construction 3.5 above, then $\Omega\in O(2)$ is the unique minimizer of $\text{max}_{s}|f(x_{s})-\Omega g(x_{s})|$.
+[…]
 Proof.
-Let $L:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L(\Omega)=\max_{s}|f(x_{s})-\Omega g(x_{s})|^{2}$. We have
-$|f(x_{s})-\Omega g(x_{s})|^{2}$ $=|f(x_{s})|^{2}+|g(x_{2})|^{2}-2f(x_{s})\cdot\Omega g(x_{s})$
-$=2-2f(x_{s})\cdot\Omega g(x_{s})$
-for all $s$.
-If $\Omega=R_{\theta}$ is a counter-clockwise rotation by $\theta$, then $f(x_{s})\cdot\Omega g(x_{s})=\cos(\gamma_{s}-\theta)$, so to minimize $L$ over $SO(2)$ we need to identify $\theta\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\gamma_{s}$ is minimized.
-If the measure of $C$ is less than $\pi$, then this is precisely achieved at the midpoint of the shortest arc $C$ containing $\{\gamma_{s}\}_{s=1}^{n}$ (furthermore, the requirement that $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$ guarantees that $C$ is unique).
-To see this, first note that the optimal $\theta$ certainly lies in $C$, and for fixed $\theta\in C$, $\cos(\gamma_{s}-\theta)$ is smallest when $\gamma_{s}$ is an endpoint of $C$.
-By symmetry, the $\theta$ which maximizes the minimum of the values $\{\cos(\gamma_{s}-\theta)\}_{s=1}^{n}$ must be half way between the two end points (otherwise, a small perturbation of $\theta$ towards the midpoint would move $\theta$ closer to the more distant endpoint while still keeping $\theta$ nearer to the closer endpoint.
+Let $L:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L(\Omega)=\sup\limits_{x\in X}|f(x)-\Omega g(x)|^{2}$. We have
+$|f(x)-\Omega g(x)|^{2}\quad$ $=\quad|f(x)|^{2}+|g(x)|^{2}-2f(x)\cdot\Omega g(x)$
+$=\quad 2-2f(x)\cdot\Omega g(x)$
+for all $x$ (here we are identifying $\mathbb{R}^{2}$ with $\mathbb{C}$ in the usual way).
+If $\Omega=R_{\theta}$ is a counter-clockwise rotation by $\theta$, then $f(x)\cdot\Omega g(x)=\cos(\gamma_{x}-\theta)$, so to minimize $L$ over $SO(2)$ we need to identify $\theta\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\gamma_{x}$ is minimized.
+If the measure of $C$ is less than $\pi$, then this is precisely achieved at the midpoint of the shortest arc $C$ containing $\{e^{i\gamma_{x}}\}_{x\in X}$ (furthermore, the requirement that $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})<\pi$ guarantees that $C$ is unique).
+To see this, first note that $e^{i\theta}$ certainly lies in $C$ for any optimal $\theta$, and for any fixed $e^{i\theta}\in C$, $d_{\mathbb{S}^{1}}(e^{i\gamma_{x}},e^{i\theta})$ is smallest when $e^{i\gamma_{x}}$ is an endpoint of $C$.
+By symmetry, the $\theta$ which maximizes the minimum of the values $\{d_{\mathbb{S}^{1}}(e^{i\gamma_{x}},e^{i\theta})\}_{x\in X}$ must be half way between the two end points (otherwise, a small perturbation of $\theta$ towards the midpoint would move $\theta$ closer to the more distant endpoint while still keeping $\theta$ nearer to the closer endpoint.
 This would yield a more optimal choice for $\theta$).
-Analogously, if $\Omega=R_{\widetilde{\theta}}r$ for some $\widetilde{\theta}\in\mathbb{R}$, one can check that $f(x_{s})\cdot\Omega g(x_{s})=\cos(\widetilde{\gamma}_{s}-\theta)$.
-Thus, to minimize $L$ over the component of $O(2)$ with matrices of negative determinant, we need to identify $\widetilde{\theta}\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\widetilde{\gamma}_{s}$ is minimized.
+Analogously, if $\Omega=R_{\widetilde{\theta}}r$ for some $\widetilde{\theta}\in\mathbb{R}$, one can check that $f(x)\cdot\Omega g(x)=\cos(\widetilde{\gamma}_{x}-\theta)$.
+Thus, to minimize $L$ over the component of $O(2)$ with matrices of negative determinant, we need to identify $\widetilde{\theta}\in\mathbb{S}^{1}$ for which the angular distance to the farthest $\widetilde{\gamma}_{x}$ is minimized.
 Finally, comparing $d_{\text{rot}}$ and $d_{\text{ref}}$ determines which component of $O(2)$ contains the optimal $\Omega$.
-To see that these two values cannot be equal, observe that if $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})>\pi$.
+To see that these two values cannot be equal, observe that if $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})<\pi$, then $\text{diam}(\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X})>\pi$.
 It follows that then $d_{\text{rot}}<\frac{\pi}{2}$ and $d_{\text{ref}}>\frac{\pi}{2}$.
-Similarly, if $\text{diam}(\{\widetilde{\gamma}_{s}\}_{s=1}^{n})<\pi$, then $\text{diam}(\{\gamma_{s}\}_{s=1}^{n})>\pi$, so $d_{\text{ref}}<d_{\text{rot}}$.
+Similarly, if $\text{diam}(\{e^{i\widetilde{\gamma}_{x}}\}_{x\in X})<\pi$, then $\text{diam}(\{e^{i\gamma_{x}}\}_{x\in X})>\pi$, so $d_{\text{ref}}<d_{\text{rot}}$.
 Corollary 3.18.
-If $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for $\varepsilon<\sqrt{2}$ and $|\pi^{-1}(U_{jk})|>1$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, then Construction 3.5 yields a witness.
+If $\mathfrak{F}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$ for $\varepsilon<\sqrt{2}$ and $|\pi^{-1}(U_{jk})|>1$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, then Construction 3.16 yields a witness.
 Proof.
-The hypotheses imply
+Let $\mathfrak{F}=\{f_{j}\}_{j\in J}$. The hypotheses imply
 $|f_{j}(x)-\Omega_{jk}f_{k}(x)|^{2}<\varepsilon^{2}$
 or, equivalently,
-$\gamma^{jk}_{s}=(f_{j}(x_{s}))\cdot(\Omega_{jk}f_{k}(x_{s}))>1-\frac{\varepsilon^{2}}{2}$
-for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$.
-Then $|\gamma^{jk}_{s}|>\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$ for all $x_{s}\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $\text{diam}(\{\gamma^{jk}_{s}\})<2\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$.
-We conclude that if $\cos^{-1}(1-\frac{\varepsilon^{2}}{2})<\frac{\pi}{2}$ (or, equivalently, if $\varepsilon<\sqrt{2}$), then each $\Omega_{jk}$ obtained using Construction 3.5 is the unique minimizer of $\text{max}_{x_{s}\in\pi^{-1}(U_{j}\cap U_{k})}|f_{j}(x_{s})-f_{k}(x_{s})|$.
-Thus by Proposition 3.4, $\Omega$ is a witness that $\{f_{j}\}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
+$\gamma^{jk}_{x}=(f_{j}(x))\cdot(\Omega_{jk}f_{k}(x))>1-\frac{\varepsilon^{2}}{2}$
+for all $x\in U_{jk}$ and $(jk)\in\mathcal{N}(\mathcal{U})$.
+Then $|\gamma^{jk}_{x}|>\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$ for all $x\in U_{jk}$, so $\text{diam}(\{\gamma^{jk}_{x}\})<2\cos^{-1}(1-\frac{\varepsilon^{2}}{2})$.
+We conclude that if $\cos^{-1}(1-\frac{\varepsilon^{2}}{2})<\frac{\pi}{2}$ (or, equivalently, if $\varepsilon<\sqrt{2}$), then each $\Omega_{jk}$ obtained using Construction 3.16 is the unique minimizer of $\max\limits_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-f_{k}(x)|$.
+Thus by Proposition 3.15, $\Omega$ is a witness that $\mathfrak{F}\in T_{\varepsilon,\delta}(\pi,\mathcal{U})$.
 We state the following technical result for later reference:
 Proposition 3.19.
-Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\}\in T_{\varepsilon_{f},\infty}(\pi,\mathcal{U})$ and $\{g_{j}\}\in T_{\varepsilon_{g},\infty}(\pi,\mathcal{U})$, and $d_{h}(f_{k}(\pi^{-1}(U_{jk})),\mathbb{S}^{1})<\delta_{1}$ (with respect to the Euclidean metric) for every ordered 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$.
-If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta_{2}$, then $d_{Z}(\Omega,\Lambda)<\frac{\sqrt{2}}{1-\delta_{1}}(\varepsilon_{f}+\varepsilon_{g}+2\delta_{2})$.
+Let $\delta<1$.
+Suppose $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witness $\{f_{j}\}_{j\in J}\in\Gamma_{\varepsilon_{f},\delta}(\pi,\mathcal{U})$ and $\{g_{j}\}_{j\in J}\in\Gamma_{\varepsilon_{g},\delta}(\pi,\mathcal{U})$, respectively, and $d_{H}(f_{k}(\pi^{-1}(U_{jk})),\mathbb{S}^{1})<\delta_{1}$ in $(\mathbb{C},|\cdot|)$ for every ordered 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$.
+If $d_{\Gamma}(\{f_{j}\}_{j\in J},\{g_{j}\}_{j\in J})<\delta_{2}$, then $d_{Z}(\Omega,\Lambda)<\frac{\sqrt{2}}{1-\delta_{1}}(\varepsilon_{f}+\varepsilon_{g}+2\delta_{2})$.
 Proof.
-Given any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, the triangle inequality implies
-$|\Omega_{jk}f_{k}(x)-\Lambda_{jk}f_{k}(x)|\leq$
+Given any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{jk})$, the triangle inequality implies
+$|\Omega_{jk}f_{k}(x)-\Lambda_{jk}f_{k}(x)|\quad\leq$
 $|\Omega_{jk}f_{k}(x)-f_{j}(x)|\ +\ |f_{j}(x)-g_{j}(x)|\ +\ |g_{j}(x)-\Lambda_{jk}g_{k}(x)|\ +\ |\Lambda_{jk}g_{k}(x)-\Lambda_{jk}f_{k}(x)|$
@@
 Corollary 3.20.
-If $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ both witness $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ then $d_{Z}(\Omega,\Lambda)<(2\sqrt{2})\alpha$.
+If $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ both witness $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}\,\alpha$.
 Corollary 3.21.
-Suppose $\Omega,\Lambda\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$.
-If $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}(\alpha+\delta)$.
-The following construction and subsequent result are due to [SP23]:
-Construction 3.11.
-Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some locally-finite open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of a paracompact space $B$.
-Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$.
-For each $j$, define $\Phi_{j}:U_{j}\to V(2)$ by
-$\Phi_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho_{0}(b)}\Omega_{0j}(b)\\ \sqrt{\rho_{1}(b)}\Omega_{1j}(b)\\ \vdots\\ \end{array}\right)$ (15)
+Suppose $\Omega,\Lambda\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witness $\mathfrak{F},\mathfrak{G}\in T_{\alpha}(\pi,\mathcal{U})$, respectively.
+If $d_{\Gamma}(\mathfrak{F},\mathfrak{G})<\delta$, then $d_{Z}(\Omega,\Lambda)<2\sqrt{2}(\alpha+\delta)$.
+The following construction and subsequent result are due to [24, Construction 4.15]:
+Construction 3.22.
+Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some locally-finite open cover $\mathcal{U}=\{U_{j}\}_{j\in\mathbb{N}}$ of a paracompact space $B$.
+Choose a partition of unity $\{\rho_{j}\}_{j\in\mathbb{N}}$ subordinate to $\mathcal{U}$.
+For each $j\in\mathbb{N}=\{0,1,\ldots\}$, define $\Psi_{j}:U_{j}\to V(2)$ by
+$\Psi_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho_{0}(b)}\Omega_{0j}(b)\\ \sqrt{\rho_{1}(b)}\Omega_{1j}(b)\\ \vdots\\ \end{array}\right)$ (21)
 Note that when $b\in U_{j}\cap U_{k}^{c}$, $\rho_{k}(b)=0$, so this map is well-defined. Let $f:B\to\mathbb{R}^{\infty\times\infty}$ be defined by
-$\widetilde{f}(b)=\sum_{j\in J}\rho_{j}(b)\Phi_{j}(b)\Phi_{j}(b)^{T}$ (16)
-Since each $b$ is contained in a finite number of sets in $\mathcal{U}$, this map is well-defined.
-For a proof of the proposition below, see [SP23]:
+$\widetilde{f}(b)=\sum_{j=0}^{\infty}\rho_{j}(b)\Psi_{j}(b)\Psi_{j}(b)^{T}$ (22)
+Since each $b$ is contained in a finite number of sets in $\mathcal{U}$, the sum above is finite and hence $\tilde{f}$ is well-defined.
+For a proof of the theorem below, see [24, Lemmas 4.16, 4.17, 4.18 and Theorem 4.21]:
 Theorem 3.23.
-For $\varepsilon<\frac{1}{2}$ and any partition of unity $\{\rho_{j}\}_{j\in J}$, Construction 3.11 yields a map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(d)^{\sqrt{2}\varepsilon})$ with the following properties:
+For $\varepsilon\geq 0$ and any partition of unity $\{\rho_{j}\}_{j\in\mathbb{N}}$, Construction 3.22 yields a map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\text{Maps}(B,\text{Gr}(d)^{\sqrt{2}\varepsilon})$ with the following properties:
   1. 1.
-    The induced map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to[B,\text{Gr}(d)^{\sqrt{2}\varepsilon}]$ factors through $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ and does not depend on choice of $\{\rho_{j}\}_{j\in J}$.
+    The induced map $\text{cl}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\left[B,\text{Gr}(d)^{\sqrt{2}\varepsilon}\right]$ factors through $\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ and does not depend on the choice of $\{\rho_{j}\}_{j\in\mathbb{N}}$.
   2. 2.
-    If $\Omega,\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Lambda)<\delta$, then $\text{cl}(\Omega)$ and $\text{cl}(\Lambda)$ become equal in $B,\text{Gr}(d)^{\sqrt{2}(\delta+\varepsilon)}]$.
+    If $\Omega,\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ satisfy $d_{Z}(\Omega,\Lambda)<\delta$, then $\text{cl}(\Omega)$ and $\text{cl}(\Lambda)$ become equal in $\left[B,\text{Gr}(d)^{\sqrt{2}(\delta+\varepsilon)}\right]$.
 3.5 Relationship To True Circle Bundles
-In this section, we show that for sufficiently small $\alpha$, any discrete approximate local trivialization $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ corresponds to a true circle bundle.
-We build on previous results of [Tin22] and [SP23] which provide conditions under which approximate $O(2)$-cocycles and classifying maps can be uniquely identified with true ones.
-The following construction and subsequent result, from [Tin22], describe a procedure for obtaining a classifying map from an approximate one:
-Construction 3.13.
-Given $A\in\mathbb{R}^{n\times n}$, compute the symmetrization $A_{s}=\frac{A+A^{T}}{2}$, then compute an orthogonal diagonalization $\Lambda D\Lambda^{T}$ of $A_{s}$ such that the eigenvalues appear in decreasing order along the diagonal of $D$.
+In this section, we show that for sufficiently small $\alpha>0$, any system of discrete approximate local angle functions $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ uniquely corresponds to a true circle bundle.
+From this we deduce conditions under which a discrete approximate circle bundle can be identified with an isomorphism class of classical circle bundles.
+We build on previous results of [30] and [24] which provide conditions under which approximate $O(2)$-cocycles and classifying maps can be uniquely identified with true ones.
+The following construction and subsequent result [30, Lemma 1] describe a procedure for obtaining a classifying map from an approximate one:
+Construction 3.24.
+Given $A\in\mathbb{R}^{n\times n}$, compute its symmetrization $A_{s}=\frac{A+A^{T}}{2}$, and an orthogonal diagonalization $A_{s}=\Lambda D\Lambda^{T}$ such that the eigenvalues of $A_{s}$ appear in decreasing order along the diagonal of $D$.
 Finally, set $\Pi(A)=\Lambda J_{2}\Lambda^{T}$, where $J_{2}$ is the $n\times n$ matrix whose only non-zero entries are the first two diagonal entries, both equal to 1.
@@
 Proposition 3.25.
-Let $\Pi:\mathbb{R}^{\infty\times\infty}\to\text{Gr}(2)$ be the function described in Construction 3.13.
-For $\varepsilon\leq\frac{\sqrt{2}}{2}$, the restriction of $\Pi$ to $\text{Gr}(2)^{\varepsilon}$ is a retraction map.
-Furthermore, the induced map $\Pi_{*}:\text{Maps}(B,\text{Gr}(2)^{\varepsilon})\to\text{Maps}(B,\text{Gr}(2))$ descends to a bijection $\Pi_{*}:[B,\text{Gr}(2)^{\varepsilon}]\to[B,\text{Gr}(2)]$.
+For $\varepsilon\leq\frac{\sqrt{2}}{2}$, Construction 3.24 yields a deformation retraction $\Pi:\text{Gr}(2)^{\varepsilon}\to\text{Gr}(2)$.
+Hence, the induced map $\Pi_{*}:\text{Maps}(B,\text{Gr}(2)^{\varepsilon})\to\text{Maps}(B,\text{Gr}(2))$ descends to a natural bijection $\Pi_{*}:[B,\text{Gr}(2)^{\varepsilon}]\to[B,\text{Gr}(2)]$.
 For later use, we also record the following additional construction and result from the same paper:
-Construction 3.15.
+Construction 3.26.
 Let $S_{n}=\{(P,A)\in\text{Gr}(2,n)\times V(2,n):\text{rank}(PA)=2\}$.
@@
 Proposition 3.27.
-Construction 3.15 defines a map $\widehat{\Pi}_{n}:S_{n}\to V(2,n)$ for each $n$, thereby inducing a map $\widehat{\Pi}:S\to V(2)$ (where $S$ is the direct limit of the sequence $S_{1}\subset S_{2}\subset\cdots$).
-Furthermore, $\widehat{\Pi}(P,A)$ is the closest matrix to $A$ in $\pi_{Gr}^{-1}(P)$ with respect to the Frobenius metric.
-Finally, we have the following construction and subsequent result from [SP23]:
-Construction 3.17.
-Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$.
-Let $\{\Phi_{j}:U_{j}\to V(2)\}_{j\in J}$ be the maps defined by (15), and let $\Pi$ be as described in Construction 3.13.
-Define $f_{\Omega}=\Pi_{*}\circ\text{cl}(\Omega)$, then for each $j$, define $\widetilde{\Phi}_{j}:U_{j}\to V(2)$ by $\widetilde{\Phi}_{j}(b)=\hat{\Pi}(f_{\Omega}(b),\Phi_{j}(b))$.
-Finally, define $\widetilde{\Omega}\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ by $\widetilde{\Omega}_{jk}(b)=\widetilde{\Phi}_{j}^{T}(b)\widetilde{\Phi}_{k}(b)$.
+Construction 3.26 defines a map $\widehat{\Pi}_{n}:S_{n}\to V(2,n)$ for each $n$, thereby inducing a map $\widehat{\Pi}:S\to V(2)$ (where $S$ is the direct limit of the sequence $S_{1}\subset S_{2}\subset\cdots$).
+Furthermore, $\widehat{\Pi}(P,A)$ is the point in $\pi_{Gr}^{-1}(P)$ closest to $A$ with respect to the Frobenius metric.
+Finally, we have the following construction and subsequent result [24, Theorem 4.27]:
+Construction 3.28.
+Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for $\varepsilon<\frac{1}{2}$.
+Let $\{\Psi_{j}:U_{j}\to V(2)\}_{j\in\mathbb{N}}$ be the maps defined by (21), and let $\Pi$ be as described in Construction 3.24.
+Define $f_{\Omega}=\Pi_{*}\circ\text{cl}(\Omega)$, and for each $j$ define $\widetilde{\Psi}_{j}:U_{j}\to V(2)$ by $\widetilde{\Psi}_{j}(b)=\hat{\Pi}(f_{\Omega}(b),\Psi_{j}(b))$.
+Finally, define $\widetilde{\Omega}\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ by $\widetilde{\Omega}_{jk}(b)=\widetilde{\Psi}_{j}(b)^{T}\widetilde{\Psi}_{k}(b)$.
 Proposition 3.29.
-For $\varepsilon<\frac{1}{2}$, Construction 3.17 yields a map $\Pi_{Z}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that the following diagram commutes:
+For $\varepsilon<\frac{1}{2}$, Construction 3.28 yields a map $\Pi_{Z}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})\to\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that the following diagram commutes:
 […]
@@
 If $\varepsilon<\frac{\sqrt{2}}{2}$, any $\varepsilon$-approximate classifying map uniquely determines a true classifying map, hence can be identified with an isomorphism class of true circle bundles.
-Similarly, if $\varepsilon<\frac{1}{2}$, any $\varepsilon$-approximate $O(2)$ Cech 1-cocycle $\Omega$ uniquely determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
+Similarly, if $\varepsilon<\frac{1}{2}$, any $\varepsilon$-approximate $O(2)$ Čech 1-cocycle $\Omega$ uniquely determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
 Proposition 3.31.
-Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$.
+Suppose $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$.
 Then any witness determines a true circle bundle, and the circle bundles determined by any two witnesses coincide.
-In this sense, $\{f_{j}\}$ uniquely determines a true circle bundle.
+In this sense, $\mathfrak{F}$ uniquely determines a true circle bundle.
 Proof.
-Any witness $\Omega$ lies in $\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ by Proposition 3.4.
+Any witness $\Omega$ lies in $\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$ by Proposition 3.15.
 Thus, if $\alpha<\frac{1}{6\sqrt{2}}$, then $(3\sqrt{2})\alpha<\frac{1}{2}$, so by Proposition 3.30 $\Omega$ determines a true circle bundle classified by $\Pi_{*}\circ\text{cl}(\Omega)$.
-By Corollary 3.9, any two witnesses $\Omega,\Omega^{\prime}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$ satisfy $d_{Z}(\Omega,\Omega^{\prime})<2\varepsilon\sqrt{2}$.
+By Corollary 3.20, any two witnesses $\Omega,\Omega^{\prime}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$ satisfy $d_{Z}(\Omega,\Omega^{\prime})<2\varepsilon\sqrt{2}$.
 Thus, by Proposition 3.23, $\Pi_{*}\circ\text{cl}(\Omega)=\Pi_{*}\circ\text{cl}(\Omega^{\prime})$ as long as $\sqrt{2}(3\varepsilon\sqrt{2}+2\varepsilon\sqrt{2})=10\varepsilon\leq\frac{\sqrt{2}}{2}$, or equivalently, $\varepsilon\leq\frac{\sqrt{2}}{20}$.
@@
 Proposition 3.32.
-Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{T}(\{f_{j}\},\{g_{j}\})<\delta$.
-If $5\alpha+2\delta<\frac{\sqrt{2}}{4}$, then $\{f_{j}\}$ and $\{g_{j}\}$ both determine the same true circle bundle.
+Suppose $\mathfrak{F},\mathfrak{G}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ and $d_{\Gamma}(\mathfrak{F},\mathfrak{G})<\delta$.
+If $5\alpha+2\delta<\frac{\sqrt{2}}{4}$, then $\mathfrak{F}$ and $\mathfrak{G}$ both determine the same true circle bundle.
 Proof.
-By Proposition 3.10, any two witnesses $\Omega_{1},\Omega_{2}\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ that $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ satisfy $d_{Z}(\Omega_{1},\Omega_{2})<2\sqrt{2}(\alpha+\delta)$.
-By Proposition 3.4, we have $\Omega_{1},\Omega_{2}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, so Proposition 3.18 implies $\text{cl}(\Omega_{1})]$ and $\text{cl}(\Omega_{2})]$ coincide in $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$, where $\varepsilon=(3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)$.
-Since $\Pi$ induces a bijection from $B,\text{Gr}(2)^{\sqrt{2}\varepsilon}]$ to $B,\text{Gr}(2)]$ by Proposition 3.19, $\Omega_{1}$ and $\Omega_{2}$ determine the same true circle bundle as long as
+By Corollary 3.21, any two witnesses $\Omega_{1},\Omega_{2}\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ for $\mathfrak{F},\mathfrak{G}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ satisfy $d_{Z}(\Omega_{1},\Omega_{2})<2\sqrt{2}(\alpha+\delta)$.
+By Proposition 3.15, we have $\Omega_{1},\Omega_{2}\in\check{Z}^{1}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$, so Proposition 3.29 implies $\text{cl}(\Omega_{1})]$ and $\text{cl}(\Omega_{2})]$ coincide in $\left[B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right]$, where $\varepsilon=(3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)$.
+Since $\Pi$ induces a bijection from $\left[B,\text{Gr}(2)^{\sqrt{2}\varepsilon}\right]$ to $B,\text{Gr}(2)]$ by Proposition 3.30, $\Omega_{1}$ and $\Omega_{2}$ determine the same true circle bundle as long as
 $\sqrt{2}\left((3\sqrt{2})\alpha+2\sqrt{2}(\alpha+\delta)\right)<\frac{\sqrt{2}}{2}$
@@
 Corollary 3.33.
-Suppose $\{f_{j}\},\{g_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ are homotopic through $T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{20}$.
-Then $\{f_{j}\}$ and $\{g_{j}\}$ determine the same true circle bundle.
+Suppose $\mathfrak{F}=\{f_{j}\}_{j\in J},\mathfrak{G}=\{g_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha<\frac{\sqrt{2}}{20}$, and that there are homotopies $\{h_{j}:U_{j}\times I\to\mathbb{S}^{1}\}_{j\in J}$ such that $h_{j}(\cdot,0)=f_{j}$, $h_{1}(\cdot,1)=g_{j}$ and $\{h_{j}(\cdot,t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for all $t\in I$.
+Then $\mathfrak{F}$ and $\mathfrak{G}$ determine the same true circle bundle.
 Proof.
-Suppose $\{h_{j}:\pi^{-1}(U_{j})\times I\to\mathbb{S}^{1}\}$ are continuous maps such that $\{h_{j}(\cdot,t)\}\in T_{\alpha}(\pi,\mathcal{U})$ for all $t$, and $h_{j}(x,0)=f_{j}(x)$ and $h_{j}(x,1)=g_{j}(x)$.
-Proposition 3.21 implies that for any $\alpha<\frac{\sqrt{2}}{20}$, there exists some $\delta>0$ such that $\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\}\in T_{\alpha}(\pi,\mathcal{U})$ determine the same true bundle as long as $d_{T}(\{h_{j}(\cdot,t)\},\{h_{j}(\cdot,t+\Delta t)\})<\delta$.
-By continuity, it follows that $\{h_{j}(\cdot,t)\}$ determines the same true bundle for all $t\in[0,1]$, and in particular that the true bundles determined by $\{f_{j}\}$ and $\{g_{j}\}$ coincide.
-Next, we provide a construction for producing a family $\{\widetilde{f}_{j}\}\in T_{0}(\pi,\mathcal{U})$ from a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ when $\alpha$ is sufficiently small.
-Construction 3.23.
-Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$.
-Let $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ be any witness, and let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
+Suppose $\{h_{j}:\pi^{-1}(U_{j})\times I\to\mathbb{S}^{1}\}_{j\in J}$ are continuous maps such that $\{h_{j}(\cdot,t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for all $t$, and $h_{j}(x,0)=f_{j}(x)$ and $h_{j}(x,1)=g_{j}(x)$.
+Proposition 3.32 implies that for any $\alpha<\frac{\sqrt{2}}{20}$, there exists some $\delta>0$ such that for any $\Delta t>0$, $\{h_{j}(\cdot,t)\}_{j\in J},\{h_{j}(\cdot,t+\Delta t)\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ determine the same true bundle as long as $d_{\Gamma}(\{h_{j}(\cdot,t)\}_{j\in J},\{h_{j}(\cdot,t+\Delta t)\}_{j\in J})<\delta$.
+By continuity, it follows that $\{h_{j}(\cdot,t)\}_{j\in J}$ determines the same true bundle for all $t\in[0,1]$, and in particular that the true bundles determined by $\{f_{j}\}_{j\in J}$ and $\{g_{j}\}_{j\in J}$ coincide.
+Remark 3.34.
+Discrete approximate local angle functions satisfying the hypotheses of Corollary 3.33 above for some $\alpha$ are said to be homotopic through $\Gamma_{\alpha}(\pi,\mathcal{U})$.
+Given a function $\pi:X\to B$ between metric spaces and an open cover $\mathcal{U}=\{U_{j}\}$ of $B$, we define $\Gamma_{0}(\pi,\mathcal{U})$ to be the set of families $\{f_{j}:U_{j}\to\mathbb{S}^{1}\}_{j\in J}$ for which there exists some Čech cocycle $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ satisfying $f_{j}(b)=\Omega_{jk}(b)f_{k}(b)$ for all $U_{jk}\in\mathcal{N}(\mathcal{U})$ and $b\in U_{jk}$.
+Note that $\Omega$ need not be locally constant.
+We provide a construction for producing a family $\{\widetilde{f}_{j}\}_{j\in J}\in\Gamma_{0}(\pi,\mathcal{U})$ from a family $\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ when $\alpha$ is sufficiently small and $\mathcal{U}=\{U_{j}\}_{j\in J}$ is locally finite.
+Our construction uses the notion of a Frechet mean [13] for angular values in $\mathbb{S}^{1}$, which generalizes the notion of a centroid in Euclidean space:
+Definition 3.35.
+Let $M$ be a complete metric space, and let $m_{1},...,m_{N}\in M$.
+For any $p\in M$, the (weighted) Frechet variance of the set $\{m_{1},...,m_{N}\}$ at $p$ with weights $\{w_{1},...,w_{N}\}$ is defined by
+$F_{V}(p)=\sum_{n=1}^{N}w_{n}d_{M}(p,x_{n})^{2}$
+The point $p$ is called a (weighted) Karcher mean of $\{m_{1},...,m_{N}\}$ if it minimizes $F_{V}$ on $M$.
+If $p$ is the unique weighted Karcher mean of $\{m_{1},...,m_{N}\}$, it is called the (weighted) Frechet mean of $\{m_{1},...,m_{N}\}$.
+Construction 3.36.
+Suppose $\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some locally-finite good open cover $\mathcal{U}$.
+Let $\Omega\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ be any witness, and let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
 Choose a partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$, and for each $j$, define $\widetilde{f}_{j}:\pi^{-1}(U_{j})\to\mathbb{S}^{1}$ by
-$\widetilde{f}_{j}(x)=\exp\left(\sum_{k\in J}\rho_{k}(\pi(x))\log\left(\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\right)\right)$ (17)
-Note that if $\pi(x)\notin U_{j}\cap U_{k}$ then $\rho_{k}(\pi(x))=0$, so the maps are well-defined.
+$\widetilde{f}_{j}(x)\quad=\quad\exp\left(\sum_{k\in J}\rho_{k}(\pi(x))\log\left(\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\right)\right)$ (23)
+Here we fix a branch of log for each $x\in X$ and each $j\in J$ as follows.
+Since $\mathcal{U}$ is locally finite, then the set $\left\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\in\mathbb{S}^{1}\,:\,k\in J\right\}$ is finite, and thus is contained in the domain of some branch of log.
+Note that if $\pi(x)\notin U_{jk}$ then $\rho_{k}(\pi(x))=0$, so $\tilde{f}_{j}(x)$ is well-defined.
 Proposition 3.37.
-Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
-Then the maps $\{\widetilde{f}_{j}\}$ described in Construction 3.23 satisfy $\widetilde{f}_{j}(x)=\widetilde{\Omega}_{jk}(\pi(x))\widetilde{f}_{k}(x)$ for all $x\in\pi^{-1}(U_{j}\cap U_{k})$ and $U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$.
+Suppose $\Omega$ witnesses that $\mathfrak{F}=\{f_{j}\}_{j\in J}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some good open cover $\mathcal{U}$.
+Then the functions $\{\widetilde{f}_{j}\}_{j\in J}$ described in Construction 3.36 form a system $\widetilde{\mathfrak{F}}\in\Gamma_{0}(\pi,\mathcal{U})$.
 Proof.
-Let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$, and assume $\Omega$ decomposes as $\Omega_{jk}(b)=\exp(\Theta_{jk}(b))r(\omega_{jk})$ for some $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$, where $\exp:\mathbb{R}\to SO(2)$ is as defined in (9) and $r:\mathbb{Z}_{2}\to O(2)$ is given by $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$.
-Then, for any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
+Let $\exp:\mathbb{R}\to SO(2)$ be as defined in (8) and define $r:\mathbb{Z}_{2}\to O(2)$ by $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$.
+Let $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
+Define $\omega=\det_{*}(\Omega)$, and for each $U_{jk}\in\mathcal{N}(\mathcal{U})$, let $R_{jk}:U_{jk}\to SO(2)$ be defined by $R_{jk}(b)=\Omega_{jk}(b)r(\omega_{jk})$.
+Since $\mathcal{U}$ is a good open cover, each intersection $U_{jk}$ is contractible, so each $R_{jk}$ has a continuous lift $\Theta_{jk}:U_{jk}\to\mathbb{R}$ satisfying $\exp(\Theta_{jk}(b))=R_{jk}(b)$.
+Thus, $\Omega$ decomposes as $\Omega_{jk}(b)=\exp(\Theta_{jk}(b))r(\omega_{jk})$.
+For each $x\in X$ and $j\in J$ such that $\pi(x)\in U_{j}$, choose a branch of $\log$ whose domain contains the (finite) set $\{\tilde{\Omega}_{jk}(\pi(x))f_{k}(x)\in\mathbb{S}^{1}:\pi(x)\in U_{k}\}$.
+Note that whenever $\pi(x)=b\in\pi^{-1}(U_{j}\cap U_{k})$ for some $j,k\in J$, we have
+$\log\left(\widetilde{\Omega}_{jk}(b)f_{k}(x)\right)\ =\ \kappa_{jk}(b)+\Theta_{jk}(b)+\omega_{jk}\log\left(f_{k}(x)\right)$
+for some $\kappa_{jk}(\pi(x))\in\mathbb{Z}=\ker(\exp)$, so
 $\tilde{\Omega}_{jk}(b)\tilde{f}_{k}(x)\quad$ $=\quad\tilde{\Omega}_{jk}(b)\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{kl}f_{l}(x)\right)\right)$
@@
 $=\quad\exp\left(\sum_{l}\rho_{l}(b)\left(\Theta_{jk}(b)+\omega_{jk}\log\left(\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)\right)$
-$=\quad\exp\left(\sum_{l}\rho_{l}(\pi(x))\log\left(\tilde{\Omega}_{jk}(b)\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
+$=\quad\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{jk}(b)\tilde{\Omega}_{kl}(b)f_{l}(x)\right)\right)$
 $=\quad\exp\left(\sum_{l}\rho_{l}(b)\log\left(\tilde{\Omega}_{jl}(b)f_{l}(x)\right)\right)$
@@
 Suppose $\Omega$ witnesses that $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U})$ and $d_{Z}(\Omega,\widetilde{\Omega})<\delta$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
-If $\delta+\alpha\sqrt{2}<2$, then for any partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and $x\in U_{j}$ for some $j\in J$, $\widetilde{f}_{j}(x)$ as defined by (17) is equal to the unique weighted Karcher mean of $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ with weights $\{\rho_{k}(\pi(x))\}_{k\in K_{x}}$, where $K_{x}=\{k\in J:x\in U_{j}\cap U_{k}\}$.
-Furthermore $d_{T}(\{f_{j}\},\{\widetilde{f}_{j}\})<\alpha+\frac{\delta}{\sqrt{2}}$, and the maps $\{\widetilde{f}_{j}\}_{j\in J}$ obtained by different choices of partition of unity are homotopic through $T_{0}(\pi,\mathcal{U})$.
+If $\delta+\alpha<\sqrt{2}$, then for any partition of unity $\{\rho_{j}\}_{j\in J}$ subordinate to $\mathcal{U}$ and $x\in U_{j}$ for some $j\in J$, $\widetilde{f}_{j}(x)$ as defined by (23) is equal to the Frechet mean of $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ with weights $\{\rho_{k}(\pi(x))\}_{k\in K_{x}}$, where $K_{x}=\{k\in J:x\in U_{jk}\}$.
+Furthermore $d_{\Gamma}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})<\alpha+\delta$, and the functions $\{\widetilde{f}_{j}\}_{j\in J}$ obtained by different choices of partition of unity are homotopic through $\Gamma_{0}(\pi,\mathcal{U})$.
 Proof.
-For any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{j}\cap U_{k})$, we have
-$|f_{j}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$ $\leq|f_{j}(x)-\Omega_{jk}f_{k}(x)|+|\Omega_{jk}f_{k}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$
-$<\alpha+\frac{|\Omega_{jk}-\widetilde{\Omega}_{jk}(\pi(x))|_{F}}{\sqrt{2}}$
-$\leq\alpha+\frac{\delta}{\sqrt{2}}$
-Thus, by Proposition A.11, $\widetilde{f}_{j}(x)$ as defined in (17) is the unique weighted Karcher mean of the set $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ as long as $\alpha+\frac{\delta}{\sqrt{2}}<\sqrt{2}$.
-Clearly we have $d_{T}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})\leq\alpha+\frac{\delta}{\sqrt{2}}$ in this case.
+For any $(jk)\in\mathcal{N}(\mathcal{U})$ and $x\in\pi^{-1}(U_{jk})$, we have
+$|f_{j}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|\quad$ $\leq\quad|f_{j}(x)-\Omega_{jk}f_{k}(x)|+|\Omega_{jk}f_{k}(x)-\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)|$
+$<\quad\alpha+\|\Omega_{jk}-\widetilde{\Omega}_{jk}(\pi(x))\|_{F}$
+$\leq\quad\alpha+\delta$
+Thus, by Proposition A.13, $\widetilde{f}_{j}(x)$ as defined in (23) is the unique weighted Karcher mean of the set $\{\widetilde{\Omega}_{jk}(\pi(x))f_{k}(x)\}_{k\in K_{x}}$ as long as $\alpha+\delta<\sqrt{2}$.
+Clearly we have $d_{\Gamma}(\{f_{j}\}_{j\in J},\{\widetilde{f}_{j}\}_{j\in J})\leq\alpha+\delta$ in this case.
 Now, suppose $\{\rho^{\prime}_{j}\}_{j\in J}$ is another partition of unity subordinate to $\mathcal{U}$.
-Each partition of unity can be viewed as the component functions of a map $\rho,\rho^{\prime}:B\to\Delta^{n-1}$.
-Since $\Delta^{n-1}$ is convex, we can form the straight line homotopy $h_{t}:B\to\Delta^{n-1}$ from $\rho$ to $\rho^{\prime}$.
-Replacing the components of $\rho$ with those of $h_{t}$ in (17) for each $t\in[0,1]$ gives a continuously varying family in $T_{0}(\pi,\mathcal{U})$, so $\{f_{j}\}$ and $\{\widetilde{f}_{j}\}$ are homotopic through $T_{0}(\pi,\mathcal{U})$.
+Each partition of unity can be viewed as the component functions of maps $\rho,\rho^{\prime}:B\to\Delta^{J}$ to the full simplex on $J$.
+Since $\Delta^{J}$ is convex, we can form the straight line homotopy $h_{t}:B\to\Delta^{J}$ from $\rho$ to $\rho^{\prime}$.
+Replacing the components of $\rho$ with those of $h_{t}$ in (23) for each $t\in[0,1]$ gives a continuously varying family in $\Gamma_{0}(\pi,\mathcal{U})$, so $\{f_{j}\}_{j\in J}$ and $\{\widetilde{f}_{j}\}_{j\in J}$ are homotopic through $\Gamma_{0}(\pi,\mathcal{U})$.
 Corollary 3.39.
-For $\alpha<\frac{\sqrt{2}}{28}$, Construction 3.23 defines a map $\Pi_{T}:T_{\alpha}(\pi,\mathcal{U})\to T_{0}(\pi,\mathcal{U})$ such that for any $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$, the families $\{f_{j}\}$ and $\Pi_{T}(\{f_{j}\})$ determine the same true bundle.
+For $\alpha<\frac{1}{28}$, Construction 3.36 defines a function $\Pi_{\Gamma}:\Gamma_{\alpha}(\pi,\mathcal{U})\to\Gamma_{0}(\pi,\mathcal{U})$ such that for any $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$, the systems $\mathfrak{F}$ and $\Pi_{\Gamma}(\mathfrak{F})$ determine the same true bundle.
 Proof.
-Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{28}<\frac{\sqrt{2}}{20}$.
-By Proposition 3.4, any witness satisfies $\Omega\in\check{Z}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, and Proposition 3.18 implies $d_{Z}(\Omega,\widetilde{\Omega})<(27\sqrt{2})\alpha$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
-The result then follows from Proposition 3.25.
+Suppose $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{1}{28}<\frac{\sqrt{2}}{28}<\frac{\sqrt{2}}{20}$.
+By Proposition 3.15, any witness satisfies $\Omega\in\check{Z}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$, and Proposition 3.29 implies $\delta=d_{Z}(\Omega,\widetilde{\Omega})<(27\sqrt{2})\alpha$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$.
+Thus,
+$\alpha+\delta\ <\ (1+27\sqrt{2})\alpha\ <\ \frac{1+27\sqrt{2}}{28}\ <\ \sqrt{2}$
+so the result follows by Proposition 3.38.
+To describe conditions under which a discrete approximate circle bundle can be uniquely identified with a class of true bundles, we introduce the following notation:
+Definition 3.40.
+For any $r>0$, $\beta,\varepsilon\geq 0$ and collection of open sets $\mathcal{U}$, define
+$\tau(r,\beta,\varepsilon,\mathcal{U})=\frac{1}{r}\left(\sup_{j\in J}\ \text{diam}(U_{j})+10\varepsilon+9\beta\right),\hskip 28.45274pt\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})=2\sin^{-1}\left(\frac{\tau(r,\beta,\varepsilon,\mathcal{U})}{2}\right)$ (24)
+Observe that if $\pi:X\to B$ is a discrete $(r,\beta,\varepsilon)$-approximate circle bundle and $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$, then
+$\tau(\Phi)\ \leq\ \tau(r,\beta,\varepsilon,\mathcal{U}),\hskip 28.45274pt\widetilde{\tau}(\Phi)\ \leq\ \widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+The following result shows that witnesses to systems of discrete approximate local angle functions emerging from trivializations subordinate to the same cover are roughly related by a change of gauge:
+Lemma 3.41.
+Let $\pi:X\to B$ be a discrete $(r,\beta,\varepsilon)$-approximate circle bundle for $r>0$, $\beta,\varepsilon\geq 0$ with trivializing cover $\mathcal{U}=\{U_{j}\}_{j\in J}$, and let $\Omega,\Omega^{\prime}\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness that $\mathfrak{F}(\Phi),\mathfrak{F}(\Phi^{\prime})\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\leq\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}$.
+There exists some $\zeta\in\check{C}^{0}(\mathcal{U};\underline{O(2)})$ such that
+$\|\Omega_{jk}^{\prime}-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}\|_{F}\quad\leq\quad 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+for all $(jk)\in\mathcal{N}(\mathcal{U})$.
+Proof.
+Define
+$\tau(\Phi,\Phi^{\prime})=\sup_{j\in J}\ \tau(\varphi_{j},\varphi_{j}^{\prime}),\hskip 28.45274pt\tilde{\tau}(\Phi,\Phi^{\prime})=2\sin^{-1}\left(\frac{\tau(\Phi,\Phi^{\prime})}{2}\right)$
+and observe
+$\tau(\Phi,\Phi^{\prime})\ \leq\ \tau(r,\beta,\varepsilon,\mathcal{U})\hskip 28.45274pt\tilde{\tau}(\Phi,\Phi^{\prime})\ \leq\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+Lemma 3.10 implies that for each $j$, there exists some $\zeta_{j}\in O(2)$ such that $d_{\infty}(f_{j}^{\prime},\zeta_{j}f_{j})\leq\tau(\Phi,\Phi^{\prime})$. Thus, for any $x\in\pi^{-1}(U_{jk})$,
+$|f_{j}^{\prime}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|\quad$ $\leq\quad|f_{j}^{\prime}(x)-\zeta_{j}f_{j}(x)|\ +\ |\zeta_{j}f_{j}(x)-\zeta_{j}\Omega_{jk}f_{k}(x)|\ +\ |\zeta_{j}\Omega_{jk}f_{k}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|$
+$=\quad|f_{j}^{\prime}(x)-\zeta_{j}f_{j}(x)|\ +\ |f_{j}(x)-\Omega_{jk}f_{k}(x)|\ +\ |\zeta_{k}f_{k}(x)-f_{k}^{\prime}(x)|$
+$\leq\quad\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi)+\tilde{\tau}(\Phi^{\prime},\Phi)$
+$=\quad 2\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi)$
+which implies
+$|\Omega_{jk}^{\prime}f_{k}^{\prime}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|\quad$ $\leq\quad|\Omega_{jk}^{\prime}f_{k}^{\prime}(x)-f_{j}^{\prime}(x)|+|f_{j}^{\prime}(x),\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|$
+$\leq\quad\tilde{\tau}(\Phi^{\prime})+(2\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi))$
+$\leq\quad 4\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+for all $x\in\pi^{-1}(U_{jk})$. Using (17) and applying Proposition A.12 to the matrix $\Omega_{jk}^{\prime}-\xi_{j}\Omega_{jk}\xi_{k}^{\prime}$ , we obtain
+$\|\Omega_{jk}^{\prime}-\xi_{j}\Omega_{jk}\xi_{k}^{\prime}\|_{F}\quad\leq\quad\frac{4\sqrt{2}}{1-2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\quad\leq\quad 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+We now arrive at the main result of this section:
+Theorem 3.42.
+Let $\xi=\pi:X\to B$ be a discrete $(r,\beta,\varepsilon)$-approximate circle bundle for some $r,\beta,\varepsilon\geq 0$, and let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be a trivializing cover.
+If $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{\sqrt{2}}{36}$, then for any $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$, the cocycle $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ is well-defined, and its cohomology class does not depend on $\Phi$.
+Proof.
+By Proposition 3.11, we have $\mathfrak{F}(\Phi)\in\Gamma_{\alpha(\Phi)}(\pi,\mathcal{U})$, where $\alpha(\Phi)$ is defined by (16).
+Thus, Proposition 3.15 implies $\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}_{(3\sqrt{2})\alpha(\Phi)}(\mathcal{U};\mathcal{C}_{O(2)})$.
+By Proposition 3.29, $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ is well-defined as long as $(3\sqrt{2})\alpha(\Phi)\leq\frac{1}{2}$, which is satisfied as long as $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{\sqrt{2}}{15}$.
+Now, suppose we have some other $\Phi^{\prime}\in T_{r,\varepsilon,\beta}(\pi,\mathcal{U})$.
+Let $\Omega=\text{wit}(\mathfrak{F}(\Phi))$ and let $\Omega^{\prime}=\text{wit}(\mathfrak{F}(\Phi^{\prime}))$ Proposition 3.30 implies that if $3\sqrt{2}\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}+d_{H}([\Omega],[\Omega^{\prime}])\leq\frac{1}{2}$, then $\Pi_{Z}(\Omega^{\prime})]=[\Pi_{Z}(\Omega)]$.
+By Lemma 3.41, $d_{H}([\Omega],[\Omega^{\prime}])\leq 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U}),$ so
+$3\sqrt{2}\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}+d_{H}([\Omega],[\Omega^{\prime}])\quad\leq\quad\frac{5}{4}(3\sqrt{2})\widetilde{\tau}(\Phi)+(5\sqrt{2})\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\quad\leq\quad(9\sqrt{2})\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
+We conclude that $\Pi_{Z}(\Omega)]=[\Pi_{Z}(\Omega^{\prime})]$ as long as $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{1}{18\sqrt{2}}=\frac{\sqrt{2}}{36}$.
+Remark 3.43.
+The cohomology class assigned to $\xi$ in the proposition above is also independent of the trivializing cover $\mathcal{U}$ following sense: suppose $\Phi^{\prime}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{V})$ for some other trivializing cover $\mathcal{V}=\{V_{j^{\prime}}\}_{j^{\prime}\in J^{\prime}}$ satisfying $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{V})<\frac{\sqrt{2}}{36}$.
+If $\mathcal{W}$ is a common refinement of $\mathcal{U}$ and $\mathcal{V}$, then the images of $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))]$ and $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi^{\prime}))]$ with respect to the induced inclusions coincide in $\check{H}^{1}(\mathcal{W};\mathcal{C}_{O(2)})$.
 4 Effective Computation Of Characteristic Classes
 In this section, we give explicit algorithms for computing the characteristic classes of a circle bundle from a discrete approximate cocycle representative.
-One can then use Construction 3.5 to compute classes from discrete approximate trivializations.
+One can then use Construction 3.16 to compute classes from discrete approximate trivializations.
 We also introduce a notion of persistence for these classes.
@@
 Relative to the reduced structure group $O(2)$, we have $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$.
-Moreover, Proposition A.9 implies $\text{Adj}_{r(\omega)}\Lambda=\Lambda^{r(\omega)}$ for any $\Lambda\in O(2)$ and $\omega=\pm 1\in\mathbb{Z}_{2}$, so the group multiplication law in $\mathbb{S}^{1}\rtimes_{r}\mathbb{Z}_{2}$ simplifies to
+Moreover, Proposition A.11 implies $\text{Adj}_{r(\omega)}(\Lambda)=\Lambda^{r(\omega)}$ for any $\Lambda\in O(2)$ and $\omega=\pm 1\in\mathbb{Z}_{2}$, so the group multiplication law in $\mathbb{S}^{1}\rtimes_{r}\mathbb{Z}_{2}$ simplifies to
 $(\Lambda_{1},\omega_{1})\cdot(\Lambda_{2},\omega_{2})=(\Lambda_{1}\Lambda_{2}^{\omega_{1}},\omega_{1}\omega_{2})$ (25)
-Thus, the cocycle condition for $(\mathbb{S}^{1}\rtimes\mathbb{Z}_{2})$-valued cocycles is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
+The cocycle condition for $(\mathbb{S}^{1}\rtimes\mathbb{Z}_{2})$-valued cocycles is therefore given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
 $\Lambda_{jk}\Lambda_{kl}^{\omega_{jk}}=\Lambda_{jl}\quad\quad\omega_{jk}\omega_{kl}=\omega_{jl}\\$ (26)
-It follows that in this context, the connecting homomorphism $\Delta_{\omega}$ described in Proposition 2.10 reduces to
-$\Delta_{\omega}([\Omega])_{jkl}=\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}$ (20)
-where $\Omega=(\varphi^{-1})_{*}(\Lambda,\Theta)$.
-For later use, we define $\text{abs}:O(2)\to SO(2)$ by $\text{abs}(\Omega)=\Omega r(\det(\Omega))$ – note $\varphi^{-1}(\Omega)=(\text{abs}(\Omega),\det(\Omega))$.
-In fact, it is easy to see that if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\geq 0$ and $\omega=\det_{*}(\Omega)$, then $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$.
+Thus, for any $\Lambda]\in\check{H}^{1}(\mathcal{U};\mathcal{C}_{O(2),\omega})$, the connecting homomorphism $\Delta_{\omega}$ described in Proposition 2.11 returns a class $\Delta_{\omega}([\Lambda])\in\check{H}^{2}(\mathcal{U};\underline{\mathbb{Z}})$ represented by the cocycle whose $(jkl)$-component is given by
+$\omega_{jk}\Theta_{kl}(b)-\Theta_{jl}(b)+\Theta_{jk}(b)$ (27)
+for all $b\in U_{j}\cap U_{k}\cap U_{l}$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$ is a lift of $\Lambda$ via $\exp_{*}$.
+For later use, we define
+$\begin{array}[]{rccl}\text{abs}:&O(2)&\longrightarrow&SO(2)\\ &A&\mapsto&Ar(\det(A))\end{array}$ (28)
+and note that $\varphi^{-1}(A)=(\text{abs}(A),\det(A))$.
+In fact, it is easy to see that if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\geq 0$ and $\omega=\det_{*}(\Omega)$, then the cochain $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{SO(2)})$ given by $\Lambda_{jk}(b)=\text{abs}(\Omega_{jk}(b))$ for all $U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$ and $b\in U_{j}\cap U_{k}$ lies in $\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$.
+We will denote the cochain $\Lambda$ so defined by $\text{abs}(\Omega)$.
+Given a real number $x$, we use the notation $\langle x\rangle$ to denote the nearest integer to $x$ (if the integer is not unique, we take the larger one).
+Similarly, for any cochain $\lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$, we denote the nearest integer-valued cochain (with respect to $d_{Z}$) by $\langle\lambda\rangle$.
 Below we give an explicit algorithm for computing characteristic classes from a discrete approximate representative.
-To prove stability, we first prove a generalization of a technical result from [SP23].
-Let $1\to F\to G\xrightarrow{\varphi}H\to 1$ be a central extension of Lie groups with H compact and F discrete.
+To prove stability, we first prove a generalization of a technical result from [24] (Theorem 6.6).
+Let $1\to F\to G\xrightarrow{\psi}H\to 1$ be a central extension of Lie groups with H compact and F discrete.
 Fix a bi-invariant Riemannian metric on $H$ and use it to metrize $H$ and $G$ with the geodesic distance and $F$ with the distance inherited from $G$.
@@
 Proposition 4.1.
-For $\varepsilon\leq\frac{\text{sys}(H)}{8}$ and any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$, we have a well-defined map $\Delta_{\omega}:\check{Z}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ given by $\Delta_{\omega}(\Lambda)=[\langle\delta^{1}_{\omega}\Theta\rangle]$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ is any lift of $\Lambda$ respect to $\varphi_{*}$.
-Furthermore, $\Delta_{\omega}$ descends to a map $\Delta_{\omega}:\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ which is stable in the sense that if $\Lambda],[\Lambda^{\prime}]\in\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$ satisfy $d_{H}([\Lambda],[\Lambda^{\prime}])<\frac{\text{sys}(H)}{8}$, then $\Delta_{\omega}([\Lambda])=\Delta_{\omega}([\Lambda^{\prime}])$.
+For $\varepsilon\leq\frac{\text{sys}(H)}{8}$ and any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$, we have a well-defined map $\beta_{\omega}:\check{Z}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ given by $\Delta_{\omega}(\Lambda)=[\langle\beta^{1}_{\omega}\Theta\rangle]$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ is any lift of $\Lambda$ respect to $\psi_{*}$ and $\langle\cdot\rangle$ denotes the unique nearest integer Furthermore, $\Delta_{\omega}$ descends to a map $\Delta_{\omega}:\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ which is stable in the sense that if $\Lambda],[\Lambda^{\prime}]\in\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$ satisfy $d_{H}([\Lambda],[\Lambda^{\prime}])<\frac{\text{sys}(H)}{8}$, then $\Delta_{\omega}([\Lambda])=\Delta_{\omega}([\Lambda^{\prime}])$.
 Remark 4.2.
@@
 Proof.
-The proof is essentially the same as that of Theorem 6.6 in [SP23].
+The proof is essentially the same as that of Theorem 6.6 in [24].
 The fundamental difference is that we are doing cohomology with local coefficients, so the coboundary maps are slightly different.
 On the other hand, because $\Gamma$ acts by isometries, the same arguments apply.
-In particular, for any $\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$, one can use the same steps shown in [SP23] to verify the 2-cochain given by $\beta_{jkl}=\langle(\delta_{\omega}^{1}\Theta)_{jkl}\rangle$ is a genuine cocycle, and the stability arguments are identical.
+In particular, for any $\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$, one can use the same steps shown in [24] to verify the 2-cochain given by $\beta_{jkl}=\langle(\delta_{\omega}^{1}\Theta)_{jkl}\rangle$ is a genuine cocycle, and the stability arguments are identical.
 An immediate result of the proposition above is the following:
 Corollary 4.3.
-For $\varepsilon\leq 1$, we have a well-defined map $\text{Cl}:\check{H}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{O(2)})\to\coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{H}^{2}(B;\mathbb{Z}_{\omega})$ given by $\Omega\longmapsto([\det_{*}\Omega],[\Delta_{\omega}(\text{abs}_{*}\Omega)])$.
+For $\varepsilon\leq 1$, we have a well-defined function
+$\text{Cl}:\check{H}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{O(2)})\to\coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{H}^{2}(B;\mathbb{Z}_{\omega})$
+given by $\Omega\longmapsto([\det_{*}\Omega],\Delta_{\omega}(\text{abs}(\Omega)))$.
 This map is 1-stable in the sense that $\text{Cl}(\Omega)=\text{Cl}(\Omega^{\prime})$ whenever $d_{H}(\Omega,\Omega^{\prime})<1$.
@@
 Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\leq 1$.
-Lemma 1.3 of [SP23] implies $\omega=\det_{*}(\Omega)$ is a cocycle and that $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any other $\Omega^{\prime}\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Omega^{\prime})<2$.
+Lemma 1.3 of [24] implies $\omega=\det_{*}(\Omega)$ is a cocycle and that $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any other $\Omega^{\prime}\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Omega^{\prime})<2$.
 If $\varepsilon<1/9$, then by Proposition 3.29, there exists a true cocycle $\Omega^{\prime}$ such that $d_{Z}(\Omega,\Omega^{\prime})<1$, so $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any such $\Omega^{\prime}$.
@@
 Thus, Proposition 4.1 gives a well-defined map $\Delta_{\omega}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ for all $\varepsilon\leq\frac{sys(SO(2)_{g})}{8}$, where $\omega=\det_{*}(\Omega)$.
-By Lemma A.18 from [SP23], $\frac{sys(SO(2)_{g})}{8}=\frac{2\pi\sqrt{2}}{8}=\frac{\pi\sqrt{2}}{4}$ with respect to geodesic distance.
-To obtain a result in terms of the Frobenius norm, we note Corollary A.17 from [SP23] implies we have an inclusion $\check{Z}_{1}^{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})\subset\check{Z}^{1}_{\frac{\pi\sqrt{2}}{4}}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})$.
-Thus, $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$, so $\Delta_{\omega}(\text{abs}_{*}(\Omega))\in\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ (and therefore $\text{Cl}(\Omega)$) is well-defined.
+By Lemma A.18 from [24], $\frac{sys(SO(2)_{g})}{8}=\frac{2\pi\sqrt{2}}{8}=\frac{\pi\sqrt{2}}{4}$ with respect to geodesic distance.
+To obtain a result in terms of the Frobenius norm, we note Corollary A.17 from [24] implies we have an inclusion $\check{Z}_{1}^{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})\subset\check{Z}^{1}_{\frac{\pi\sqrt{2}}{4}}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})$.
+Thus, $\text{abs}(\Omega)\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$, so $\Delta_{\omega}(\text{abs}(\Omega))\in\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ (and therefore $\text{Cl}(\Omega)$) is well-defined.
 […]
@@
 Because the isomorphism class of a circle bundle is completely determined by its characteristic classes, Proposition 4.1 implies that if $\varepsilon<\frac{1}{9}$, then any true cocycle $\widetilde{\Omega}$ such that $d_{Z}(\Omega,\widetilde{\Omega})<1$ must satisfy $\text{Cl}(\widetilde{\Omega})=\text{Cl}(\Omega)$.
-Below is an explicit algorithm for computing characteristic class representatives.
-Input: A cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ subordinate to a finite open cover $\mathcal{U}$.
-Output: A pair of cochains $\omega\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}\in\check{C}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
-  1. 1.
-    Compute $\omega_{jk}=\det(\Omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
-  2. 2.
-    For each $(jk)\in\mathcal{N}(\mathcal{U})$, compute $\Lambda_{jk}=\Omega_{jk}r(\omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, where $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$.
-Then, choose a lift $\Theta_{jk}\in\mathbb{R}$ of each $\Lambda_{jk}$ via the $\exp$ map defined in (9).
-  3. 3.
-    Compute $\tilde{e}_{jkl}=\langle\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}\rangle$ for each $(jkl)\in\mathcal{N}(\mathcal{U})$.
+Below is an explicit algorithm for computing characteristic class representatives:
+Algorithm 1 (Computation Of Characteristic Classes).
+1:
+2:A cochain $\Omega\in\check{C}^{1}\!\left(\mathcal{U};\underline{O(2)}\right)$ subordinate to a good, finite open cover $\mathcal{U}$.
+3:A pair of cochains $\omega\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}\in\check{C}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
+4:Compute $\omega_{jk}=\det(\Omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
+5:For each $(jk)\in\mathcal{N}(\mathcal{U})$, set $\Lambda_{jk}=\Omega_{jk}\,r(\omega_{jk})$ with $r(\pm 1)=\begin{bmatrix}1&0\\[2.0pt] 0&\pm 1\end{bmatrix}$, then choose a lift $\Theta_{jk}\in\mathbb{R}$ via the $\exp$ map in (8).
+6:For each $(jkl)\in\mathcal{N}(\mathcal{U})$, compute $\tilde{e}_{jkl}=\big\langle\,\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}\,\big\rangle$.
 Corollary 4.5.
-If $\mathcal{U}$ is a finite good open cover and $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\underline{O(2)})$ for some $\varepsilon<\frac{1}{2}$, the outputs $\omega$ and $\tilde{e}$ of the algorithm above are cocycles.
-If $\varepsilon<\frac{1}{9}$, then $\omega$ and $\tilde{e}$ represent the characteristic classes of $\Pi_{Z}(\Omega)$.
+If $\mathcal{U}$ is a good, finite open cover and $\Omega\in\check{Z}^{1}_{\varepsilon}\left(\mathcal{U};\underline{O(2)}\right)$ for some $\varepsilon<\frac{1}{9}$, the outputs $\omega$ and $\tilde{e}$ of Algorithm 1 represent the characteristic classes associated with $\Pi_{Z}(\Omega)]\in\check{H}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$.
+Proof.
+By Corollary 4.3, the classes $\det_{*}(\Omega)]$ and $\Delta_{\omega}(\text{abs}(\Omega))$ are respectively the first Stiefel-Whitney class and twisted Euler class of the true circle bundle represented by $\Pi_{Z}(\Omega)$.
+Expression (27) shows that $\widetilde{e}$ (as defined in 1 above) represents $\Delta_{\omega}(\text{abs}(\Omega))$.
 In the case where $B$ is a closed connected 2-manifold, the following describes how to explicitly compute the twisted Euler number associated with a twisted Euler class:
-Input: A finite good open cover $\mathcal{U}$ of a closed connected surface $B$ and cocycles $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, $\tilde{e}\in\check{Z}^{1}(\mathcal{U};\mathbb{Z}_{\omega})$.
-Output: The twisted Euler number $\tilde{\eta}$ associated to $\tilde{e}$ (up to choice of sign).
-  1. 1.
-    For $p=2,3$, construct the matrix representation $D_{\omega,p}$ for the twisted boundary operator $\partial_{\omega,p}:\check{C}_{p+1}(\mathcal{U};\mathbb{Z}_{\omega})\to\check{C}_{p}(\mathcal{U};\mathbb{Z}_{\omega})$ with respect to some total orderings on the simplices in each dimension (e.g., lexicographical).
+Algorithm 2 (Computation Of Twisted Euler Number).
+1:
+2:A finite good open cover $\mathcal{U}$ of a closed connected surface $B$ and cocycles $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$, $\tilde{e}\in\check{Z}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
+3:The twisted Euler number $\tilde{\eta}$ associated to $\tilde{e}$ (up to choice of sign).
+4:For $p=2,3$, construct the matrix representation $D_{\omega,p}$ for the twisted boundary operator $\partial_{\omega,p}:\check{C}_{p+1}(\mathcal{U};\mathbb{Z}_{\omega})\to\check{C}_{p}(\mathcal{U};\mathbb{Z}_{\omega})$ with respect to some total orderings on the simplices in each dimension (e.g., lexicographical).
 Explicitly, the non-zero entries of $D_{\omega,p}$ are given by
-    $(D_{\omega,1})_{jkl}^{kl}=\omega_{jk},\quad(D_{\omega,1})_{jkl}^{jl}=-1,\quad(D_{\omega,1})_{jkl}^{jk}=1\\$ (21)
+$(D_{\omega,1})_{jkl}^{kl}=\omega_{jk},\quad(D_{\omega,1})_{jkl}^{jl}=-1,\quad(D_{\omega,1})_{jkl}^{jk}=1$ (29)
 for each $(jkl)\in\mathcal{N}(\mathcal{U})$ and
@@
 for each $(jklm)\in\mathcal{N}(\mathcal{U})$.
-  2. 2.
-    Compute the Smith normal form $LD_{\omega,2}R=S$ of $D_{\omega,2}$ over $\mathbb{Z}$ and let $k$ denote the number of zero columns in $S$).
+5:Compute the Smith normal form $LD_{\omega,2}R=S$ of $D_{\omega,2}$ over $\mathbb{Z}$ and let $k$ denote the number of zero columns in $S$).
 If $k=1$, set $\mu_{B,\omega}$ denote the last column of $R$ and proceed to Step 4.
 Otherwise, let $K$ be the submatrix of $R$ containing the last $k$ columns, and let $B$ equal the last $k$ rows of the product $R^{-1}D_{\omega,3}$.
-  3. 3.
-    Compute the Smith normal form $L^{\prime}BR^{\prime}=S^{\prime}$ of $B$. Let $v$ denote the last column of $R^{\prime}$, and set $\mu_{B,\omega}=Kv$.
-  4. 4.
-    Compute $\tilde{\eta}=\sum_{(jkl)\in\mathcal{N}(\mathcal{U})}\tilde{e}_{jkl}\cdot(\mu_{B,\omega})_{jkl}$.
+6:Compute the Smith normal form $L^{\prime}BR^{\prime}=S^{\prime}$ of $B$. Let $v$ denote the last column of $R^{\prime}$, and set $\mu_{B,\omega}=Kv$.
+7:Compute $\tilde{\eta}=\sum_{(jkl)\in\mathcal{N}(\mathcal{U})}\tilde{e}_{jkl}\cdot(\mu_{B,\omega})_{jkl}$.
 Remark 4.6.
-Although Čech theory is typically formulated in terms of cohomology, there is a natural dual notion of Čech homology with the property that $\check{H}_{*}(\mathcal{U};\mathcal{L})\cong H_{*}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of a space $B$ and any local system $\mathcal{L}$ of abelian groups (in particular, given a good open cover $\mathcal{U}$, we can define $\check{C}_{k}(\mathcal{U};\mathcal{L})=\text{Hom}_{\mathbb{Z}}(\check{C}^{k}(\mathcal{U};\mathcal{L}),\mathbb{Z})$ and consider the associated homology groups – these groups then coincide with the corresponding singular homology).
+Although Čech theory is typically formulated in terms of cohomology, there is a natural dual notion of Čech homology with the property that $\check{H}_{*}(\mathcal{U};\mathcal{L})\cong H_{*}(B;\mathcal{L})$ for any good open cover $\mathcal{U}$ of a space $B$ and any local system $\mathcal{L}$ of abelian groups (in particular, given a good open cover $\mathcal{U}$, we can define $\check{C}_{k}(\mathcal{U};\mathcal{L})=\text{Hom}_{\mathbb{Z}}(\check{C}^{k}(\mathcal{U};\mathcal{L}),\mathbb{Z})$ and consider the associated homology groups – these groups then coincide with the corresponding singular homology.
+See [7] for more details).
 We can therefore interpret $D_{\omega}$ in the construction above as the matrix representation of a twisted boundary operator whose kernel is generated by the associated twisted fundamental class.
 Proposition 4.7.
-The algorithm above computes the twisted Euler number associated to a twisted Euler class (up to choice of sign).
+If $B$ is a closed connected surface, the algorithm above computes the twisted Euler number associated to a twisted Euler class (up to choice of sign).
 Proof.
 Since $B$ is a compact connected surface, we have $\text{rank}(H_{2}(B;\mathbb{Z}_{\omega}))=1$, and any generator coincides with the fundamental class up to sign.
-If $\mathcal{U}$ has no quadruple intersections, then $H_{2}(B;\mathbb{Z}_{\omega})=Z_{2}(B;\mathbb{Z}_{\omega})=\text{ker}\partial_{\omega,2}$, so the last column $r$ of the matrix $R$ above represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$ (specifically, the vector $r$ corresponds to the simplicial cycle $\sum_{U_{jkl}\in\mathcal{N}(\mathcal{U})^{(2)}}r_{jkl}\cdot U_{jkl}$).
+If $\mathcal{U}$ has no quadruple intersections, then $H_{2}(B;\mathbb{Z}_{\omega})=Z_{2}(B;\mathbb{Z}_{\omega})=\text{ker}\partial_{\omega,2}$, so the last column $r$ of the matrix $R$ above represents a generator of $H_{2}(B;\mathbb{Z}_{\omega})$ (specifically, the vector $r$ corresponds to the simplicial cycle $\sum_{(jkl)\in\mathcal{N}(\mathcal{U})^{(2)}}r_{jkl}\cdot U_{jkl}$).
 If $k>1$, then the columns of $K$ form a basis for $\ker D_{\omega,2}$.
@@
 5 Persistence Of Characteristic Classes
-Discrete approximate trivializations constructed from a real-world dataset are unlikely to satisfy the fairly restrictive criteria required for our theoretical guarantees to hold.
-Due to practical considerations such as non-uniform sampling density or the presence of outliers for which the feature map is ill-defined, we may have reason to believe that some of the local measurements incorporated into our model are more reliable than others.
+Due to practical considerations such as non-uniform sampling density or the presence of outliers for which our feature map is ill-defined, we may have reason to believe that some of the local measurements incorporated into our model are more reliable than others.
 In other cases, we may want to make a deliberate choice to ignore or modify certain parts of the data so that the restricted discrete approximate cocycle representative becomes trivial.
-Given a family of discrete approximate local trivializations $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ and a witness $\Omega$, we therefore introduce a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ which measures the alignment quality of each pair $f_{j}|_{\pi^{-1}(U_{j}\cap U_{k})}$ and $f_{k}|_{\pi^{-1}(U_{j}\cap U_{k})}$ by the corresponding $\Omega_{jk}$.
-We then construct an induced simplex-wise filtration $\mathcal{W}$ on $\mathcal{N}(\mathcal{U})$ and compute the persistence of $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ with respect to $\mathcal{W}$.
-In Section 6, we show how to construct a coordinatization map which is in some sense compatible with the restriction of $\Omega$ to any subcomplex $W^{r}$ in $\mathcal{W}$.
+Given a family of discrete approximate local angle functions $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ and a witness $\Omega$, we therefore introduce a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ which measures the alignment quality of each pair $f_{j}|_{\pi^{-1}(U_{j}\cap U_{k})}$ and $f_{k}|_{\pi^{-1}(U_{j}\cap U_{k})}$ by the corresponding $\Omega_{jk}$.
+We then construct an induced a simplex-wise filtration $\mathcal{W}^{\Omega}$ on $\mathcal{N}(\mathcal{U})$ and compute the persistence of $w_{1}(\Omega)$ and $\tilde{e}(\Omega)$ with respect to $\mathcal{W}^{\Omega}$.
+In Section 6, we show how to construct a coordinatization map which is in some sense compatible with the restriction of $\Omega$ to any subcomplex $W^{r}$ in $\mathcal{W}^{\Omega}$.
 5.1 Weights Filtration
 Definition 5.1.
-Suppose $\pi:X\to B$ is a map between metric spaces, $\mathcal{U}$ is an open cover of $B$ and $\Omega\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$.
-For each $U_{jk}\in\mathcal{N}(\mathcal{U})^{(1)}$, define
-$w(U_{jk})=\frac{1}{|\pi^{-1}(U_{jk})|}\sum_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-\Omega_{jk}(\pi(x))f_{k}(x)|$ (22)
-Extend $w$ to a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ by setting $w(U_{j})=0$ for all $j$ and inductively defining $w$ for higher simplices by $w(U_{\sigma})=\max_{\sigma^{\prime}\subset\sigma}w(U_{\sigma^{\prime}})$.
+Suppose $\pi:X\to B$ is a function between metric spaces, $\mathcal{U}$ is a good open cover of $B$ and $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})\cong C^{1}(\mathcal{N}(\mathcal{U});O(2))$ witnesses that $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\geq 0$.
+For each $(jk)\in\mathcal{N}(\mathcal{U})^{(1)}$, define
+$w(jk)=\frac{1}{|\pi^{-1}(U_{jk})|}\sum_{x\in\pi^{-1}(U_{jk})}|f_{j}(x)-\Omega_{jk}f_{k}(x)|$ (30)
+Extend $w$ to a weight function $w:\mathcal{N}(\mathcal{U})\to\mathbb{R}_{\geq 0}$ by setting $w(j)=0$ for all $j$ and inductively defining $w$ for higher simplices by $w(\sigma)=\max\limits_{\sigma^{\prime}\subsetneq\sigma}w(\sigma^{\prime})$.
 Order the simplices of $\mathcal{N}(\mathcal{U})$ by (increasing) weight, then by (increasing) dimension, then lexicographically.
@@
 Definition 5.3.
-Suppose $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$.
+Suppose $\Omega\in\check{C}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witnesses that $\{f_{j}\}\in\Gamma_{\alpha}(\pi,\mathcal{U})$.
 Let $\mathcal{W}^{\Omega}=\{W_{r}\}_{r=1}^{|\mathcal{N}(\mathcal{U})|}$ denote the associated weights filtration.
-Given a cochain $\Lambda\in\check{C}^{p}(\mathcal{U};\mathcal{F})$ for some presheaf of groups $\mathcal{F}$ and $p\geq 1$, we define the cobirth of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$, where $r$ is the largest index such that $\Lambda|_{W_{r}}$ is a cocycle.
-Similarly, we define the codeath of $\Lambda$ to be the pair $(r,w(U_{\sigma_{r}}))$ for $r$ the largest index such that $\Lambda|_{W_{r}}$ is a coboundary.
+Given a cochain $\Lambda\in\check{C}^{p}(\mathcal{U};\mathcal{F})$ for a presheaf of groups $\mathcal{F}$ and $p\geq 1$, we define the cobirth of $\Lambda$ to be the pair $(r,w(\sigma_{r}))$, where $r$ is the largest index such that $\Lambda|_{W_{r}}$ is a cocycle (assume $p\leq 1$ if $\mathcal{F}$ is a sheaf of non-abelian groups).
+Similarly, we define the codeath of $\Lambda$ to be the pair $(r,w(\sigma_{r}))$ for $r$ the largest index such that $\Lambda|_{W_{r}}$ is a coboundary.
 (a) Full nerve
@@
 Construction 5.4.
-Suppose $G$ is a PID and $G_{\omega}$ is a local system with stalk $G$.
-If $\mathcal{U}$ is finite, one can efficiently compute the cobirth and codeath of $\Lambda$ as follows: let $D$ denote the matrix representation of the coboundary operator $\delta_{\omega}:\check{C}^{*}(\mathcal{U};G_{\omega})\to\check{C}^{*}(\mathcal{U};G_{\omega})$ when the simplices are arranged in increasing order according to the weights filtration.
-Define $\hat{\Lambda}\in\mathbb{R}^{|\mathcal{N}(\mathcal{U})|}$ such that $\hat{\Lambda}_{r}=\Lambda(U_{\sigma_{r}})$ whenever $\sigma_{r}$ is a $p$-simplex and $\hat{\Lambda}_{r}=0$ otherwise.
-It is easy to see that the cobirth of $\Lambda$ is 1 less than the smallest index $r_{b}$ such that $(D\hat{\Lambda})_{b_{p}}$ is nonzero (if none exists, the cobirth is $|\mathcal{N}(\mathcal{U})|$).
+Suppose $R$ is a PID and $R_{\omega}$ is a local system with stalk $R$.
+If $\mathcal{U}$ is finite, one can efficiently compute the cobirth and codeath of $\Lambda$ as follows: let $D$ denote the matrix representation of the coboundary operator $\delta_{\omega}:\check{C}^{*}(\mathcal{U};R_{\omega})\to\check{C}^{*}(\mathcal{U};R_{\omega})$ when the simplices are arranged in increasing order according to the weights filtration.
+Define the vector $\hat{\Lambda}\in\mathbb{R}^{|\mathcal{N}(\mathcal{U})|}$ such that $\hat{\Lambda}_{r}=\Lambda(U_{\sigma_{r}})$ whenever $\sigma_{r}$ is a $p$-simplex and $\hat{\Lambda}_{r}=0$ otherwise.
+It is easy to see that the cobirth of $\Lambda$ is 1 less than the smallest index $r_{b}$ such that $(D\hat{\Lambda})_{r_{b}}$ is nonzero (if none exists, the cobirth is $|\mathcal{N}(\mathcal{U})|$).
 To determine the codeath, write $S=LDR$, where $L$ and $R$ are invertible and $S$ is the Smith normal form of $D$.
 For $p>0$, let $r_{p}$ denote the largest index of a row of $S$ corresponding to a $p$-simplex and containing a non-zero entry.
-The codeath of $\Lambda$ is 1 less than the largest index $d_{b}>r_{p}$ such that $(L\hat{\Lambda})_{d_{b}}$ is nonzero (if none exists, then the codeath equals the cobirth).
+The codeath of $\Lambda$ is 1 less than the largest index $r_{d}>r_{p}$ such that $(L\hat{\Lambda})_{r_{d}}$ is nonzero (if none exists, then the codeath equals the cobirth).
 Given a cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$, we can use Construction 5.4 to compute the cobirth index $b_{sw}$ and codeath index $d_{sw}$ of $\omega=\det_{*}(\Omega)\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ (here the local system is trivial).
@@
 For instance, if the cobirth of a class representative is $\infty$ (or all the weights are fairly small and close together), one may still want to introduce cuts in the base space in order to artificially create a trivial bundle which admits a global coordinate system.
-With this in mind, we offer a recursive approach for constructing a family $\{B^{r}\}$ of spaces and corresponding open covers $\{\mathcal{U}^{r}\}$ such that $W^{r}=\mathcal{N}(\mathcal{U}^{r})$.
+With this in mind, we offer a recursive approach for constructing a family of spaces $\{B^{r}\}$ and corresponding open covers $\{\mathcal{U}^{r}\}$ such that $W^{r}=\mathcal{N}(\mathcal{U}^{r})$ for each $r$.
 Construction 5.5.
-Let $\mathcal{W}^{\Omega}$ denote the weights filtration associated with some $\Omega\in C^{k}(\mathcal{N}(\mathcal{U});\underline{O(2)})$.
-Suppose $W^{r}\in\mathcal{W}^{\Omega}$ for $r>1$, and we have a space $B^{r}$ and open cover $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j\in J}$ such that $B_{r}=\bigcup_{j\in J}U^{r}_{j}$ and $W^{r}=\mathcal{N}(\mathcal{U}^{r})$.
+Let $\mathcal{W}^{\Omega}$ denote the weights filtration associated with $\Omega\in\check{C}^{k}(\mathcal{U};\underline{O(2)})$.
+Suppose $W^{r}\in\mathcal{W}^{\Omega}$ for $r>1$, and we have a space $B^{r}$ and open cover $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j\in J}$ such that $B^{r}=\bigcup_{j\in J}U^{r}_{j}$ and $W^{r}=\mathcal{N}(\mathcal{U}^{r})$.
 Let $\sigma_{r}\in\mathcal{N}(\mathcal{U})$ denote the simplex which appears in $W^{r}$ but not $W^{r-1}$.
@@
 6 Coordinatization
-Suppose $\mathcal{U}=\{U_{j}\}_{j=1}^{n}$ is a finite open cover of a metric space $B$ and $\Omega\in\check{Z}_{\varepsilon}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{\varphi_{j}\}_{j=1}^{n}\in T_{\alpha}(\pi,\mathcal{U})$ for some map $\pi:X\to B$ ($X$ a finite metric space).
-Let $\mathcal{W}=\{W^{r}\}$ denote the associated weights filtration of $\mathcal{N}(\mathcal{U})$.
+Suppose $\mathcal{U}=\{U_{j}\}_{j=1}^{n}$ is a good finite open cover of a metric space $B$ and $\Omega\in\check{Z}_{\varepsilon}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witnesses that $\{f_{j}\}_{j=1}^{n}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some function $\pi:X\to B$ ($X$ a finite metric space).
+Let $\mathcal{W}^{\Omega}=\{W^{r}\}$ denote the associated weights filtration of $\mathcal{N}(\mathcal{U})$.
 For each $r$, we show how to construct a coordinatization map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ for $d\leq 2n$ which is in some sense compatible with $\Omega|_{W^{r}}$.
-This gives a menu of options for models for the dataset.
-In the special case when $\Omega|_{W^{r}}$ is an approximate coboundary, we construct a ’global trivialization’ map $F^{r}:X\to B\times\mathbb{S}^{1}$.
+This gives a menu of options of models for the dataset.
+In the special case when $\Omega|_{W^{r}}$ is an approximate coboundary, we construct a ’global trivialization’ map $F^{r}:X\to B^{r}\times\mathbb{S}^{1}$.
   1. 1.
-    Step 1: (Choose A Filtration Stage) Choose a filtration stage $W^{r}\in\mathcal{W}$ of $\mathcal{N}(\mathcal{U})$.
+    Step 1: (Choose A Filtration Stage) Choose a filtration stage $W^{r}\in\mathcal{W}^{\Omega}$ of $\mathcal{N}(\mathcal{U})$.
 Use Construction 5.5 to construct the sets $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j=1}^{n}$, then choose a partition of unity $\{\rho^{r}_{j}\}_{j=1}^{n}$ subordinate to $\mathcal{U}^{r}$.
   2. 2.
-    Step 2: (Dimensionality Reduction) For each $j$, define $\widehat{\Phi}^{r}_{j}:U^{r}_{j}\to V(2,2n)$ by
-    $\widehat{\Phi}^{r}_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho^{r}_{0}(b)}\Omega_{0j}\\ \vdots\\ \sqrt{\rho^{r}_{n}(b)}\Omega_{nj}\\ \end{array}\right)$ (23)
-    Choose a value $d\leq 2n$ and apply Principal Stiefel Coordinates to the point cloud $X_{V}=\bigcup_{j}\Phi_{j}^{r}(U_{j}^{r})\subset V(2,2n)$ to obtain a map $\text{psc}:X_{V}\to V(2,d)$.
+    Step 2: (Dimensionality Reduction) For each $j$, define $\widehat{\Psi}^{r}_{j}:U^{r}_{j}\to V(2,2n)$ by
+    $\widehat{\Psi}^{r}_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho^{r}_{0}(b)}\Omega_{0j}\\ \vdots\\ \sqrt{\rho^{r}_{n}(b)}\Omega_{nj}\\ \end{array}\right)$ (31)
+    Choose a value $d\leq 2n$ and apply Principal Stiefel Coordinates [14] to the point cloud $X_{V}=\bigcup_{j}\Phi_{j}^{r}(U_{j}^{r})\subset V(2,2n)$ to obtain a map $\text{psc}:X_{V}\to V(2,d)$.
   3. 3.
-    Step 3: (Classifying Map) For each $j$, let $\widetilde{\Phi}_{j}^{r}=\text{psc}\circ\widehat{\Phi}_{j}^{r}:U_{j}^{r}\to V(2,d)$. Define $\tilde{f}^{r}:B^{r}\to\mathbb{R}^{d\times d}$ by
-    $\widetilde{f}^{r}(b)=\sum_{j=1}^{n}\rho^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)^{T}$ (24)
-    and let $f^{r}=\Pi\circ\tilde{f}^{r}:B^{r}\to Gr(2,d)$.
+    Step 3: (Classifying Map) For each $j$, let $\widetilde{\Psi}_{j}^{r}=\text{psc}\circ\widehat{\Psi}_{j}^{r}:U_{j}^{r}\to V(2,d)$. Define $\widetilde{g}^{r}:B^{r}\to\mathbb{R}^{d\times d}$ by
+    $\widetilde{g}^{r}(b)=\sum_{j=1}^{n}\rho^{r}_{j}(b)\widetilde{\Psi}^{r}_{j}(b)\widetilde{\Psi}^{r}_{j}(b)^{T}$ (32)
+    and let $g^{r}=\Pi\circ\tilde{g}^{r}:B^{r}\to Gr(2,d)$.
   4. 4.
-    Step 4: (Cocycle Projection) For each $j$, define $\Phi^{r}_{j}:U^{r}_{j}\to V(2,d)$ by $\Phi^{r}_{j}(b)=\widehat{\Pi}(f^{r}(b),\widetilde{\Phi}^{r}_{j}(b))$, where $\widehat{\Pi}$ is the map from Construction 3.15.
-Then, define $\Omega^{r}\in\check{Z}^{1}(\mathcal{U}^{r};\mathcal{C}_{O(2)})$ by $\Omega^{r}_{jk}(b)=(\Phi^{r}_{j}(b))^{T}\Phi^{r}_{k}(b)$.
+    Step 4: (Cocycle Projection) For each $j$, define $\Psi^{r}_{j}:U^{r}_{j}\to V(2,d)$ by $\Psi^{r}_{j}(b)=\widehat{\Pi}(g^{r}(b),\widetilde{\Psi}^{r}_{j}(b))$, where $\widehat{\Pi}$ is the map from Construction 3.26.
+Then, define $\Omega^{r}\in\check{Z}^{1}(\mathcal{U}^{r};\mathcal{C}_{O(2)})$ by $\Omega^{r}_{jk}(b)=(\Psi^{r}_{j}(b))^{T}\Psi^{r}_{k}(b)$.
   5. 5.
     Step 5: (Bundle Map) Define $F^{r}_{j}:\pi^{-1}(U^{r}_{j})\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ by
-    $F_{j}^{r}(x)=\Phi^{r}_{j}(\pi(x))\exp\left(\sum_{k=0}^{n}\rho_{k}^{r}(\pi(x))\log\Omega_{jk}^{r}(\pi(x))\varphi_{k}(x)\right)$ (25)
+    $F_{j}^{r}(x)=\Psi^{r}_{j}(\pi(x))\exp\left(\sum_{k=0}^{n}\rho_{k}^{r}(\pi(x))\log\Omega_{jk}^{r}(\pi(x))f_{k}(x)\right)$ (33)
     The $F^{r}_{j}$’s agree on overlaps, yielding a globally defined map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$.
 6.1 Coordinatization Pipeline For Trivial Bundles
-In the special case when a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U}^{r})$ is witnessed by a discrete approximate coboundary $\Omega$, we construct a global trivialization map $\varphi:X\to B\times\mathbb{S}^{1}$.
+In the special case when a family $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U}^{r})$ is witnessed by a discrete approximate coboundary $\Omega$, we construct a global trivialization map $\varphi:X\to B\times\mathbb{S}^{1}$.
 In particular, suppose $\omega=\text{det}_{*}(\Omega)\in\check{Z}^{1}(\mathcal{U}^{r};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}=\Delta_{\omega}(\Omega)\in\check{Z}^{2}(\mathcal{U}^{r};\mathbb{Z}_{\omega})$ are coboundaries.
@@
     Construct a solution to the linear system $\delta^{0}\phi=\omega$ over $\mathbb{Z}_{2}$.
-Define $\hat{\Omega}=\phi\cdot\Omega\in\check{Z}^{1}(\mathcal{U}^{r};\underline{SO(2)})$, and let $\hat{f}_{j}=\phi_{j}\circ f_{j}$ for all $j$.
+Define $\hat{\Omega}=\phi\cdot\Omega\in\check{Z}^{1}\left(\mathcal{U}^{r};\underline{SO(2)}\right)$, and let $\hat{f}_{j}=\phi_{j}\circ f_{j}$ for all $j$.
   2. 2.
@@
 Remark 6.1.
-Since trivial bundles are flat, we can often use Singer’s method [WS13] in practice to obtain a locally-constant true coboundary $\widetilde{\Omega}$ which is close to an approximate one.
+Since trivial bundles are flat, we can often use Singer’s method [33] in practice to obtain a locally-constant true coboundary $\widetilde{\Omega}$ which is close to an approximate one.
 Though we have no theoretical guarantees on stability, the fact that the result is locally-constant guarantees that the resulting trivialization map does not have unnecessary variation.
-7 Implementations
+7 Experiments
 Here we demonstrate the practical viability and utility of the algorithms we have described in this paper through a series of implementations on real and synthetic datasets.
-In each case, we used the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations.
 A summary of the results of each experiment can be found in Table 2.
-Note that the witness quality $\alpha$ in each case was well outside the range required to meet our theoretical hypotheses, which suggests that our approach is stable for a larger class of objects than the current results apply to.
-Lastly, further details about our implementation on optical flow data can be found in an upcoming companion paper.
+Note that the roughness value $\alpha$ in each case was well outside the range required to meet our theoretical hypotheses, which suggests that our approach is stable for a larger class of objects than those covered by the present analysis.
+Further details about our implementation on optical flow data can be found in the companion paper [32].
+Additional examples, implementation details and documentation are provided in the accompanying open-source software package.
 Experiment Base Space Ambient Dim Sample Size SW1 $\widetilde{\eta}$
 Optical Flow Patches $\mathbb{RP}^{1}$ 18 25,000 0 0
-Prism Densities $\mathbb{RP}^{2}$ 15,408 25,000 $\neq 0$ $\pm 3$
-Star Densities $\mathbb{S}^{2}$ 15,408 25,000 0 $\pm 10$
-Table 1: A summary of the setups for the three experiments.
-Experiment $\varepsilon_{\text{trivs}}$ $\delta$ $\alpha$ $\varepsilon_{\text{cocycle}}$ $d_{Z}(\Omega,\Pi_{Z}(\Omega))$
-Optical Flow Patches 0.976 0 0.976 0 0
-Prism Densities 0.582 0.272 0.799 0.250 0.248
-Star Densities 1.120 0.620 2.949 0.501 0.471
+Folded Klein Bottle $\mathbb{S}^{1}$ 8 10,000 $\neq 0$ 0
+Prism Densities $\mathbb{RP}^{2}$ 15,408 10,000 $\neq 0$ $\pm 3$
+Table 1: A summary of the setups for the experiments.
+Each entry in the SW1 column indicates whether the Stiefel-Whitney class of the corresponding circle bundle model vanishes.
+Entries in the $\widetilde{\eta}$ column indicate the twisted Euler numbers of the models.
+Experiment $\varepsilon$ $\delta$ $\alpha$ $\varepsilon_{\text{cocycle}}$ $d_{Z}(\Omega,\Pi_{Z}(\Omega))$
+Optical Flow Patches 0.207 0 0.207 0 0
+Folded Klein Bottle 0.207 0 0.207 0 0
+Prism Densities 0.380 0.545 0.837 0.312 0.295
 Table 2: Summary of experimental results.
+The values of $\varepsilon$, $\delta$ and $\alpha$ for each system of discrete approximate angle functions are shown.
+For each experiment, the constructed witness $\Omega$ was found to lie in $\check{Z}^{1}_{\varepsilon_{\text{cocycle}}}(\mathcal{U};\underline{O(2)})$.
 Experiment $(\mathcal{U})$ $w_{\text{max}}$ $b_{\omega}$ $d_{\omega}$ $b_{\widetilde{e}}$ $d_{\widetilde{e}}$ $(W^{r})^{(1)}$ $(W^{r})^{(2)}$
-Optical Flow Patches 12 0.976 0.976 0.976 0.976 0.976 12/12 0/0
-Prism Densities 73 0.987 0.987 0.231 0.987 0.231 70/216 4/73
-Star Densities 146 1.939 1.939 1.939 1.939 1.912 431/432 286/288
+Optical Flow Patches 12 0.088 $\infty$ $\infty$ $\infty$ $\infty$ 12/12 0/0
+Folded Klein Bottle 12 0.088 $\infty$ $\infty$ $\infty$ $\infty$ 12/12 0/0
+Prism Densities 60 0.176 $\infty$ 0.176 $\infty$ 0.088 67/177 11/118
 Table 3: Summary of persistence information for the computed characteristic class representatives in each experiment.
+Here $w_{\text{max}}$ denotes the maximum weight among all simplices in $\mathcal{N}(\mathcal{U})$.
+The cobirth and codeath weights of the Stiefel-Whitney and twisted Euler classes are respectively denoted by $b_{\omega}$, $d_{\omega}$, $b_{\widetilde{e}}$ and $d_{\widetilde{e}}$.
+$W^{r}$ denotes the stage of the weights filtration corresponding to $\min\{d_{\omega},d_{\widetilde{e}}\}$; the number of 1 and 2-simplices in $W^{r}$ are denoted by $(W^{r})^{(1)}$ and $(W^{r})^{(2)}$, respectively.
 7.1 Optical Flow Patches
-Optical flow refers to the apparent motion of objects between consecutive frames of a video [HS81].
+Optical flow refers to the apparent motion of objects between consecutive frames of a video [11].
 This perceived motion is characterized by an assignment of a vector $v\in\mathbb{R}^{2}$ to each pixel in each frame.
-The vector at pixel $(x,y)$ in frame $t$ points to the location in frame $t+1$ where pixel $(x,y)$ appears to move.
-The estimation of optical flow from a video is central to many computer vision tasks, including motion tracking, object segmentation, and video compression [MG15], [Zha+21].
-The Sintel dataset [But+12], derived from the open-source animated short film “Sintel” by the Blender Foundation, was created as a high-quality benchmark for testing and training optical flow models.
+The vector at pixel $(x,y)$ in frame $t$, points to the location in frame $t+1$ where pixel $(x,y)$ appears to move.
+The estimation of optical flow from a video is central to many computer vision tasks, including motion tracking, object segmentation, and video compression [16], [34].
+The Sintel dataset [4], derived from the open-source animated short film “Sintel” by the Blender Foundation, was created as a high-quality benchmark for testing and training optical flow models.
 Figure 4: A frame from the Sintel video with sample (scaled) optical flow vectors shown
-In [Ada+20], Adams et al.
-studied the nonlinear statistics of 3 x 3 high-contrast optical flow patches from [But+12].
-They proposed that a dense core subset $X$ of (normalized) high-contrast patches is concentrated around a 2-manifold in $\mathbb{R}^{18}$ homeomorphic to $\mathbb{T}^{2}$.
+In [1], Adams et al.
+studied the nonlinear statistics of 3 x 3 high-contrast optical flow patches from [4].
+They proposed that a dense core subset $X$ of (normalized) high-contrast patches is concentrated around a 2-manifold in $\mathbb{R}^{18}$ homeomorphic to the 2-dimensional torus $\mathbb{T}^{2}$.
 To support their claim, they introduced a feature map $\pi_{\text{flow}}:X\to\mathbb{RP}^{1}$ which roughly measures a patch’s predominant axis of flow (see Figure 5).
-They then offered evidence that the data in $X$ over any small interval in $\mathbb{RP}^{1}$ has circular topology, clearly supporting the hypothesis that an underlying circle bundle structure is present.
+They then offered evidence that the data in $X$ over any small interval in $\mathbb{RP}^{1}$ has circular topology, supporting the hypothesis that an underlying circle bundle structure is present.
 Figure 5: A sample of mean-centered and normalized high-contrast optical flow patches labeled by predominant direction (represented as an angle in $0,\pi)$)
@@
 The total spaces of the two bundles are the torus and the Klein bottle.
-To confirm the torus model proposed by [Ada+20], we used the same pre-processing to obtain a sample of 50,000 normalized high-contrast optical flow patches from the Sintel dataset and refine to a sample $X_{\text{flow}}$ of 25,000 patches from the dense core subset $X(300,50)$.
+To confirm and provide coordinates for the torus model proposed by [1], we used the same pre-processing to obtain a sample of 50,000 normalized high-contrast optical flow patches from the Sintel dataset and refine to a sample $X_{\text{flow}}$ of 25,000 patches from the dense core subset $X(300,50)$.
 We then computed the predominant axis of flow $\pi_{\text{flow}}(x)\in\mathbb{RP}^{1}$ for each patch $x\in X_{\text{flow}}$.
-Figure 5 shows the persistence diagram for a random subsample of 1000 points from $X_{\text{flow}}$.
-For comparison, we also computed persistence diagrams for a random 1000-point sampling from $X(300,30)$ and a noisy synthetic sampling of the proposed torus model.
+Figure 6 shows the persistence diagram for a random subsample of 1,000 points from $X_{\text{flow}}$.
+For comparison, we also computed persistence diagrams for a random 1,000-point sampling from $X(300,30)$ and a noisy synthetic sampling of the theoretical proposed torus model.
 Observe that the synthetic dataset exhibits the Betti signature of $\mathbb{T}^{2}$ ($\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$) over a large range of scales, while the most persistent Betti signature for the samples of $X(300,50)$ and $X(300,30)$ is $\beta_{0}=1$, $\beta_{1}=1$, $\beta_{2}=0$.
-Figure 5:
-On the other hand, as observed in [Ada+20], the local picture of $X_{\text{flow}}$ with respect to $\pi_{\text{flow}}$ tells a different story.
+Figure 6: Persistence diagrams for the top 50% and 30% densest subsets of the original dataset (left, middle), and those of a synthetically generated dataset from the hypothesized torus model (right).
+On the other hand, as observed in [1], the local picture of $X_{\text{flow}}$ with respect to $\pi_{\text{flow}}$ tells a different story.
 We covered $\mathbb{RP}^{1}$ with 12 open intervals of the form $U_{j}=(\frac{j\pi}{12}-\frac{\pi}{16},\frac{j\pi}{12}+\frac{\pi}{16})$ (note that $\frac{\pi}{16}$ is the maximal radius for which the nerve of $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ has no triple intersections).
@@
 Figure 7: PCA and persistence diagrams for the sets $\{\pi^{-1}(U_{j})\}_{j=1}^{12}$
-We computed a family $\{f_{j}\}_{j=1}^{12}$ of discrete approximate local trivializations for $X_{\text{flow}}$ subordinate to $\mathcal{U}$ using the sparse circular coordinates algorithm, then obtained a witness $\Omega\in\check{C}^{1}(\mathcal{N}(\mathcal{U});O(2))$ using the procedure outlined in Construction 3.5.
-Figure 7 shows a visualization of $\mathcal{N}(\mathcal{U})$; the weight $w(\Omega_{jk})$ of each edge $(jk)$ is shown in black, and the values assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)$ are shown in red.
+We computed a family $\{f_{j}\}_{j=1}^{12}$ of discrete approximate local trivializations for $X_{\text{flow}}$ subordinate to $\mathcal{U}$ using PCA, then obtained a witness $\Omega\in\check{C}^{1}(\mathcal{N}(\mathcal{U});O(2))$ using the procedure outlined in Construction 3.16.
+Figure 8 shows a visualization of $\mathcal{N}(\mathcal{U})$; the weight $w(jk)$ of each edge $(jk)$ is shown in black, and the values assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)$ are shown in red.
 A 0-cochain $\phi\in C^{0}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ satisfying $\delta^{0}\phi=\omega$ is shown in blue, confirming a topological torus structure underlying the data.
@@
 Each edge is labeled by its associated weight.
-The components of the Stiefel-Whitney class representative $\omega=\text{det}_{*}(\Omega)$ are shown in blue, and a potential $\phi$ for $\omega$ is shown in red.
+The components of the Stiefel-Whitney class representative $\omega$ are shown in blue, and a potential $\phi$ for $\omega$ is shown in red.
 Once we verified the global triviality of the underlying bundle structure, we used the coordinatization pipeline described in Section 6 to obtain a global coordinate system for $X_{\text{flow}}$.
-To our surprise, we discovered that a significant portion of the patches had ambiguous predominant flow directions despite being high-contrast (see Figure 8).
-Such patches are not accounted for in the torus model proposed by [Ada+20].
-Upon further investigation, we found that the fibers of the predominant direction map $\pi$ are geometrically more like cylinders than circles, and that this extra degree of freedom locally parametrizes variations in directional strength among patches in $X_{\text{flow}}$.
-We present the details of our unexpected findings and a refined model for the dataset in an upcoming paper.
-The generalized model we describe also accounts for other families of high-contrast patches not featured in the original torus model which are particularly relevant for motion tracking analysis.
+It is important to note that a significant portion of the patches had ambiguous predominant flow directions despite being high-contrast (see Figure 9).
+Such patches are not accounted for in the torus model proposed by [1].
+Upon further investigation, we found that the fibers of the predominant direction map $\pi_{\text{flow}}$ are geometrically more like cylinders than circles, and that this extra degree of freedom locally parametrizes variations in directional strength among patches in $X_{\text{flow}}$.
+The generalized model we describe in our upcoming work also accounts for other families of high-contrast patches not featured in the original torus model which are particularly relevant for motion tracking analysis.
 Figure 9: A sample of high-contrast optical flow patches for which the predominant direction is essentially arbitrary
@@
 Patches near the left and right edges also roughly coincide, which reflects the global torus structure.
-Figure 9: (Left) The optical flow torus model proposed by [Ada+20].
-(Right) A sample of coordinatized patches from $X_{\text{flow}}$ with high directional strength, arranged by predominant flow direction (x) and assigned fiber coordinate (y).
+Figure 10: A sample of coordinatized patches from $X_{\text{flow}}$ with high directional strength, arranged by predominant flow direction (x) and assigned fiber coordinate (y).
+Compare with Figure 7 in [1]
 From a modeling perspective, the results of our analysis of $X_{\text{flow}}$ demonstrate how the interplay between local and global measurements depends crucially on the geometry of the chosen feature map.
 On the other hand, a careful fiberwise analysis helped us to both confirm an existing model and identify subtle, informative additional structure in the data which had previously been missed.
-See our upcoming paper for additional details.
+For additional details, see our companion paper [32].
+7.2 Folded Klein Bottle
+We demonstrate our pipeline on a noisy synthetic sampling of a 2-manifold $M\subset\mathbb{R}^{8}$ with the topology of a Klein bottle.
+Unlike our previous example, the dataset does not come equipped with any obvious choice of projection map, so we use the DREiMac library’s topologically-flavored circular coordinates algorithm [19], [20] to obtain an $\mathbb{S}^{1}$-valued base map.
+The data in the fibers of the resulting projection map is concentrated around topological (but not geometric) circles.
+We therefore also use the DREiMac algorithm to construct local circular coordinates.
+The model manifold $M$ is defined as follows: let $\gamma:\mathbb{R}\to\mathbb{R}^{4}$ be defined by $\gamma(t)=(\sin(t),\cos(t),\sin(2t),0)$ for all $t\in\mathbb{R}$, and let $F=\gamma(\mathbb{R})\subset\mathbb{R}^{4}$.
+Note that $F$ is topologically a circle, but geometrically it is folded.
+In particular, $F$ is invariant under a reflection in the $x$ and $y$ coordinates.
+Now, define $R:[0,2\pi]\to SO(4)$ by
+$R(t)=\left(\begin{array}[]{cccc}a&b&0&d\\ b&a&0&-d\\ 0&0&1&0\\ -d&d&0&c\\ \end{array}\right)$
+where
+$\tau=\frac{t}{2},\hskip 28.45274pta=\frac{1+\cos(\tau)}{2},\hskip 28.45274ptb=\frac{1-\cos(\tau)}{2}$
+$c=\cos(\tau),\hskip 28.45274ptd=\frac{\sqrt{2}}{2}\sin(\tau)$
+Observe in particular that $R(0)=I$, $R(2\pi)$ is the matrix which maps $(x,y,z,w)$ to (y,x,z,w)(y,x,z,-w) and $R(t)$ fixes the $z$-axis for all $t$. Finally, define $\widetilde{M}$ by
+$\widetilde{M}=\{(u,v)\in\mathbb{R}^{4}\times\mathbb{R}^{4}:u=\gamma(t),\ v\in R(t)F\}$
+and apply an $SO(8)$ rotation to obtain the manifold $M$ (the rotation is simply to intermix the coordinates).
+We generated a dataset $X_{\text{kb}}\subset\mathbb{R}^{8}$ by sampling $10,000$ points from $M$ with a Gaussian noise level of $\sigma=0.05$.
+Figure 11 shows persistence diagrams with $\mathbb{Z}_{2}$- and $\mathbb{Z}_{3}$-coefficients for a subsample of $X_{\text{kb}}$ (recall that the Klein bottle has Betti numbers $\beta_{0}=0$, $\beta_{1}=1$, $\beta_{2}=0$, with 2-torsion classes in homology dimensions 1 and 2):
+Figure 11: Persistence diagrams for a subsample of $X_{kb}$
+We used the $\mathbb{Z}_{3}$-persistent class as input for the DREiMac circular coordinates algorithm to produce a map $\pi_{\text{kb}}:X_{\text{kb}}\to\mathbb{S}^{1}$.
+We then constructed a cover of $\mathbb{S}^{1}$ by metric balls $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ around equally-spaced landmarks.
+Figure 12 shows PCA projections of the data in $\pi^{-1}_{\text{kb}}(U_{j})$ for several $j$, confirming that PCA could not be used to produce local circular coordinates.
+The points are colored according to angular coordinates computed using DREiMac:
+Figure 12: PCA projections of $\pi_{\text{kb}}^{-1}(U_{j})$ for several $j$.
+A characteristic class computation then confirmed the anticipated global Klein bottle structure; a visualization of the nerve decorated by the non-trivial Stiefel-Whitney class representative is shown in Figure 13.
+Figure 13: The nerve of the open cover $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ of $\mathbb{S}^{1}$.
+Each edge is labeled by its associated weight.
+The components of the Stiefel-Whitney class representative $\omega$ are shown in blue.
+Note the class is non-trivial on the full nerve, reflecting the underlying Klein bottle structure.
+After the heaviest edge is removed, the class becomes trivial; a potential for the restricted cocycle is shown in red.
 7.3 3D Densities
@@
 The ambient dimension for such data is inevitably very large, and distances computed using the standard Euclidean metric often fail to reflect one’s intuition about how densities should compare (for example, two densities which differ by a small translation may be very far apart with respect to Euclidean distance).
-One would ideally opt for an Earth-mover-type distance such as the Wasserstein metric [RTG00], but these metrics are prohibitively expensive to compute, so an effective local-to-global approach to topological modeling which circumvents the need to construct a full distance matrix is highly desirable.
+One would ideally opt for an Earth-mover-type distance such as the Wasserstein metric [22], but these metrics are prohibitively expensive to compute, so an effective local-to-global approach to topological modeling which circumvents the need to construct a full distance matrix is highly desirable.
+Furthermore, for our experiment below, we sample from a 3-manifold embedded in a space of 3D densities, and persistence diagrams for dimension 3 are computationally intractable using any metric; the only way to certify the global topology is with local computations.
 Consider the space $\mathcal{D}$ of real-valued functions on the unit ball $\mathbb{B}^{3}$.
 The space $\mathcal{D}$ has a natural right $SO(3)$-action defined by $(\rho\cdot A)(x)=\rho(A^{T}x)$ for all $\rho\in\mathcal{D}$, $A\in SO(3)$ and $x\in\mathbb{B}^{3}$.
-In particular, the orbit of any fixed $\rho_{0}\in\mathcal{D}$ is diffeomorphic to a quotient of $SO(3)$ by the corresponding stabilizer subgroup.
+In particular, the orbit of any $\rho_{0}\in\mathcal{D}$ is diffeomorphic to a quotient of $SO(3)$ by the corresponding stabilizer subgroup.
 For our demonstration, we sampled the $SO(3)$-orbit of a density $\rho_{\text{prism}}\in\mathcal{D}$ concentrated around the boundary of a triangular prism $P\subset\mathbb{B}^{3}$ (see Figure 14).
@@
 In Section A.2, we show $(\mathcal{O}_{\text{prism}},\pi,\mathbb{RP}^{2})$ is a non-orientable circle bundle with twisted Euler number $\pm 3$.
-We generated a synthetic dataset by applying 25,000 random rotations to $\rho_{\text{prism}}$ and sampling the resulting densities on a lattice $L$ (in particular, we restricted a 32 x 32 x 32 lattice on $1,1]^{3}$ to $\mathbb{B}^{3}$).
+We generated a synthetic dataset by applying 5,000 random rotations to $\rho_{\text{prism}}$ and sampling the resulting densities on a lattice $L$ (in particular, we restricted a 32 x 32 x 32 lattice on $1,1]^{3}$ to $\mathbb{B}^{3}$).
 The sampled density values were then flattened to vectors $X_{\text{prism}}\subset\mathbb{R}^{32^{3}}$.
@@
 Setting $\pi(v_{\rho})\in\mathbb{RP}^{2}$ equal to the axis associated with the minimum eigenvalue of $\widetilde{M}_{\rho}$ yields a well-defined map $\pi:X_{\text{prism}}\to\mathbb{RP}^{2}$.
-To classify and coordinatize the resulting discrete approximate circle bundle, we constructed a cover $\mathcal{U}=\{U_{j}\}_{j=1}^{73}$ of $\mathbb{RP}^{2}$ and used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations, then used Construction 3.5 to obtain a witness $\Omega\in C^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ and compute characteristic class representatives.
-As suggested by Figure 12, we found that the standard Euclidean metric was sufficient to model the small portion of $X_{\text{prism}}$ in each $\pi^{-1}(U_{j})$.
-Figure 13(a) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$.
-The node representing each $U_{j}\in\mathcal{U}$ is labeled by the rotated prism used to generate a sample $x_{j}\in X_{\text{prism}}$ whose base projection $\pi(x_{j})\in\mathbb{RP}^{2}$ was near the center of $U_{j}$.
+To classify the resulting discrete approximate circle bundle and infer the topology of the 3-manifold underlying $X_{\text{prism}}$, we constructed a cover $\mathcal{U}=\{U_{j}\}_{j=1}^{60}$ of $\mathbb{RP}^{2}$; as suggested by Figure 16, we found (to our surprise) that PCA was sufficient to compute local circular coordinates on the small portion of $X_{\text{prism}}$ in each $\pi^{-1}(U_{j})$.
+Figure 17(a)(a) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$.
+Each vertex $j\in\mathcal{N(U)}$ is labeled by the rotated prism used to generate a sample $x_{j}\in X_{\text{prism}}$ whose base projection $\pi(x_{j})\in\mathbb{RP}^{2}$ was near the center of $U_{j}$.
 Note in particular that the prisms at nodes which are connected by an edge have similar axes of symmetry.
-The edge representing each 1-simplex $U_{jk}\in\mathcal{N}(\mathcal{U})$ is colored according to the value assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$, red for 1-1).
+The edge representing each 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$ is colored according to the value assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$, red for 1-1).
 We verified that $\omega$ was not a coboundary, confirming that one cannot consistently choose an orientation for the data (since $H^{1}(\mathbb{RP}^{2};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$, verifying non-triviality is sufficient to uniquely determine the orientation class).
-With respect to $\omega$, the twisted Euler number of $\Omega$ was found to be 3-3, in agreement with our theoretical model.
+With respect to $\omega$, the twisted Euler number of $\Omega$ was found to be $\pm 3$, in agreement with our theoretical model.
+This confirmed the global topology of the underlying 3-manifold up to homeomorphism.
 Using the algorithm described in Section 5, we computed the cobirth and codeath of each characteristic class representative (see Table 3).
-Each representative was found to be a cocycle on the entire nerve of $\mathcal{U}$, but the two cocycles only became coboundaries after 146 of the 216 1-simplices in $\mathcal{N}(\mathcal{U})$ were removed.
+Note that each representative was found to be a cocycle on the entire nerve of $\mathcal{U}$.
+The twisted Euler class representative became a coboundary after a single edge was removed from $\mathcal{N}(\mathcal{U})$.
+This makes sense, because introducing a cut in $\mathbb{RP}^{2}$ yields a space which is homotopy equivalent to $\mathbb{S}^{1}$, hence has no homology in dimension $2$.
+On the other hand, the orientation class representative only restricted to a coboundary after 110 of the 177 1-simplices in $\mathcal{N}(\mathcal{U})$ were removed.
 One can interpret this result as a reflection of the compatibly non-orientable structures of the bundle and the base space.
@@
 The figure on the right shows the 1-skeleton of the maximal subcomplex of $\mathcal{N}(\mathcal{U})$ on which the characteristic class representatives restricted to coboundaries; the nodes which correspond to open sets containing the base projections of points in $\widetilde{X}_{\text{prism}}$ are highlighted.
-To illustrate, we used the coordinatization algorithm described in Section 6 to compute global coordinates for $X_{\text{prism}}$ compatible with the maximal stage $W\subset\mathcal{N}(\mathcal{U})$ of the weights filtration on which the characteristic class representatives became coboundaries.
-We isolated the set of densities $\widetilde{X}_{\text{prism}}\subset X_{\text{prism}}$ whose base point projections satisfied $|z|<0.1$, then projected these base points to the equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$.
-Figure 13(b) shows the 1-skeleton of $W$; the nodes which correspond to open sets containing the base projections of points in $\widetilde{X}_{\text{prism}}$ are highlighted.
+To illustrate further, we isolated the set of densities $\widetilde{X}_{\text{prism}}\subset X_{\text{prism}}$ whose base point projections satisfied $|z|<0.1$, then projected these base points to the equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$.
+Using the same cover of $\mathbb{RP}^{1}$ as we used to analyze the optical flow dataset, we recomputed local coordinates and approximate transition matrices and verified that the resulting discrete approximate cocycle produced a non-trivial orientation class representative.
+After dropping the heaviest edge from the nerve, we used the coordinatization algorithm described in Section 6 to compute global coordinates for $\widetilde{X}_{\text{prism}}$ compatible with the resulting subcomplex.
 Figure 18 shows a sample of $\widetilde{X}_{\text{prism}}$ arranged according to assigned base angles in $\mathbb{RP}^{1}$ and assigned fiber angles in $\mathbb{S}^{1}$.
@@
 Figure 18: A collection of coordinatized samples from $\widetilde{X}_{\text{prism}}$ arranged according to base projection angle (x) and assigned fiber coordinate (y).
-The dotted vertical line indicates an orientation flip in the coordinate system.
-Finally, we used the pipeline described in Section 6 to construct a point cloud embedding $F:X_{\text{prism}}\to V(2)$ which respects the global topology implied by the characteristic classes (recall these were well-defined on the entire nerve $\mathcal{N}(\mathcal{U})$).
-In this sense, our pipeline serves a similar function as Isomap, but with the preservation of topological (rather than geometric) features in mind.
+Finally, to demonstrate our bundle map algorithm, we used the procedure described in Section 6 to construct a point cloud embedding $F:X_{\text{prism}}\to V(2,60)\times_{O(2)}\mathbb{S}^{1}\subset\mathbb{R}^{120}$ for the full dataset which respects the global topology implied by the characteristic classes (recall these were well-defined on the entire nerve $\mathcal{N}(\mathcal{U})$).
+One could potentially reduce the ambient dimension further by applying Principal Stiefel Coordinates to the dataset of frames in $V(2,60)$ used to produce $F$.
 The mean projection errors for different projection dimensions are shown in Figure 19 .
 Figure 19:
-7.3 Geometric Signal Unwrapping
-Our final implementation provides an example of how the techniques and algorithms developed here can be integrated with existing topological modeling tools to form a larger local-to-global TDA framework.
-We consider a second dataset of discretized densities, sampled from the $SO(3)$-orbit of the density $\rho_{\text{star}}\in\mathcal{D}$ shown in Figure 16.
-Note that $\rho_{\text{star}}$ has a single 5-fold rotational symmetry in the $yz$-plane, so $\mathcal{O}_{\text{star}}$ is homeomorphic to the lens space $SO(3)/\mathbb{Z}_{5}\cong L(10,1)$.
-Applying the same projection map as we used for $\mathcal{O}_{\text{prism}}$ yields a bundle $(\mathcal{O}_{\text{star}},\pi,\mathbb{RP}^{2})$ with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$.
-Figure 16: A sample of 3D densities from $X_{\text{star}}$.
-Each density is represented by its 2D projection along the $x$-axis as well as the corresponding rotated triangle mesh used to generate the density.
-Figure 17: A PCA visualization of the data in a ’fat fiber’ of $X_{\text{star}}$ (right) over a small neighborhood in $\mathbb{RP}^{2}$ (left), portrayed here by a hemisphere with boundary identifications.
-[…]
-As in the previous experiment, we sampled 25,000 densities from $\mathcal{O}_{\text{star}}$ and formed a dataset $X_{\text{star}}\subset\mathbb{R}^{32^{3}}$ from the discretizations.
-We used a Mapper-like pipeline to cluster the data $X_{j}=X_{j}^{+}\coprod X_{j}^{-}$ over each open set $U_{j}$ in the same cover of $\mathbb{RP}^{2}$ (the sign labels for each pair of local clusters $X_{j}^{\pm}$ were assigned arbitrarily).
-Figure 18(a) shows a graph $G$ whose nodes are the local clusters $\{X_{j}^{\pm}\}_{j=1}^{73}$ and whose edges represent intersections between clusters.
-Edges between local clusters with the same sign label are shown in blue, and intersections between local clusters with opposite sign labels are shown in red.
-Figure 18(b) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$ with edges colored according to the values assigned by the induced connectivity cocycle $\nu\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$ and red for 1-1).
-Note that each edge in Figure 18(b) corresponds to a pair of edges in Figure 18(a) of the same color.
-Figure 19 gives a more detailed visualization of $G$; each local cluster $X_{j}^{\pm}$ is labeled by the rotated star pyramid used to generate a sample density it contains.
-For readability, edges connecting nodes with opposite sign labels are not shown explicitly, but their presence is indicated by light gray lines.
-Lastly, observe that $G$ can be canonically identified with the 1-skeleton of the nerve of a good open cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$.
-We verified that $\nu$ was a non-trivial cocycle, confirming that $X_{\text{star}}$ could not be globally separated into two connected components and that $\pi|_{X_{\text{star}}}$ could instead be lifted to a map $\tilde{\pi}:X_{\text{star}}\to\mathbb{S}^{2}$.
-To explicitly construct $\tilde{\pi}$, we selected a random $v_{\rho}\in X_{1}^{+}$ and chose an arbitrary lift $\tilde{\pi}(v_{\rho})\in\mathbb{S}^{2}$ of $\pi(v_{\rho})\in\mathbb{RP}^{2}$.
-For every other $v_{\rho^{\prime}}\in X_{1}^{+}$, we then defined $\tilde{\pi}(v_{\rho^{\prime}})$ to be the lift‘ of $\pi(v_{\rho^{\prime}})$ whose inner product with $\tilde{\pi}(v_{\rho})$ is positive.
-Similarly, for each $v_{\rho^{\prime}}\in X_{1}^{-}$, we defined $\tilde{\pi}(v_{\rho^{\prime}})$ so that $\tilde{\pi}(v_{\rho^{\prime}})\cdot\tilde{\pi}(v_{\rho})<0$.
-We then propagated this choice of lift across the entire dataset using a breadth-first search of $G$.
-Finally, we used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations for $X_{\text{star}}$ over $\mathbb{S}^{2}$, then used the algorithms described in Section 4 to obtain representatives of the orientation and twisted Euler classes.
-The results are summarized in Table 2.
-Note the Stiefel-Whitney class representative we obtained was trivial, and we obtained a twisted Euler number of $10$, as expected.
-(a) Local Cluster Graph $G$
-(b) Connectivity Cocycle $\nu$
-Figure 18: Visualizations of (a) the graph $G$ which represents the connectivity relations among local clusters in $X_{\text{star}}$.
-(b) The 1-skeleton of the nerve $\mathcal{N}(\mathcal{U})$ of the open cover of $\mathbb{RP}^{2}$ used to analyze $X_{\text{star}}$ (right).
-Each 1-simplex of $\mathcal{N}(\mathcal{U})$ is colored according to the assigned value of the connectivity class representative $\nu$ (blue for 1, red for -1), and the pair of corresponding edges in $G$ are assigned this same color.
-Figure 19: A detailed visualization of the local cluster graph $G$ which double covers the 1-skeleton of $\mathcal{N}(\mathcal{U})$.
-Nodes corresponding to local clusters labeled with a ++ are shown on the left, and the corresponding clusters labeled by a - are shown on the right.
-Edges which connect local clusters with the same sign label are shown in blue.
-For readability, edges which connect clusters with opposite sign labels are not shown explicitly, but are indicated with gray lines.
-Note the diagram above can equivalently be interpreted as representing the nerve of the cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$ which we ultimately used to model $X_{\text{star}}$ as a discrete approximate circle bundle.
-The process of lifting a bundle with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$ to a genuine circle bundle over a suitable covering space can be viewed as a form of geometric signal unwrapping.
-In other fields, classical signal unwrapping techniques (e.g., in interferometry, MRI, or phase retrieval) are used to resolve ambiguities in angular measurements (see, for example, [GP98]); our construction provides a topological analog in which the discrete fiber ambiguity is resolved via a covering, yielding a continuous, globally well-defined structure.
 8 Conclusion and Future Work
@@
 We also proposed a coordinatization scheme that mimics the construction of a bundle map to a universal bundle, which can be tuned to capture different global topologies using a persistence-based approach on characteristic classes.
-A natural direction for future research is to establish a recovery theorem providing conditions under which a finite metric space $X$ and a feature map $\pi:X\to B$ uniquely determine a true circle bundle, independent of the choice of discrete, approximate local trivializations.
 Beyond circle bundles, the theory and algorithms presented here could be extended in several ways.
@@
 Here we present some background definitions and results which are used throughout the paper.
-A.1 The Group $\text{Homeo}(\mathbb{S}^{1})$
-In this section, we prove some technical results about $\text{Homeo}(\mathbb{S}^{1})$.
-We denote the connected components of orientation-preserving and -reversing homeomorphisms by $\text{Homeo}_{+}(\mathbb{S}^{1})$ and $\text{Homeo}_{-}(\mathbb{S}^{1})$, respectively.
-A model for the universal cover $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$ of each component is given by
-$\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}=\{f\in\text{Homeo}(\mathbb{R}):f(x+1)=f(x)\pm 1\}$ (29)
-We then define $\widetilde{\text{Homeo}(\mathbb{S}^{1})}=\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}\coprod\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$.
-We denote the group of isometries of $\mathbb{S}^{1}$ by $\text{Iso}(\mathbb{S}^{1})<\text{Homeo}(\mathbb{S}^{1})$ and denote the connected components by $\text{Iso}_{\pm}(\mathbb{S}^{1})$.
-We have a natural identification of $\text{Iso}_{+}(\mathbb{S}^{1})$ with $\mathbb{S}^{1}$ itself via the map $R^{+}:\mathbb{S}^{1}\to\text{Iso}_{+}(\mathbb{S}^{1})$ which sends $e^{i\theta}$ the counter-clockwise rotation $R^{+}_{\theta}$ (equivalently, $R^{+}_{\theta}$ denotes left-translation in $\mathbb{S}^{1}$ by $e^{i\theta}$).
-Similarly, every orientation-reversing isometry is expressible as $R_{\theta}^{-}$ for some $\theta$, where $R_{\theta}^{-}(z)=e^{i\theta}\bar{z}$ for all $z\in\mathbb{S}^{1}$.
-We can model the universal covers $\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})$ as subsets of $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$:
-$\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})=\{T^{\pm}_{a}(x)=\pm x+a\ :\ a\in\mathbb{R}\}$ (30)
+A.1 Approximate Isometries Of The Circle
+The main result in this section is Lemma 3.10, which roughly states that discrete approximate local trivialization functions on a common domain are approximately related by an isometry of $\mathbb{S}^{1}$.
+We first review some definitions and introduce technical results which will be used in the proof.
+Definition A.1.
+Let $d_{\mathbb{S}^{1}}$ denote geodesic distance on $\mathbb{S}^{1}$, and let $d_{\infty}$ be the sup metric on the space $\text{Fun}(\mathbb{S}^{1},\mathbb{S}^{1})$.
+Note that for $t\in[0,2\pi]$ one has
+$d_{\mathbb{S}^{1}}(e^{is},e^{it})=\min\{|s-t|,\,2\pi-|s-t|\}\in[0,\pi]$
+The next three lemmas are used in the proof of Proposition A.5:
+Lemma A.2.
+Suppose $f,g\in\text{Homeo}(\mathbb{S}^{1})$ such that $\deg(f)\neq\deg(g)$. Then, there exists some $z\in\mathbb{S}^{1}$ such that f(z)=g(z)f(z)=-g(z).
+Proof.
+Let $h=(-g)^{-1}\circ f\in\text{Homeo}(\mathbb{S}^{1})$. Note that multiplication by 1-1 is a degree-1 homeomorphism, so the properties of degree imply
+$\deg(h)\quad=\quad\deg((-g)^{-1})\cdot\deg(f)\quad=\quad\deg(-g)\cdot\deg(f)\quad=\quad\deg(g)\cdot\deg(f)\quad=\quad-1$
+One can show that every orientation-reversing homeomorphism of $\mathbb{S}^{1}$ has a fixed point (this is an easy consequence of the Intermediate Value Theorem applied to a continuous lift), and if $z_{0}$ is a fixed point of $h$, then $(-g^{-1})\circ f(z_{0})=z_{0}$ (or equivalently, $f(z_{0})=-g(z_{0})$).
+Lemma A.3.
+Let $(X,d)$ be a metric space. For any fixed $x_{0}\in X$, the assignment $x\longmapsto d(x_{0},x)$ is $1$-Lipschitz:
+$|d(x_{0},x)-d(x_{0},x^{\prime})|\leq d(x,x^{\prime})\hskip 28.45274ptx,x^{\prime}\in X$
+Proof.
+Define $f(x)=d(x_{0},x)$ for all $x\in X$. For any $x,x^{\prime}\in X$, the triangle inequality implies
+$d(x,x^{\prime})\geq d(x_{0},x)-d(x_{0},x^{\prime})\hskip 28.45274ptd(x,x^{\prime})\geq d(x_{0},x^{\prime})-d(x_{0},x)$
+The result is immediate.
+Lemma A.4.
+Let $a,b\in\mathbb{S}^{1}$ such that $d_{\mathbb{S}^{1}}(a,b)=\frac{\pi}{2}$. Suppose $z,w\in\mathbb{S}^{1}$ such that
+$\max\{\left|d_{\mathbb{S}^{1}}(z,a)-d_{\mathbb{S}^{1}}(w,a)\right|,\quad\left|d_{\mathbb{S}^{1}}(z,b)-d_{\mathbb{S}^{1}}(w,b)\right|\}\quad\leq\quad\delta$
+for some $\delta<\frac{\pi}{2}$. Then $d_{\mathbb{S}^{1}}(z,w)\leq\delta$.
+Proof.
+Since $d_{\mathbb{S}^{1}}$ is $O(2)$-invariant, it suffices to prove the lemma for the case $a=1$ and $b=i$.
+Let $s,t\in[0,2\pi)$ and $z=e^{it},w=e^{is}\in\mathbb{S}^{1}$.
+We first show $d_{\mathbb{S}^{1}}(e^{it},e^{is})\leq\delta$ in the case when $t<\frac{\pi}{2}$ and $s<\pi$.
+We then argue that the hypothesis $\delta<\frac{\pi}{2}$ implies $e^{is}$ and $e^{it}$ cannot lie in opposite quadrants and use $O(2)$-invariance to conclude that the identity holds for all possible $s,t$.
+Suppose $t\in[0,\frac{\pi}{2})$. If $s\in[0,\pi)$, then
+$d_{\mathbb{S}^{1}}(e^{it},e^{is})\quad=\quad|t-s|\quad=\quad\left|d_{\mathbb{S}^{1}}(e^{it},1)-d_{\mathbb{S}^{1}}(e^{is},1)\right|\quad\leq\quad\delta$
+The case where $s\in[\frac{3\pi}{2},2\pi)$ reduces to the case above by applying a reflection, so $d_{\mathbb{S}^{1}}(e^{it},e^{is})\leq\delta$ in this case by $O(2)$-invariance of $d_{\mathbb{S}^{1}}$.
+The case where $s\in[\pi,\frac{3\pi}{2})$ is forbidden by the hypothesis $\delta<\frac{\pi}{2}$.
+To see this, suppose $s\in[\pi,\frac{3\pi}{2})$.
+Then,
+$\delta\quad\geq\quad\left|d_{\mathbb{S}^{1}}(e^{it},1)-d_{\mathbb{S}^{1}}(e^{is},1)\right|\quad=\quad|t-(2\pi-s)|\quad=\quad|2\pi-(s+t)|$ (37)
+$\delta\quad\geq\quad\left|d_{\mathbb{S}^{1}}(e^{it},i)-d_{\mathbb{S}^{1}}(e^{is},i)\right|\quad=\quad\left|\left(\frac{\pi}{2}-t\right)-\left(s-\frac{\pi}{2}\right)\right|\quad=\quad|\pi-(s+t)|$ (38)
+Since $\delta<\frac{\pi}{2}$, inequalities (37) and (38) are contradictory.
+We conclude $d_{\mathbb{S}^{1}}(e^{it},e^{is})$ for all feasible $s$ when $t\in[0,\frac{\pi}{2})$.
+Thus, by $O(2)$-invariance of $d_{\mathbb{S}^{1}}$, the result holds for all possible $s,t\in[0,2\pi)$.
 Proposition A.5.
-Suppose $\tilde{f}\in\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}$ is $L$-Lipschitz. Then, $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})\leq\frac{L-1}{L}$.
+Suppose $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is any function satisfying $\text{dist}(h)\leq\varepsilon$ for some $\varepsilon<\frac{\pi}{4}$. Then,
+  1. 1.
+    There exists some $\Omega\in O(2)$ such that $d_{\infty}(\Omega,h)\leq 2\varepsilon$,
+  2. 2.
+    The determinant of any such $\Omega$ is the same,
+  3. 3.
+    The map $d_{\infty}(h,\cdot):O(2)\to\mathbb{R}_{\geq 0}$ has a unique minimizer.
 Proof.
-Assume $\tilde{f}\in\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}$ (the proof for $\tilde{f}\in\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$ is similar).
-For any $x\in(0,1)$, we have $\tilde{f}(x)-\tilde{f}(0)\leq Lx$ and $\tilde{f}(1)-\tilde{f}(x)\leq L(1-x)$.
-On the other hand, we also have $\tilde{f}(0)<\tilde{f}(x)<\tilde{f}(1)$ for any $x\in(0,1)$, so the point $(x,\tilde{f}(x))$ is constrained to lie in a polygonal region.
-In particular, the quantity $|y-(x+\tilde{f}(0))|$ is maximized at the corner points $(\frac{1}{L},\frac{L-1}{L})$ and $(1-\frac{1}{L},\frac{L-1}{L})$.
-It follows that $d_{\infty}(\tilde{f},T^{+}_{\tilde{f}(0)})\leq\frac{L}{L-1}$.
-Proposition A.2.
-[…]
+First, observe that there is a unique $\Omega\in O(2)$ which satisfies $\Omega\circ h(1)=1$ and $d_{\mathbb{S}^{1}}(\Omega\circ h(i),i)\leq\varepsilon$.
+Let $\tilde{h}=\Omega\circ h$; to prove 1 above, it suffices to show $d_{\infty}(\tilde{h},\text{Id})\leq 2\varepsilon$.
+To see this, note that for any $z\in\mathbb{S}^{1}$, we have
+$\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),1)-d_{\mathbb{S}^{1}}(z,1)\right|\quad=\quad\left|d_{\mathbb{S}^{1}}(h(z),h(1))-d_{\mathbb{S}^{1}}(z,1)\right|\quad\leq\quad\varepsilon$ (39)
+and
+$\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),i)-d_{\mathbb{S}^{1}}(z,i)\right|\quad$ $\leq\quad\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),i)-d_{\mathbb{S}^{1}}(\tilde{h}(z),\tilde{h}(i))\right|\quad+\quad\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),\tilde{h}(i))-d_{\mathbb{S}^{1}}(z,i)\right|$
+$\leq\quad d_{\mathbb{S}^{1}}(i,\tilde{h}(i))\quad+\quad\left|d_{\mathbb{S}^{1}}(h(z),h(i))-d_{\mathbb{S}^{1}}(z,i)\right|$
+$\leq\quad 2\varepsilon$
+By Lemma A.4, $d_{\mathbb{S}^{1}}(\tilde{h}(z),z)\leq 2\varepsilon$ for all $z\in\mathbb{S}^{1}$, so $d_{\infty}(\tilde{h},\text{Id})\leq 2\varepsilon$.
+2. Suppose $\Omega^{\prime}\in O(2)$ satisfies $d_{\infty}(h,\Omega^{\prime})\leq 2\varepsilon$ and $\det(\Omega^{\prime})\neq\det(\Omega)$. By the triangle inequality,
+$d_{\infty}(\Omega^{\prime},\Omega)\quad\leq\quad d_{\infty}(\Omega^{\prime},h)\ +\ d_{\infty}(h,\Omega)\quad\leq\quad 4\varepsilon\quad<\quad\pi$
+On the other hand, Lemma A.2 implies $d_{\infty}(\Omega^{\prime},\Omega^{-1})=\pi$, a contradiction. We conclude $\det(\Omega^{\prime})=\det(\Omega^{-1})=\det(\Omega)$.
+3.
+Let $L_{h}:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L_{h}(\Omega)=d_{\infty}(h,\Omega)$.
+By compactness, at least one minimizer of $L_{h}$ exists.
+By 1, the set of minimizers is contained in $S_{\varepsilon}=L_{h}^{-1}([0,2\varepsilon])$, and 2 implies $S_{\varepsilon}$ is contained in a single connected component of $O(2)$.
+We assume minimizers of $L_{h}$ have positive determinant (if not, the same result is obtained by first pre-composing $h$ with any reflection in $O(2)$).
+For $t\in[0,2\pi)$, define $s(t)\in[0,2\pi)$ such that $h(e^{it})=e^{is(t)}$, and let g(t)=s(t)tg(t)=s(t)-t. If $R_{\theta}$ minimizes $L_{h}$ for some $\theta\in[0,2\pi)$, then
+$d_{\mathbb{S}^{1}}(e^{ig(t)},e^{i\theta})\quad=\quad d_{\mathbb{S}^{1}}(e^{i(s(t)-t)},e^{i\theta})\quad=\quad d_{\mathbb{S}^{1}}(e^{is(t)},e^{i(\theta+t)})\quad\leq\quad L_{h}(R_{\theta})\quad\leq\quad 4\varepsilon\quad<\quad\frac{\pi}{2}$
+for all $t\in[0,2\pi)$.
+It follows that the image $e^{ig([0,2\pi))}$ is contained in an arc of diameter at most $4\varepsilon<\pi$ containing $e^{i\theta}$.
+The unique minimizer of $L_{h}$ is then $R_{\theta^{*}}$, where $e^{i\theta^{*}}$ is the (unique) midpoint of the smallest arc containing $e^{ig([0,2\pi))}$.
+From uniqueness of the minimizer, we conclude $\theta^{*}=\theta\mod 2\pi$.
+We finally arrive at the main result of this section:
+Lemma A.6.
+Let $\pi:X\to B$ be a function between metric spaces.
+Suppose $\{\varphi_{j}:(\pi^{-1}(U_{j}),\pi,U_{j})\to(U_{j}\times\mathbb{S}^{1}_{r_{j}},p_{B},U_{j})\}_{j=1}^{2}$ are discrete $(r,\beta,\varepsilon)$-approximate local trivializations for some $r_{1},r_{2}\geq r$ and open sets $U_{1},U_{2}\subseteq B$ such that $\bar{B}_{\beta}(\pi(x))\subset U_{12}=U_{1}\cap U_{2}$ for some $x\in X$.
+Then,
+$|r_{1}-r_{2}|\leq\frac{2}{\pi}(\varepsilon+\beta)$
+and there exists some $\Omega_{12}\in O(2)$ such that
+$|f_{1}(x)-\Omega_{12}f_{2}(x)|\quad\leq\quad\tilde{\tau}(\varphi_{1},\varphi_{2})$
+for all $x\in\pi^{-1}(U_{12})$, where $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$ for $j=1,2$ and $\widetilde{\tau}(\varphi_{1},\varphi_{2})$ is defined by (14).
 Proof.
-Assume $f$ is orientation-preserving (the proof when $f$ reverses orientation is similar), and let $\tilde{f}$ be any lift of $f$.
-Let $w,z\in\mathbb{S}^{1}$, and let $\tilde{w},\tilde{z}\in\mathbb{R}$ be lifts of $w$ and $z$ such that $d_{\mathbb{R}}(\tilde{w},\tilde{z})\leq\frac{1}{2}$ (such a pair of lifts always exists).
-We have
-[…]
-[…]
-Proposition A.3.
-Suppose $\tilde{f},\tilde{g}\in\widetilde{\text{Homeo}(\mathbb{S}^{1})}$ are lifts of $f,g\in\text{Homeo}(\mathbb{S}^{1})$, respectively.
-If $d_{\infty}(\tilde{f},\tilde{g})<\varepsilon$, then $d_{\infty}(f,g)<2\pi\varepsilon$.
-Proof.
-Choose any $z\in\mathbb{S}^{1}$ and let $\tilde{z}\in\mathbb{R}$ be any lift. Then,
-$d_{\mathbb{S}^{1}}(f(z),g(z))\$$)\displaystyle=\ d_{\mathbb{S}^{1}}(\exp\circ\tilde{f}(\tilde{z}),\ \exp\circ\tilde{g}(\tilde{z}$))
-$)\displaystyle\leq\ 2\pi d_{\mathbb{R}}(\tilde{f}(\tilde{z}),\ \tilde{g}(\tilde{z}$))
-$<2\pi\varepsilon$
-Proposition A.4.
-If $f\in\text{Homeo}_{\pm}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,R_{\log f(1)}^{\pm})\leq 2\pi(\frac{L-1}{L})$.
-Proof.
-By Proposition A.2, any lift $\tilde{f}$ of $f$ is $L$-Lipschitz, so by Proposition A.1, any lift satisfies $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})<\frac{L-1}{L}$.
-Finally, using Proposition A.3, we obtain $d_{\infty}(f,R^{\pm}_{\log f(0)})=d_{\infty}(f,\exp\circ T_{\tilde{f}(0)})<2\pi(\frac{L-1}{L})$.
-A.2 Circle Bundles Over $\mathbb{S}^{2}$ and $\mathbb{RP}^{2}$
-In this section, we provide the details of the constructions and classifications of the two non-trivial circle bundle models described in the Implementations section.
-Since both models are associated with quotients of $SO(3)$, we begin by describing principal $\mathbb{S}^{1}$-bundle structures for $SO(3)$ and its universal cover $\mathbb{S}^{3}$.
+For $j=1,2$, let $\psi_{j}:B\times\mathbb{S}^{1}\to X$ be an almost-inverse for $\varphi_{j}$.
+Let $\pi_{j}=p_{B}\circ\varphi_{j}$ and $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$.
+Choose some $x_{12}\in\pi^{-1}(U_{1}\cap U_{2})$ such that $\bar{B}_{\beta}(\pi(x_{12}))\subset U_{1}\cap U_{2}$, let $b_{12}=\pi(x_{12})$ and let $b_{j}=p_{B}\circ\varphi_{j}(x_{12})$ for $j=1,2$.
+For any $z\in\mathbb{S}^{1}$, we have $d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{12})\leq\frac{\beta}{2}$, so $\psi_{2}(b_{2},z)\in\bar{B}_{\frac{\beta}{2}}(b_{12})\subseteq\bar{B}_{\beta}(b_{12})\subseteq U_{jk}$.
+We therefore have a well-defined function $\Lambda_{12}:\mathbb{S}^{1}_{r_{2}}\to\mathbb{S}^{1}_{r_{1}}$ given by $\Lambda_{12}(z)=f_{1}\circ\psi_{2}(b_{2},z)$ for all $z\in\mathbb{S}^{1}$.
+We now estimate $\text{dist}(\Lambda_{12})$ in order to invoke Proposition A.5.
+Let $T_{12}:\mathbb{S}^{1}\to U_{1}\times\mathbb{S}^{1}$ be defined by $T_{12}(z)=\varphi_{1}\circ\psi_{2}(b_{2},z)$ for all $z\in\mathbb{S}^{1}$.
+Note that for any $z\in\mathbb{S}^{1}$,
+$d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{12})\quad\leq\quad d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{2})+d_{B}(b_{2},b_{12})\quad\leq\quad\beta$
+so $\psi_{2}(b_{2},z)\in\bar{B}_{2\beta}(b_{12})\subset\pi^{-1}(U_{1})$ and $T_{12}(z)$ is well-defined. Moreover, for any $z,z^{\prime}\in\mathbb{S}^{1}$,
+$d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime}))\$$=\ d_{B}(\pi_{1}\circ\psi_{2}(b_{2},z),\ \pi_{1}\circ\psi_{2}(b_{2},z^{\prime}))$
+$\leq d_{B}(\pi_{1}\circ\psi_{2}(b_{2},z),\ \pi\circ\psi_{2}(b_{2},z))$
+$+d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{2})$
+$+d_{B}(b_{2},\pi\circ\psi_{2}(b_{2},z^{\prime}))$
+$+d_{B}(\pi\circ\psi_{2}(b_{2},z^{\prime}),\pi_{1}\circ\psi_{2}(b_{2},z^{\prime}))$
+$\leq 2\beta$
+Since $\text{dist}(\varphi_{1}),\text{dist}(\psi_{2})\leq\varepsilon$, we have
+$|d_{\infty}^{r_{1}}(T_{12}(z),T_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})|\quad\leq\quad 2\varepsilon$
+which in particular implies
+$d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))\quad\leq\quad d^{r_{1}}_{\infty}(T_{12}(z),T_{12}(z^{\prime}))\quad\leq\quad d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})+2\varepsilon$ (40)
+for any $z,z^{\prime}\in\mathbb{S}^{1}$. On the other hand, since
+$d^{r_{1}}_{\infty}(T_{12}(z),T_{12}(z^{\prime}))\quad=\quad\max\{d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime})),\hskip 14.22636ptd^{r_{1}}_{\mathbb{S}^{1}}(\Lambda_{12}(z),\ \Lambda_{12}(z^{\prime}))\}$
+we can apply the algebraic identity $v\geq\max\{u,v\}-u$ (for any $u,v\geq 0$) to obtain
+$d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\ \Lambda_{12}(z^{\prime}))\$$\geq\ d_{\infty}^{r_{1}}(T_{12}(z),T_{12}(z^{\prime}))-d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime}))$
+$\geq(d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})-2\varepsilon)-2\beta$
+$=d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})-2(\varepsilon+\beta)$
+Combining with (40), we conclude $\text{dist}(\Lambda_{12})\leq 2(\varepsilon+\beta)$ as a function from $\mathbb{S}^{1}_{r_{2}}$ to $\mathbb{S}^{1}_{r_{1}}$.
+As an immediate consequence, we find that for any $z,z^{\prime}\in\mathbb{S}^{1}$,
+$\pi r_{2}-2(\varepsilon+\beta)\quad\leq\quad d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))\quad\leq\quad\pi r_{1}$
+which implies $\pi r_{2}-2(\varepsilon+\beta)\leq\pi r_{1}$, or $\pi(r_{2}-r_{1})\leq 2(\varepsilon+\beta)$.
+Swapping the roles of $r_{1}$ and $r_{2}$ yields $\pi(r_{1}-r_{2})\leq 2(\varepsilon+\beta)$, so
+$|r_{2}-r_{1}|\quad\leq\quad\frac{2}{\pi}(\varepsilon+\beta)$
+Now, observe
+$2(\varepsilon+\beta)\quad\geq\quad\left|d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})\right|\quad=\quad\left|r_{1}d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-r_{2}d_{\mathbb{S}^{1}}(z,z^{\prime})\right|$
+which implies
+$\left|d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}(z,z^{\prime})\right|\quad$ $\leq\quad\left|d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-\frac{r_{2}}{r_{1}}\ d_{\mathbb{S}^{1}}(z,z^{\prime})\right|+\left|\frac{r_{2}}{r_{1}}d_{\mathbb{S}^{1}}(z,z^{\prime})-d_{\mathbb{S}^{1}}(z,z^{\prime})\right|$
+$\leq\quad\frac{2}{r_{1}}(\varepsilon+\beta)+\frac{d_{\mathbb{S}^{1}}(z,z^{\prime})}{r_{1}}|r_{2}-r_{1}|$
+$\leq\quad\frac{2}{r_{1}}(\varepsilon+\beta)+\frac{\pi}{r_{1}}|r_{2}-r_{1}|$
+$\leq\quad\frac{4}{r_{1}}(\varepsilon+\beta)$
+Since $r_{1}>\frac{16}{\pi}(\varepsilon+\beta)$ by definition of a discrete $(r,\beta,\varepsilon)$-approximate local trivialization, the inequality above implies that, as a function from $\mathbb{S}^{1}$ to $\mathbb{S}^{1}$ (with the usual unscaled geodesic distance), $\text{dist}(\Lambda_{12})<\frac{\pi}{4}$.
+Proposition A.5 then implies that the loss function $\Omega\longmapsto d_{\infty}(\Lambda_{12},\Omega)$ has a unique minimizer $\Omega_{12}$ in $O(2)$, and that $d_{\infty}(\Lambda_{12},\Omega_{12})\leq\frac{8}{r_{1}}(\varepsilon+\beta)$.
+Now, let $\tilde{x}\in\pi^{-1}(U_{12})$ be arbitrary.
+Let $\varphi_{2}(\tilde{x})=(\tilde{b},\tilde{z})\in U_{2}\times\mathbb{S}^{1}_{r_{2}}$ and let $\tilde{y}=\psi_{2}(b_{2},\tilde{z}_{2})\in\pi^{-1}(U_{12})$.
+Observe,
+$d_{B}(\tilde{b}_{2},b_{2})\quad$ $\leq\quad d_{B}(\tilde{b}_{2},\pi(\tilde{x}))+d_{B}(\pi(\tilde{x}),\pi(x_{12}))+d_{B}(\pi(x_{12}),b_{2})$
+$\leq\quad\frac{\beta}{2}+\text{diam}(U_{12})+\frac{\beta}{2}$
+$=\quad\text{diam}(U_{12})+\beta$
+By the codistortion property of $\varphi_{2}$ and $\psi_{2}$,
+$d_{X}(\tilde{x},\tilde{y})\quad\leq\quad d_{\infty}^{r_{2}}(\varphi_{2}(\tilde{x}),(b_{2},\tilde{z}_{2}))+\varepsilon\quad\leq\quad d_{B}(\tilde{b}_{2},b_{2})+\varepsilon\quad\leq\quad\text{diam}(U_{12})+\varepsilon+\beta$
+and since $\text{dist}(\varphi_{1})\leq\varepsilon$,
+$d_{\mathbb{S}^{1}}^{r_{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))\quad\leq\quad d_{\infty}^{r_{1}}(\varphi_{1}(\tilde{x}),\varphi_{1}(\tilde{y}))\quad\leq\quad d_{X}(\tilde{x},\tilde{y})+\varepsilon\quad\leq\quad\text{diam}(U_{12})+2\varepsilon+\beta$
+or, equivalently,
+$d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))\quad\leq\quad\frac{1}{r_{1}}\left(\text{diam}(U_{12})+2\varepsilon+\beta\right)$
+Finally, noting that $)f_{1}(\tilde{y})=\Lambda_{12}(f_{2}(\tilde{x}$)), we obtain
+$)\displaystyle d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),\Omega_{12}f_{2}(\tilde{x}$$)\quad$ $)\displaystyle\leq\quad d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))+d_{\mathbb{S}^{1}}(f_{1}(\tilde{y}),\Omega_{12}f_{2}(\tilde{x}$))
+$)\displaystyle\leq\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+2\varepsilon+\beta)+d_{\mathbb{S}^{1}}(\Lambda_{12}(f_{2}(\tilde{x})),\Omega_{12}f_{2}(\tilde{x}$))
+$\leq\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+2\varepsilon+\beta+8(\varepsilon+\beta))$
+$=\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+10\varepsilon+9\beta)$
+$\leq\quad\tau(\varphi_{1},\varphi_{2})$
+A.2 Circle Bundles Over $\mathbb{RP}^{2}$
+In this section, we provide some details of the construction and classification of the circle bundle underlying the dataset used in Section 7.3.
+The bundle is obtained as a quotient of the classical principal $\mathbb{S}^{1}$-bundle over $\mathbb{S}^{2}$ with total space $SO(3)$, briefly described below:
 For any $v\in\mathbb{S}^{2}$ and $\theta\in\mathbb{R}$ let $R_{v,\theta}\in SO(3)$ denote the counterclockwise rotation by $\theta$ about $v$.
@@
 Indeed, Adj is not a homeomorphism, nor is it $\mathbb{S}^{1}$-equivariant (in fact, one has $\text{Adj}_{q\cdot e^{i\theta}}=\text{Adj}_{q}\cdot e^{i(2\theta)}$).
-The Hopf bundle famously has Euler number -1 with respect to the standard orientation on $\mathbb{S}^{2}$ (for a proof, see [Hus94]).
-The following proposition immediately implies that the Euler number of the principal bundle $(SO(3),\pi,\mathbb{S}^{2})$ is 2-2, and will also be useful for classifying the circle bundles from our models in Section 7:
+The Hopf bundle famously has Euler number -1 with respect to the standard orientation on $\mathbb{S}^{2}$ (for a proof, see [12]).
+The following proposition immediately implies that the Euler number of the principal bundle $(SO(3),\pi,\mathbb{S}^{2})$ is 2-2, and will also be useful for classifying the circle bundle model underlying the density dataset in Section 7:
 Proposition A.7.
@@
 Proof.
-Let $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{S}^{1}})$ be a cocycle representative of $\xi$ subordinate to some open cover $\mathcal{U}$ of $B$.
-By the results in Section 2.1, $\varphi_{p}^{*}\xi$ is a principal $\mathbb{S}^{1}$-bundle which is represented by $\varphi_{*}\Omega$.
-If $\theta\in\check{C}^{1}(\mathcal{U};\mathbb{R})$ is any lift of $\widetilde{\Omega}$ (i.e., $\Omega_{jk}(b)=e^{i\theta_{jk}(b)}$ for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$), then
-$(\varphi_{*}\Omega)_{jk}(b)=\varphi\circ\Omega_{jk}(b)=\varphi(e^{i\theta_{jk}(b)})=e^{i(p\theta_{jk}(b))}$
-for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $p\theta$ is a lift of $\varphi_{*}\Omega$. The components of the Euler class $e(\varphi_{*}\xi)$ are then given by
-$e(\varphi_{*}\xi)_{jkl}$ $=p\theta_{jk}(b)+p\theta_{kl}(b)-p\theta_{jl}(b)$
-$=p(\theta_{jk}(b)+\theta_{kl}(b)-\theta_{jl}(b))$
-$=pe(\xi)_{jkl}$
-We conclude $e(\varphi_{*}\xi)=pe(\xi)$.
+See, for example, Chapter 4 of  [12].
 Corollary A.8.
@@
 Proposition A.9.
-Suppose $\xi=(E,\pi,B)$ is a principal $G$-bundle, and $E$ and $B$ have free, proper right actions by a group $H$ such that:
+Suppose $\xi=(E,\pi,B)$ is a principal $G$-bundle, and $E$ and $B$ have free, proper right actions by a topological group $H$ such that:
   1. 1.
@@
 We conclude $(\mathcal{O}_{x},\widetilde{\pi},\mathbb{S}^{2})$ is a principal $\mathbb{S}^{1}$-bundle with Euler number 2p-2p.
-Lastly, suppose that in addition to a $p$-fold rotational symmetry in the plane perpendicular to $e_{1}$, $x\in X$ also has an additional 2-fold rotational symmetry with respect to this plane (imagine a prism).
+Lastly, suppose that in addition to a $p$-fold rotational symmetry in the plane perpendicular to $e_{1}$, $x\in X$ also has an additional 2-fold rotational symmetry with respect to this plane (as for the triangle prism density functions in Section 7.3).
 Then $\text{stab}(x)$ is the non-abelian group generated by $\{R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\}$, where $R_{e_{1},\theta}R_{e_{2},\pi}=R_{e_{2},\pi}R_{e_{1},-\theta}$.
@@
 A.3 Miscellaneous
+Here we provide several miscellaneous results which are used throughout the paper.
 Proposition A.11.
@@
 Suppose $V=\{v_{j}\}_{j=0}^{n}\subset\mathbb{S}^{1}\subset\mathbb{R}^{2}$ such that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$.
-Then $V$ has a unique weighted Karcher mean with respect to the geodesic distance on $\mathbb{S}^{1}$ (and any weights).
+Then $V$ has a unique weighted Karcher mean with respect to the geodesic distance on $\mathbb{S}^{1}$ and any weights.
+In particular, given weights $w_{0},...,w_{n}$, the unique Karcher mean is given by
+$\mu=\exp\left(\sum_{j=0}^{n}w_{j}\log(v_{j})\right)$
+where $\log$ denotes any branch of $\log$ whose domain contains $V$.
 Proof.
@@
 Since $\mathbb{S}^{1}$ is geodesically convex on any subset of diameter less than $\pi$, $V$ has a unique weighted Karcher mean with respect to any weights.
+In particular, we can choose lifts $\{\phi_{j}\}_{j=0}^{n}\subset\mathbb{R}$ of $\{v_{j}\}_{j=0}^{n}$ which lie in an interval $I\subset\mathbb{R}$ of length less than $\frac{1}{2}$.
+The Frechet variance of $V$ around any $p=e^{i\phi_{p}}\in\exp(I)\subset\mathbb{S}^{1}$ is given by
+$F_{V}(p)$ $=\sum_{j=0}^{n}w_{j}d_{\mathbb{S}^{1}}(p,v_{j})^{2}$
+$=\sum_{j=0}^{n}w_{j}d_{\mathbb{S}^{1}}(e^{i\phi_{p}},e^{i\phi_{j}})^{2}$
+$=\sum_{j=0}^{n}w_{j}d_{\mathbb{R}}(\phi_{p},\phi_{j})^{2}$
+which is the weighted variance of $\{\phi_{j}\}_{j=0}^{n}$ about $\phi_{p}$.
+This is uniquely minimized at $\phi_{p}=\sum_{j=0}^{n}w_{j}\phi_{j}$, so $F_{V}$ has a minimizer at $p=\exp(\sum_{j=0}^{n}w_{j}\phi_{j})$.
+Since we have shown that weighted Karcher mean is unique under the hypotheses, $p$ is the unique weighted Karcher mean of $V$.